### v3

# Rigid graph products

Abstract.
We prove rigidity properties for von Neumann algebraic graph products. We introduce the notion of rigid graphs and define a class of II1-factors named $\mathcal{C}_{\textrm{Rigid}}$. For von Neumann algebras in this class we show a unique rigid graph product decomposition. In particular, we obtain unique prime factorization results and unique free product decomposition results for new classes of von Neumann algebras. Furthermore, we show that for many graph products of II1-factors, including the hyperfinite II1-factor, we can, up to a constant 2, retrieve the radius of the graph from the graph product. We also prove several technical results concerning relative amenability and embeddings of (quasi)-normalizers in graph products. Furthermore, we give sufficient conditions for a graph product to be nuclear and characterize strong solidity, primeness and free-indecomposability for graph products.
1. Introduction
The advent of Popa’s deformation-rigidity theory has led to major applications to the structure of von Neumann algebras and their decomposability properties for crossed products, tensor products and free products. For instance, in [55] Ozawa and Popa studied the notion of strongly solid von Neumann algebras (see Definition 7.1) and proved that the free group factors possess this property. Consequently, these von Neumann algebras do not admit certain crossed product decompositions, and they are prime factors (see Definition 8.1), meaning that they can not decompose as tensor products in non-trivial way (see also [54], [60]). More general prime factorization results were then obtained in e.g. [20, 22, 41, 47, 56, 58, 65, 66]. In the same spirit, decompositions of von Neumann algebras in terms of free products and Kurosh type results were studied in e.g. [42, 44, 52, 58].
This paper contributes to decomposability and rigidity results for von Neumann algebras that appear as graph products. The von Neumann algebraic graph product was introduced in [14],[51]. Given a simple graph $\Gamma$ and for each vertex $v\in\Gamma$ a von Neumann algebra $M_{v}$ with a normal faithful state $\varphi_{v}$, one can construct the graph product von Neumann algebra $M_{\Gamma}:=*_{v,\Gamma}(M_{v},\varphi_{v})$, see Section 2.6 for the exact definition. The construction of graph products naturally generalizes the notion of free products and tensor products. Indeed, if $\Gamma$ has no edges then $M_{\Gamma}$ is simply the von Neumann algebraic free product $_{v\in\Gamma}(M_{v},\varphi_{v})$, while if $\Gamma$ is a complete graph then $M_{\Gamma}$ is the tensor product $\overline{\bigotimes}_{v\in\Gamma}M_{v}$. The construction of graph products was originally introduced by Green in [35] for the setting of groups. For groups, the graph product $G_{\Gamma}=*_{v,\Gamma}G_{v}$ generalizes both free products and direct sums. The two notions of graph products naturally correspond since for group von Neumann algebras we have $\mathcal{L}(*_{v,\Gamma}G_{v})=*_{v,\Gamma}\mathcal{L}(G_{v})$.
We will prove rigidity results for graph products of von Neumann algebras. We first discuss our main result A which establishes unique rigid graph product decompositions. Thereafter, we give new unique prime factorization results and unique free product decomposion results. Furthermore, we state results that characterize primeness, free indecomposability and strong solidity for graph products. Hereafter, we present other main results that are needed in the proofs. Last, we give an overview of the structure of the paper.
1.1. Unique rigid graph product decomposition
Our main result, A, concerns the question whether from the graph product $_{v,\Gamma}(M_{v},\varphi_{v})$ we can, under some conditions, retrieve the graph $\Gamma$ and the vertex von Neumann algebras $M_{v}$. Such questions have already been studied for graph products of groups. In [35, Theorem 4.12] Green showed the following rigidity result, which for graph products $_{v,\Gamma}G_{v}$ of prime cycles $G_{v}$ asserts that the graph $\Gamma$ and the vertex groups $G_{v}$ can be retrieved from the graph product group.
Theorem (Green). Let $\Gamma,\Lambda$ be finite graphs, $G_{\Gamma}:=*_{v,\Gamma}G_{v}$ and $H_{\Lambda}:=*_{w,\Lambda}H_{w}$ be graph products of groups $G_{v}:=\mathbb{Z}/p_{v}\mathbb{Z}$ and $H_{w}:=\mathbb{Z}/q_{w}\mathbb{Z}$ with some prime numbers $p_{v},q_{w}$. If $G_{\Gamma}$ and $H_{\Lambda}$ are isomorphic, then there is a graph isomorphism $\alpha:\Gamma\to\Lambda$ such that $H_{\alpha(v)}\simeq G_{v}$.
In the current paper we prove an analogy of this result for graph products $M_{\Gamma}=*_{v,\Gamma}(M_{v},\tau_{v})$ of tracial von Neumann algebras $(M_{v},\tau_{v})$. Earlier rigidity results for von Neumann algebraic graph products have already been proven in [17, Theorem A and C] for group von Neumann algebras $M_{v}:=\mathcal{L}(G_{v})$ for certain discrete property (T) groups $G_{v}$ and for graphs $\Gamma$ from a class called CC1. In our main result, A, we also prove rigidity results for graph products of von Neumann algebras $M_{\Gamma}=*_{v,\Gamma}(M_{v},\tau_{v})$. Our result compares to [17, 18] as follows. On the one hand we cover a much richer class of graphs than CC1 and our vertex von Neumann algebras $M_{v}$ come from a different class than [17, 18]. In our paper $M_{v}$ are not even necessarily group von Neumann algebras. On the other hand the type of rigidity obtained in [17, 18] is stronger as it recovers the groups up to isomorphism, and not just the von Neumann algebras. Furthermore, [17, 18] obtains a so-called superrigidity result, meaning that the group can be recovered from an isomorphism of $\mathcal{L}(G)$ with any other group von Neumann algebra, whereas our rigidity results are usually for an isomorphism of two von Neumann algebras in the class $\mathcal{C}_{{\rm Rigid}}$ introduced below. Such a superrigidity result is simply not true in the context of the current paper as we argue in Remark 6.6.
The condition we impose on the vertex von Neumann algebras $M_{v}$ is that they lie in the class $\mathcal{C}_{{\rm Vertex}}$ of all non-amenable II1-factors that satisfy property strong (AO) (see Definition 6.4) and have separable preduals. This is a natural class of von Neumann algebras including the (interpolated) free group factors $\mathcal{L}(\mathbb{F}_{t})$ for $1<t<\infty$, the group von Neumann algebras $\mathcal{L}(G)$ of non-amenable hyperbolic icc groups $G$ [38], $q$-Gaussian von Neumann algebras $M_{q}(H_{\mathbb{R}})$ associated with real Hilbert spaces $H_{\mathbb{R}}$ with $2\leq\dim(H_{\mathbb{R}})<\infty$ [7, Remark 4.5], [49], free orthogonal quantum groups [73] as well as several common series of easy quantum groups and free wreath products of quantum groups [12, Theorem 0.5].
The condition we impose on the graph $\Gamma$ is that each vertex $v$ in $\Gamma$ satisfies $\textrm{Link}(\textrm{Link}(v))=\{v\}$ (for the definition of Link see Section 2.4). Such graphs, which we call rigid, form a large natural class of graphs containing for example complete graphs and cyclic graphs with at least 5 vertices. We also observe that all graphs in CC1 are rigid (see Remark 3.10). We stress that some restrictions on the graphs need to be imposed. Indeed, for general graphs $\Gamma$, and graph products $M_{\Gamma}=*_{v,\Gamma}(M_{v},\tau_{v})$ with $M_{v}\in\mathcal{C}_{{\rm Vertex}}$, it is not possible to retrieve the graphs $\Gamma$ from $M_{\Gamma}$ (see Remark 6.6). This is due to the fact that the free product $(M_{v},\tau_{v})*(M_{w},\tau_{w})$ of factors $M_{v},M_{w}\in\mathcal{C}_{{\rm Vertex}}$ again lies again in the class $\mathcal{C}_{{\rm Vertex}}$ (see Remark 6.5).
We now state our main result which shows rigidity for the class $\mathcal{C}_{{\rm Rigid}}$ of all graph products $M_{\Gamma}=*_{v,\Gamma}(M_{v},\tau_{v})$ with $\Gamma$ non-empty, rigid graphs and with $M_{v}\in\mathcal{C}_{{\rm Vertex}}$.
Theorem A (Theorem 6.18 and Theorem 8.5).
Let $\Gamma$ be a rigid graph and for $v\in\Gamma$ let $M_{v}$ be von Neumann algebras in class $\mathcal{C}_{{\rm Vertex}}$ with faithful normal state $\tau_{v}$. Let $M_{\Gamma}=\ast_{v,\Gamma}(M_{v},\tau_{v})$ be their graph product. Suppose there is another graph product decomposition of $M_{\Gamma}$ over another rigid graph $\Lambda$ and other von Neumann algebras $N_{w}\in\mathcal{C}_{{\rm Vertex}}$, $w\in\Lambda$, i.e. $M_{\Gamma}=*_{w,\Lambda}(N_{w},\tau_{w})$. Then there is a graph isomorphism $\alpha:\Gamma\rightarrow\Lambda$, and for each $v\in\Gamma$ there is a unitary $u_{v}\in M_{\Gamma}$ and a real number $0<t_{v}<\infty$ such that:
(1.1) $M_{\textrm{Star}(v)}=u_{v}^{*}N_{\textrm{Star}(\alpha(v))}u_{v}$ $and$ $M_{v}\simeq N_{\alpha(v)}^{t_{v}}.$
Furthermore, for the connected component $\Gamma_{v}\subseteq\Gamma$ of any vertex $v\in\Gamma$, we have $M_{\Gamma_{v}}=u_{v}^{*}N_{\alpha(\Gamma_{v})}u_{v}$; and for any irreducible component $\Gamma_{0}\subseteq\Gamma$, $\exists t_{0}\in(0,\infty)$ such that $M_{\Gamma_{0}}\simeq N_{\alpha(\Gamma_{0})}^{t_{0}}$.
We remark that in the setting of [17, Theorem 7.9], it is possible to obtain unitary conjugacy between the vertex von Neumann algebras $M_{v}=\mathcal{L}(G_{v})$. In our setting it is generally only possible to obtain isomorphisms up to amplification between the vertex von Neumann algebras. The reason is that the tensor product $M_{v}\overline{\otimes}M_{w}$ of II1-factors is isomorphic to the tensor product $M_{v}^{t}\overline{\otimes}M_{w}^{1/t}$ for any $0<t<\infty$. For certain subgraphs $\Gamma_{0}\subseteq\Gamma$ we do however obtain unitary conjugacy of the graph products $M_{\Gamma_{0}}$ to $N_{\alpha(\Gamma_{0})}$ inside $M_{\Gamma}$. Indeed, this is the case when $\Gamma_{0}$ is a connected component of $\Gamma$ or is of the form $\Gamma_{0}=\textrm{Star}(v)$ for some vertex $v$ of $\Gamma$ (for the definition of Star see Section 2.4). Moreover, for $\Gamma_{0}$ an irreducible component of $\Gamma$ we are able to show that $M_{\Gamma_{0}}$ is isomorphic to a amplification of $N_{\alpha(\Gamma_{0})}$.
1.2. Unique prime factorization
For classes of von Neumann algebras we are interested in unique prime factorization results. Recall that a II1-factor $M$ is prime if it can not decompose as a tensor product $M=M_{1}\overline{\otimes}M_{2}$ of diffuse factors $M_{1},M_{2}$. The first example of a prime factor was given by Popa in [60]. Thereafter, Ge showed in [34] that $\mathcal{L}(\mathbb{F}_{n})$ is a prime factor for $n\geq 2$ by computing Voiculescu’s free entropy. Later, in [54] Ozawa introduced a new property, called solidity, which for non-amenable factors implies primeness. He showed that all finite von Neumann algebras satisfying the Akemann-Ostrand property are solid. We note that in particular all von Neummann algebras in $\mathcal{C}_{{\rm Vertex}}$ are prime. There are many more examples of prime factors, see e.g. [9, 19, 20, 22, 30, 58, 65, 66].
Given a class $\mathcal{C}$ of von Neumann algebras, a natural question is whether any von Neumann algebra $M\in\mathcal{C}$ has a tensor product decomposition $M=M_{1}\overline{\otimes}\cdots\overline{\otimes}M_{m}$ for some $m\geq 1$ and prime factors $M_{1},\ldots,M_{m}\in\mathcal{C}$, which is called prime factorization inside $\mathcal{C}$, and whether the prime factorization is unique. This is to say, given another prime factorization $M=N_{1}\overline{\otimes}\cdots\overline{\otimes}N_{n}$, with $n\geq 1$ and prime factors $N_{1},\ldots,N_{n}\in\mathcal{C}$, do we have $n=m$ and, up to permutation of the indices, any $M_{i}$ is isomorphic to an amplification of $N_{i}$. The first unique prime factorization (UPF) results were established by Ozawa and Popa in [56] for tensor products of group von Neumann algebras $\mathcal{L}(G_{v})$ for certain groups $G_{v}$. The groups they considered included non-amenable, icc groups that are hyperbolic or are discrete subgroups of connected simple Lie groups of rank one. Later, in [47] Isono studied UPF results for free quantum group factors. Thereafter, by combining results from [56] and [47], Houdayer and Isono showed in [41] more general UPF results for tensor products of factors from a class called $\mathcal{C}_{{\rm(AO)}}$. We note that our class $\mathcal{C}_{{\rm Vertex}}$ is very similar to $\mathcal{C}_{{\rm(AO)}}$ and that $\mathcal{C}_{{\rm Vertex}}\subseteq\mathcal{C}_{{\rm(AO)}}$. In the setting of graph products, UPF results have been obtained in [19, Theorem 6.16] under the condition that the vertex von Neumann algebras are group von Neumann algebras.
We observe that we can use A to obtain UPF results. Indeed, let $\mathcal{C}_{{\rm Complete}}$ be the class of all tensor products of von Neumann algebras in $\mathcal{C}_{{\rm Vertex}}$. If in A we restrict our attention to complete graphs (which are rigid) then we precisely obtain UPF results for the class $\mathcal{C}_{{\rm Complete}}$ (see Corollary 6.20). This partially retrieves the UPF results from [41]. To obtain more general UPF results we prove the following result which characterizes primeness for graph products of II1-factors (see also Theorems 8.11 and 8.12 in the case the vertex von Neumann algebras are not II1-factors).
Theorem B (Theorem 8.4).
Let $\Gamma$ be a finite simple graph of size $|\Gamma|\geq 2$. For any $v\in\Gamma$, let $M_{v}$ be a II1-factor. The graph product $M_{\Gamma}=*_{v,\Gamma}(M_{v},\tau_{v})$ is prime if and only if $\Gamma$ is irreducible.
We then use A and B to prove the following theorem which covers UPF results for a new class of von Neumann algebras (see Remark 8.7).
Theorem C (Theorem 8.6).
Any von Neumann algebra $M\in\mathcal{C}_{{\rm Rigid}}^{f}$ has a prime factorization inside $\mathcal{C}_{{\rm Rigid}}^{f}$, i.e.
(1.2) $M=M_{1}\overline{\otimes}\cdots\overline{\otimes}M_{m},$
for some $1\leq m<\infty$ and prime factors $M_{1},\ldots,M_{m}\in\mathcal{C}_{{\rm Rigid}}^{f}$.
Suppose $M$ has another prime factorization inside $\mathcal{C}_{{\rm Rigid}}^{f}$, i.e.
(1.3) $M=N_{1}\overline{\otimes}\cdots\overline{\otimes}N_{n},$
for some $1\leq n<\infty$, and prime factors $N_{1},\ldots,N_{n}\in\mathcal{C}_{{\rm Rigid}}^{f}$. Then $m=n$ and there is a permutation $\sigma$ of $\{1,\ldots,m\}$ such that $M_{i}$ is stably isomorphic to $N_{\sigma(i)}$ for $1\leq i\leq m$.
1.3. Unique free product decomposition
In [52] Ozawa extended the results [56] for tensor products to the setting of free products. In particular, he showed for $M=M_{1}*\cdots*M_{m}$ a von Neumann algebraic free products of non-prime, non-amenable, semiexact finite factors $M_{1},\ldots,M_{m}$ that if $M=N_{1}*\cdots*N_{n}$ is another free product decomposition into non-prime, non-amenable, semiexact finite factors $N_{1},\ldots,N_{n}$, then $m=n$ and, up to permutation of the indices, $M_{i}$ unitarily conjugates to $N_{i}$ inside $M$ for each $1<i<m$. This can be seen as a von Neumann algebraic version of the Kurosh isomorphism theorem [48], which states that any discrete group uniquely decomposes as a free product of freely-indecomposable groups. Versions of Ozawa’s result were later shown for other classes of von Neumann algebras, see [2],[44],[58]. In [42] these results were then extended by Houdayer and Ueda to a single, large class of von Neumann algebras. Other Kurosh type theorems have recently been obtained in [29, Corollary 8.1], [27, Corollary 1.8].
In the current paper we obtain unique free product decomposition results for a new class of von Neumann algebras. First, we prove the following result which characterizes precisely when a graph product $M_{\Gamma}=*_{v,\Gamma}(M_{v},\tau_{v})$ can decompose as tracial free product of II1-factors.
Theorem D (Theorem 9.1).
Let $\Gamma$ be a simple graph of size $|\Gamma|\geq 2$, and for each $v\in\Gamma$ let $M_{v}$ be II1-factor with separable predual. Then the graph product $M_{\Gamma}:=*_{v,\Gamma}(M_{v},\tau_{v})$ can decompose as a tracial free product $M_{\Gamma}=(M_{1},\tau_{1})*(M_{2},\tau_{2})$ of II1-factors $M_{1}$$M_{2}$ if and only if $\Gamma$ is not connected.
Using A and D we obtain unique free product decomposition for the class $\mathcal{C}_{{\rm Rigid}}\setminus\mathcal{C}_{{\rm Vertex}}$.
Theorem E (Theorem 9.2).
Any von Neumann algebra $M\in\mathcal{C}_{{\rm Rigid}}\setminus\mathcal{C}_{{\rm Vertex}}$ can decompose as a tracial free product inside $\mathcal{C}_{{\rm Rigid}}\setminus\mathcal{C}_{{\rm Vertex}}$, i.e.
(1.4) $M=*_{i\in I}M_{i},$
for some index set $I$, and for every $i\in I$ a factor $M_{i}\in\mathcal{C}_{{\rm Rigid}}\setminus\mathcal{C}_{{\rm Vertex}}$ that can not decompose as any tracial free product of II1-factors.
Suppose $M$ can decompose as another tracial free product inside $\mathcal{C}_{{\rm Rigid}}\setminus\mathcal{C}_{{\rm Vertex}}$, i.e.
$M=*_{j\in J}N_{j}$
for another index set $J$ and for every $j\in J$ a factor $N_{j}\in\mathcal{C}_{{\rm Rigid}}\setminus\mathcal{C}_{{\rm Vertex}}$ that can not decompose as tracial free product of II1-factors. Then $|I|=|J|$ and there is a bijection $\sigma$ between $J$ and $I$ such that $N_{j}$ unitarily conjugates to $M_{\sigma(j)}$ in $M$.
Let us remark that von Neumann algebras in the class $\mathcal{C}_{{\rm Complete}}\setminus\mathcal{C}_{{\rm Vertex}}$ are examples of non-prime, non-amenable, semiexact, finite factors. Thus Ozawa’s result in particular asserts a unique free product decomposition for free products of factors in $\mathcal{C}_{{\rm Complete}}\setminus\mathcal{C}_{{\rm Vertex}}$. The same result is also covered by E since any free product of factors in $\mathcal{C}_{{\rm Complete}}\setminus\mathcal{C}_{{\rm Vertex}}$ lies in the class $\mathcal{C}_{{\rm Rigid}}\setminus\mathcal{C}_{{\rm Vertex}}$. We observe that, in contrast to Ozawa’s result, in E it is possible for the factors $M_{i},i\in I$ to be prime. More generally, we remark that the result of E is not covered by the result from [42] (see Remark 9.4). Thus our examples of unique free product decompositions are again new.
1.4. Graph radius rigidity
We are interested in the question whether from the graph product $M_{\Gamma}=*_{v,\Gamma}(M_{v},\tau_{v})$ of II1-factors $M_{v}$ we can retrieve the radius of the graph $\Gamma$. To study this question we introduce the notion of the radius of a von Neumann algebra $M$ (see Definition 10.3). As we show in the following theorem, we are in many cases able to estimate the radius of the von Neumann algebra $M_{\Gamma}$ with the radius of the graph $\Gamma$.
Theorem F (Theorem 10.6 and Theorem 10.11).
Let $\Gamma$ be a non-complete simple graph. For $v\in\Gamma$ let $M_{v}$ be a II1-factor and let $M_{\Gamma}=*_{v,\Gamma}(M_{v},\tau_{v})$ be the tracial graph product. Suppose one of the following holds true:
  1. (1)
    For all $v\in\Gamma$ the vertex algebra $M_{v}$ possesses strong (AO) and has separable predual.
  2. (2)
    For all $v\in\Gamma$ we have $M_{v}=\mathcal{L}(G_{v})$ for some countable icc group $G_{v}$.
Then,
$\operatorname{Radius}(\Gamma)-2\leq\operatorname{Radius}(M_{\Gamma})\leq\max\{2,\operatorname{Radius}(\Gamma)\}.$
The above result allows us to distinguish certain von Neumann algebras coming from graph products. In particular, for graph products $R_{\Gamma_{i}}=*_{v,\Gamma_{i}}(R_{v},\tau_{v})$ of hyperfinite II1-factors $R_{v}$, we are able to show that $R_{\Gamma_{1}}\not\simeq R_{\Gamma_{2}}$ whenever $2\leq\operatorname{Radius}(\Gamma_{1})<\operatorname{Radius}(\Gamma_{2})-2$ (see Remark 10.7).
We remark that when $\Lambda_{i}$ for $i=1,2$ are graphs of size $2\leq|\Lambda_{1}|=:n<|\Lambda_{2}|=:m$ and with no edges, then $R_{\Lambda_{1}}=\mathcal{L}(\mathbb{F}_{n})$ and $R_{\Lambda_{2}}=\mathcal{L}(\mathbb{F}_{m})$ by [32]. In this case, it is very hard to distinguish $R_{\Lambda_{1}}$ from $R_{\Lambda_{2}}$ as this is precisely the free factor problem. Of course, F is of no use here since $\operatorname{Radius}(\Lambda_{1})=\infty=\operatorname{Radius}(\Lambda_{2})$.
1.5. Strong solidity
For a finite von Neumann algebra $M$ the notion of strong solidity was introduced by Ozawa and Popa in [55]. This property, which in particular implies solidity, asserts that for any diffuse amenable von Neumann subalgebra $A\subseteq M$, its normalizers $\operatorname{Nor}_{M}(A)$ generates a von Neumann algebra that is amenable. This property implies that for a non-amenable von Neumann algebra it does not have a Cartan subalgebra, and hence can not decompose as a crossed product in a natural way. In [55], it was shown in that the free group factors $\mathcal{L}(\mathbb{F}_{t})$ are strong solidity. Nowadays, many examples of strongly solid von Neumann algebras are known, see e.g. [12, 21, 28, 45, 59]. Moreover, we remark that using the resolution of the Peterson-Thom conjecture (see [36], [4] ,[3]), it has been shown in [37] that the free group factors even satisfy a strengthened version of strong solidity.
In this paper we study strong solidity for graph products of von Neumann algebras. In [6] strong solidity was characterized for group von Neumann algebras $\mathcal{L}(\mathcal{W}_{\Gamma})$ of right-angled Coxeter groups $\mathcal{W}_{\Gamma}$. Using similar techniques, we characterize strong solidity for arbitrary graph products over finite graphs.
Theorem G (Theorem 7.7).
Let $\Gamma$ be a finite simple graph, and for each $v\in\Gamma$ let $M_{v}$ ($\not$$\mathbb{C}$) be a von Neumann algebra with normal faithful trace $\tau_{v}$. Then $M_{\Gamma}$ is strongly solid if and only if the following conditions are satisfied:
  1. (1)
    For each vertex $v\in\Gamma$ the von Neumann algebra $M_{v}$ is strongly solid;
  2. (2)
    For each subgraph $\Lambda\subseteq\Gamma$ with $M_{\Lambda}$ non-amenable, we have that $M_{\textrm{Link}(\Lambda)}$ is not diffuse;
  3. (3)
    For each subgraph $\Lambda\subseteq\Gamma$ with $M_{\Lambda}$ non-amenable and diffuse, we have moreover that $M_{\textrm{Link}(\Lambda)}$ is atomic.
We remark that for a large class of vertex von Neumann algebras $M_{v}$ it can be verified whether the conditions (1),(2) and (3) hold true for the graph products $M_{\Lambda}$ and $M_{\textrm{Link}(\Lambda)}$. In particular, for right-angled Hecke von Neumann algebras this characterizes strong solidity (using Theorem 7.12 from [15], [64]). Partial results in this direction had already been obtained in [8] and [11].
1.6. Other results
The proofs of the stated theorems require several main results that are of independent interest, which we present here. Firstly, we give sufficient conditions for a graph product of unital C-algebras to be nuclear. This is a generalizion of Ozawa’s result for free products [53] and is needed in the proof of A.
Theorem H (Theorem 4.4).
Let $\Gamma$ be a simple graph. Let $A_{\Gamma}=*^{\min}_{v,\Gamma}(A_{v},\varphi_{v})$ be the reduced C-algebraic graph product of nuclear, unital C-algebras $A_{v}$ with GNS-faithful state $\varphi_{v}$. Let $\mathcal{H}_{v}:=L^{2}(A_{v},\varphi_{v})$ and let $\pi_{v}:A_{v}\to\mathbb{B}(\mathcal{H}_{v})$ be the GNS-representation. If for any $v\in\Gamma$, $\pi_{v}(A_{v})$ contains the space of compact operators $\mathbb{K}(\mathcal{H}_{v})$, then $A_{\Gamma}$ is nuclear.
The following result is the graph product analogue of [41, Theorem 5.1] and [52, Theorem 3.3.], and is crucial in the proof of A for establishing the graph isomorphism.
Theorem I (Theorem 6.15).
Let $\Gamma$ be a simple graph, $(M_{\Gamma},\tau)=\ast_{v,\Gamma}(M_{v},\tau_{v})$ be the graph product of finite von Neumann algebras $M_{v}$ that satisfy condition strong (AO) and have separable preduals. Let $Q\subseteq M_{\Gamma}$ be a diffuse von Neumann subalgebra. At least one of the following holds:
  1. (1)
    The relative commutant $Q^{\prime}\cap M_{\Gamma}$ is amenable;
  2. (2)
    There exists $\Gamma_{0}\subseteq\Gamma$ such that $Q\prec_{M_{\Gamma}}M_{\Gamma_{0}}$ and $\textrm{Link}(\Gamma_{0})\not$$\varnothing$.
The following result concerning relative amenability extends [6, Theorem 3.7] to the setting of graph product. This result is needed in the proof of the characterizations given in B, D and G.
Theorem J (Theorem 5.3).
Let $\Gamma$ be a simple graph with subgraphs $\Gamma_{1},\Gamma_{2}\subseteq\Gamma$. For each $v\in\Gamma$ let $(M_{v},\tau_{v})$ be a von Neumann algebra with a normal faithful trace. Let $P\subset M_{\Gamma}$ be a von Neumann subalgebra that is amenable relative to $M_{\Gamma_{i}}$ inside $M_{\Gamma}$ for $i=1,2$. Then $P$ is amenable relative to $M_{\Gamma_{1}\cap\Gamma_{2}}$ inside $M_{\Gamma}$.
1.7. Paper overview
In Section 2 we recall Popa’s interwtining-by-bimodule technique and introduce our notation for simple graphs and for graph products. In Section 3 we introduce the notion of rigid graphs and study some basic properties. Here, we also define graph products of graphs and precisely characterize when a graph product of graphs is rigid. In Section 4 we prove H which establishes sufficient conditions for a graph product to be nuclear. In Section 5 we prove some technical results concerning graph products. In particular, using calculation for iterated conditional expectations in graph products we prove J regarding relative-amenability, and prove some embedding results for graph products. In Section 6 we prove I which we then use to prove the major part of A. In Section 7 we prove G which characterizes strong solidity for graph products. In Section 8 we prove B which characterizes primeness in graph products. Moreover, we also complete the proof of A and we prove C which establishes UPF results for the class $\mathcal{C}_{{\rm Rigid}}$. In Section 9 we prove D which characterizes free-indecomposability for graph products and we prove E which establishes unique free product decomposition results for the class $\mathcal{C}_{{\rm Rigid}}\setminus\mathcal{C}_{{\rm Vertex}}$. Last, in Section 10 we define the radius of a von Neumann algebra and prove F which for graph products provides good estimates on the radius of the graph.
Comments
After our results have appeared on the arXiv new rigidity results for graph products have appeared in [13], [31] and [39] for new classes of graph products. In particular [13], [31] also consider infinite graphs. After these results appeared we revised this manuscript and updated earlier statements for finite graphs that hold for infinite graphs as well with the same proof.
Acknowledgements
The authors wish to thank David Jekel and Ben Hayes for communicating several suggestions for improvement of our paper.
2. Preliminaries
2.1. General notation
For a Hilbert space $\mathcal{H}$ we denote $\mathbb{B}(\mathcal{H})$ for the space of bounded operators on $\mathcal{H}$ and denote $\mathbb{K}(\mathcal{H})$ for the space of compact operators on $\mathcal{H}$. For a von Neumann algebra $M$ we denote $\mathcal{Z}(M)=M\cap M^{\prime}$ for the center and denote $\mathcal{U}(M)$ for the set of all unitaries. Furthermore, for $0<t<\infty$ and a II1-factor $M$ we denote $M^{t}$ for the amplification of $M$ by $t$. For $u\in\mathcal{U}(\mathbb{B}(\mathcal{H}))$ we write $\operatorname{Ad}_{u}(x)=uxu^{\ast},x\in\mathbb{B}(\mathcal{H})$. Inclusions of von Neumann algebras are always assumed to be unital inclusions. We write $1_{M}$ for the unit of a von Neumann algebra $M$. For a von Neumann subalgebra $A\subseteq 1_{A}M1_{A}$ we denote the set of normalizers and quasi-normalizers respectively by
$\operatorname{Nor}_{M}(A)$ $:=\{u\in\mathcal{U}(M):u^{*}Au=A\},$
$\operatorname{qNor}_{M}(A)$ $:=\{x\in M:\exists x_{1},\ldots,x_{n},y_{1},\ldots,y_{m}\in M\text{ s.t. }Ax\subseteq\sum_{i=1}^{n}x_{i}A\text{ and }xA\subseteq\sum_{i=1}^{m}Ay_{i}\}.$
Tensor products of von Neumann algebras are defined as the strong closure of the their spacial tensor products.
2.2. Jones extension
Let $(M,\tau)$ be a tracial von Neumann algebra and $Q\subseteq M$ a von Neumann subalgebra. We denote $\mathbb{E}_{Q}:M\to Q$ for the trace-preserving conditional expectation, and denote $e_{Q}:L^{2}(M,\tau)\to L^{2}(Q,\tau)$ for its $L^{2}$-extension. We regard $M\subseteq\mathbb{B}(L^{2}(M,\tau))$ as a subalgebra, and denote $\langle M,e_{Q}\rangle$ for the Jones extension which is the von Neumann algebra generated by $M\cup\{e_{Q}\}$.
2.3. Popa’s intertwining-by-bimodule theory
We recall the following definition from the fundamental work of [61, 62]. In this section we let $M$ be a finite von Neumann algebra.
Definition 2.1 (Embedding $A\prec_{M}B$).
For von Neumann subalgebras $A\subseteq 1_{A}M1_{A},B\subseteq 1_{B}M1_{B}$ we will say that $A$ embeds in $B$ inside $M$ (denoted by $A\prec_{M}B$) if one of the following equivalent statements holds:
  1. (1)
    There exist projections $p\in A$, $q\in B$, a normal $\ast$-homomorphism $\theta:pAp\to qBq$ and a non-zero partial isometry $v\in qMp$ such that $\theta(x)v=vx$ for all $x\in pAp$;
  2. (2)
    There exists no net of unitaries $(u_{i})_{i}$ in $A$ such that for any $x,y\in 1_{A}M1_{B}$ we have that $\|\mathbb{E}_{B}(x^{\ast}u_{i}y)\|_{2}\rightarrow 0$;
  3. (3)
    There exists a Hilbert AA-$B$ bimodule $\mathcal{K}\subseteq L^{2}(M,\tau)$ such that $\dim_{B}\mathcal{K}<\infty$.
We say that $A$ embeds stably in $B$ inside $M$ (denoted by $A\prec_{M}^{s}B$) if for any projection $r\in A^{\prime}\cap M$ we have $Ar\prec_{M}B$.
Lemma 2.2 ( Lemma 2.4 in [30], see also [69]).
Let $(M,\tau)$ be a tracial von Neumann algebra and let $P\subseteq 1_{P}M1_{P}$, $Q\subseteq 1_{Q}M1_{Q}$ and $R\subseteq 1_{R}M1_{R}$ be von Neumann subalgebras. Then the following hold:
  1. (1)
    Assume that $P\prec_{M}Q$ and $Q\prec_{M}^{s}R$. Then $P\prec_{M}R$;
  2. (2)
    Assume that, for any non-zero projection $z\in\operatorname{Nor}_{1_{P}M1_{P}}(P)^{\prime}\cap 1_{P}M1_{P}\subseteq\mathcal{Z}(P^{\prime}\cap 1_{P}M1_{P})$, we have $Pz\prec_{M}Q$. Then $P\prec_{M}^{s}Q$.
In particular, we note that if $Q^{\prime}\cap 1_{Q}M1_{Q}$ is a factor and $P\prec_{M}Q$ and $Q\prec_{M}R$ then $P\prec_{M}R$.
2.4. Simple graphs
A simple graph $\Gamma$ is an undirected graph without double edges and without edges between a vertex and itself. We write $v\in\Gamma$ for saying that $v$ is a vertex of $\Gamma$. We write $\Lambda\subseteq\Gamma$ to say that $\Lambda$ is a subgraph of $\Gamma$, meaning that the vertices of $\Lambda$ form a subset of the vertices of $\Gamma$, and two vertices in $\Lambda$ share an edge iff they share an edge in $\Gamma$. For $v\in\Gamma$ we set
$\begin{split}\textrm{Link}_{\Gamma}(v)=&\{w\in\Gamma\mid v\textrm{ and }w\textrm{ share an edge}\},\\ \textrm{Star}_{\Gamma}(v)=&\{v\}\cup\textrm{Link}_{\Gamma}(v).\end{split}$
This entails in particular that $v\not\in\textrm{Link}_{\Gamma}(v)$. For $\Lambda\subseteq\Gamma$ we set $\textrm{Link}_{\Gamma}(\Lambda)=\bigcap_{v\in\Lambda}\textrm{Link}(v)$ with the convention $\textrm{Link}_{\Gamma}(\varnothing)=\Gamma$. We remark that $v\in\textrm{Link}_{\Gamma}(\textrm{Link}_{\Gamma}(v))$. When the graph $\Gamma$ is fixed, we will leave out the subscript and simply write $\textrm{Link}(v),\textrm{Star}(v),\textrm{Link}(\Lambda)$. We denote $|\Gamma|$ for the size of the graph, i.e. the number of vertices. We will call a graph $\Gamma$ irreducible if we can not write $\Gamma=\Gamma_{1}\cup\Gamma_{2}$ for some non-empty subgraphs $\Gamma_{1},\Gamma_{2}\subseteq\Gamma$ with $\textrm{Link}(\Gamma_{1})=\Gamma_{2}$. We call a graph $\Gamma$ connected if there exists a path between any two distinct vertices $v,w\in\Gamma$. An irreducible component of a graph $\Gamma$ is a non-empty subgraph $\Lambda\subseteq\Gamma$ that is irreducible and satisfies $\textrm{Link}_{\Gamma}(\Lambda)=\Gamma\setminus\Lambda$. A connected component of a graph $\Gamma$ is a subgraph $\Lambda\subseteq\Gamma$ that is connected and satisfies for $v\in\Lambda$ that $\textrm{Link}_{\Gamma}(v)\subseteq\Lambda$. We call a graph $\Gamma$ complete if any two vertices in $\Gamma$ share an edge. A complete subgraph $\Lambda\subseteq\Gamma$ is called a clique of $\Gamma$.
2.5. Right-angled Coxeter groups
Let $\Gamma$ be a simple graph. We let $\mathcal{W}_{\Gamma}$ be the right-angled Coxeter group corresponding to the graph $\Gamma$, that is, $\mathcal{W}_{\Gamma}$ is the group generated by the vertex set of $\Gamma$, subject to the relations that $vw=wv$ whenever $v,w$ share an edge in $\Gamma$ and $v^{2}=e$ with $e$ the group unit (thus $\mathcal{W}_{\Gamma}$ is equal to the graph product group $_{v,\Gamma}(\mathbb{Z}/2\mathbb{Z})$). For $\mathbf{v}\in\mathcal{W}_{\Gamma}$ we write $|\mathbf{v}|$ for the length of $\mathbf{v}$. For $\mathbf{v}_{1},\ldots,\mathbf{v}_{n}\in\mathcal{W}_{\Gamma}$ we say that the expression $\mathbf{v}=\mathbf{v}_{1}\cdots\mathbf{v}_{n}$ is reduced if $|\mathbf{v}|=|\mathbf{v}_{1}|+\ldots+|\mathbf{v}_{n}|$. We call a word $\mathbf{w}\in\mathcal{W}_{\Gamma}$ a clique word if it has a reduced expression $\mathbf{w}=w_{1}\cdots w_{n}$ with $w_{i}\in\Gamma$ such that $w_{i}$ commutes with $w_{j}$ for any $1\leq i,j\leq n$. For a reduced word $\mathbf{v}=v_{1}\cdots v_{n}\in\mathcal{W}_{\Gamma}$ we write
$\textrm{Link}_{\Gamma}(\mathbf{v})=\bigcap_{1\leq i\leq n}\textrm{Link}_{\Gamma}(v_{i}).$
For a subset $S\subseteq\mathcal{W}_{\Gamma}$ we denote
[…]
We apply this notation when $S\subseteq\Gamma\subseteq\mathcal{W}_{\Gamma}$ is a subgraph of $\Gamma$ or when $S=\{\mathbf{u}\}\subseteq\mathcal{W}_{\Gamma}$ is a singleton. In the latter case we simply write $\mathcal{W}(\mathbf{u})$ respectively $\mathcal{W}^{\prime}(\mathbf{u})$ for $\mathcal{W}(\{\mathbf{u}\})$ respectively $\mathcal{W}^{\prime}(\{\mathbf{u}\})$.
2.6. Graph products
We introduce the notion of operator algebraic graph products as in [14], [51], where we mainly follow the first reference.
[…]
Remark 2.3.
We recall the following notational convention, that appeared first in [15, top of page 8]. In case $(v_{1}^{\prime},\ldots,v_{n}^{\prime})$ is another reduced representative for $\mathbf{v}$ there is a unique permutation $\sigma$ of the $\{1,\ldots,n\}$ such that $v_{\sigma(i)}=v_{i}^{\prime}$ and if $i<j$ are such that $v_{i}=v_{j}$ then $\sigma(i)<\sigma(j)$. Thus there exists a unique unitary map
$U_{\sigma}:\mathring{\mathcal{H}}_{\mathbf{v}}:=\mathring{\mathcal{H}}_{v_{1}}\otimes\cdots\otimes\mathring{\mathcal{H}}_{v_{n}}\rightarrow\mathring{\mathcal{H}}_{\mathbf{v}}^{\prime}:=\mathring{\mathcal{H}}_{v_{1}^{\prime}}\otimes\cdots\otimes\mathring{\mathcal{H}}_{v_{n}^{\prime}},$
mapping $\xi_{1}\otimes\ldots\otimes\xi_{n}$ to $\xi_{\sigma(1)}\otimes\ldots\otimes\xi_{\sigma(n)}$. We will from this point identify the spaces $\mathring{\mathcal{H}}_{\mathbf{v}}$ and $\mathring{\mathcal{H}}_{\mathbf{v}}^{\prime}$ through $U_{\sigma}$ and omit $U_{\sigma}$ in the notation. We say that vectors are identified this way through shuffle equivalence.
We denote
$\mathcal{H}_{\Gamma}:=\bigoplus_{\mathbf{v}\in\mathcal{W}_{\Gamma}}\mathring{\mathcal{H}}_{\mathbf{v}},$
where we let $\Omega=1$ in $\mathring{\mathcal{H}}_{e}$ and $\mathring{\mathcal{H}}_{e}=\mathbb{C}$. $\mathcal{H}_{\Gamma}$ is the graph product of the vertex Hilbert spaces $(\mathcal{H}_{v},\xi_{v})$. Furthermore, for $S\subseteq\mathcal{W}_{\Gamma}$ we write
$\mathcal{H}(S):=\bigoplus_{\mathbf{v}\in\mathcal{W}(S)}\mathring{\mathcal{H}}_{\mathbf{v}}$ $and$ $\mathcal{H}^{\prime}(S):=\bigoplus_{\mathbf{v}\in\mathcal{W}^{\prime}(S)}\mathring{\mathcal{H}}_{\mathbf{v}}.$
Again, we apply this notation mainly when $S\subseteq\Gamma\subseteq\mathcal{W}_{\Gamma}$ is a subgraph of $\Gamma$ or when $S=\{\mathbf{u}\}\subseteq\mathcal{W}_{\Gamma}$ is a singleton and in the later case we simply write $\mathcal{H}(\mathbf{u})$ respectively $\mathcal{H}^{\prime}(\mathbf{u})$ for $\mathcal{H}(\{\mathbf{u}\})$ respectively $\mathcal{H}^{\prime}(\{\mathbf{u}\})$. For $\mathbf{v}_{1},\ldots,\mathbf{v}_{n}\in\mathcal{W}_{\Gamma}$ with $\mathbf{v}=\mathbf{v}_{1}\cdots\mathbf{v}_{n}$ reduced, we let
$\mathcal{Q}_{(\mathbf{v}_{1},\ldots,\mathbf{v}_{n})}:\mathring{\mathcal{H}}_{\mathbf{v}_{1}}\otimes\cdots\otimes\mathring{\mathcal{H}}_{\mathbf{v}_{n}}\to\mathring{\mathcal{H}}_{\mathbf{v}}$
be the natural unitary from [5]. Observe that simply $\mathcal{Q}_{(\mathbf{v}_{1},\ldots,\mathbf{v}_{n})}(\eta_{1}\otimes\cdots\otimes\eta_{n})\simeq\eta_{1}\otimes\cdots\otimes\eta_{n}$ up to shuffle equivalence when $\mathbf{v}_{i}\not=e$ for all $1\leq i\leq n$.
For a subgraph $\Lambda\subseteq\Gamma$ we define two unitaries:
$U_{\Lambda}:\mathcal{H}_{\Lambda}\otimes\mathcal{H}(\Lambda)\to\mathcal{H}_{\Gamma}$ as $U_{\Lambda}|_{\mathring{\mathcal{H}}_{\mathbf{u}}\otimes\mathring{\mathcal{H}}_{\mathbf{w}}}$ $=\mathcal{Q}_{(\mathbf{u},\mathbf{w})}$ $\text{ for }\mathbf{u}\in\mathcal{W}_{\Lambda},\mathbf{w}\in\mathcal{W}(\Lambda),$
$U_{\Lambda}^{\prime}:\mathcal{H}^{\prime}(\Lambda)\otimes\mathcal{H}_{\Lambda}\to\mathcal{H}_{\Gamma}$ as $U_{\Lambda}|_{\mathring{\mathcal{H}}_{\mathbf{u}}\otimes\mathring{\mathcal{H}}_{\mathbf{w}}}$ $=\mathcal{Q}_{(\mathbf{u},\mathbf{w})}$ $\text{ for }\mathbf{u}\in\mathcal{W}^{\prime}(\Lambda),\mathbf{w}\in\mathcal{W}_{\Lambda},$
and for $u\in\Gamma$ simply write $U_{u}$ respectively $U_{u}^{\prime}$ instead of $U_{\{u\}}$ respectively $U_{\{u\}}^{\prime}$.
As in [14] we define the embeddings
$\lambda_{v}:\mathbb{B}(\mathcal{H}_{v})\to\mathbb{B}(\mathcal{H}_{\Gamma})$ as $\lambda_{v}(a)$ $=U_{v}(a\otimes 1)U_{v}^{*},$
$\rho_{v}:\mathbb{B}(\mathcal{H}_{v})\to\mathbb{B}(\mathcal{H}_{\Gamma})$ as $\rho_{v}(a)$ $=U_{v}^{\prime}(1\otimes a)(U_{v}^{\prime})^{*}.$
As in [5], for a word $\mathbf{v}\in\mathcal{W}_{\Gamma}\setminus\{e\}$ with minimal representative $(v_{1},\ldots v_{n})$ we denote
[…]
for the algebraic tensor product and put $\mathring{\mathbf{A}}_{e}=\mathbb{C}1$. We denote $\mathbf{A}_{\Gamma}:=\bigoplus_{\mathbf{v}\in\mathcal{W}_{\Gamma}}\mathring{\mathbf{A}}_{\mathbf{v}}$ for the algebraic direct sum. We let $\lambda:\mathbf{A}_{\Gamma}\to\mathbb{B}(\mathcal{H}_{\Gamma})$ be the linear embedding, from [5, Equation (18)], which is defined as
$a_{1}\otimes\ldots\otimes a_{n}\mapsto\lambda_{v_{1}}(a_{1})\ldots\lambda_{v_{n}}(a_{n}),$
[…]
$\mathbb{E}_{A_{\Lambda}}(a_{1}\ldots a_{n})=\left\{\begin{array}[]{cc}a_{1}\ldots a_{n},&\forall i,v_{i}\in\Lambda;\\ 0,&\textrm{ otherwise.}\end{array}\right$.
Graph products of von Neumann algebras. In case $A_{v},v\in\Gamma$ is a von Neumannn algebra, usually denoted by $M_{v}$, equipped with normal faithful state $\varphi_{v}$, the graph product von Neumann algebra $M_{\Gamma}=*_{v,\Gamma}(M_{v},\varphi_{v})$ is constructed as the strong operator topology closure of the reduced C-algebraic graph product constructed above. Moreover, we define $\mathring{M}_{\mathbf{v}}$ as the closure of $\lambda(\mathring{\mathbf{M}}_{\mathbf{v}})$ in the strong operator topology. We also define the graph product state $\varphi_{\Gamma}(x)=\langle x\Omega,\Omega\rangle$ which restricts to $\varphi_{v}\circ\lambda_{v}^{-1}$ on $\lambda_{v}(M_{v})$. The conditional expectation $\mathbb{E}_{A_{\Lambda}}$ extends to a normal conditional expectation $\mathbb{E}_{M_{\Lambda}}:M_{\Gamma}\rightarrow M_{\Lambda}$.
[…]
(2.1) $A_{\Gamma}=A_{\Gamma_{1}}\ast_{A_{\Lambda}}A_{\Gamma_{2}},\qquad M_{\Gamma}=M_{\Gamma_{1}}\ast_{M_{\Lambda}}M_{\Gamma_{2}},$
for both the reduced C-algebraic and von Neumann algebraic graph products.
2.7. Creation and annihilation operators
Let $\Gamma$ be a simple graph as before and for each $v\in\Gamma$ let $M_{v}$ be a von Neumann algebra with faithful normal state $\varphi_{v}$. Let $M_{\Gamma}$ be the graph product von Neumann algebra. As in [5, Equation (27)], for $\mathbf{w}\in\mathcal{W}_{\Gamma}$ we denote
(2.4) $\mathcal{S}_{\mathbf{w}}:=\biggl{\{}(\mathbf{w}_{1},\mathbf{w}_{2},\mathbf{w}_{3})\in\mathcal{W}_{\Gamma}^{3}:\begin{array}[]{l}\mathbf{w}=\mathbf{w}_{1}\mathbf{w}_{2}\mathbf{w}_{3}\text{ reduced},\\ \mathbf{w}_{2}\text{ is a clique word.}\end{array}\biggr{\}}$
and $\mathcal{S}_{\Gamma}=\bigcup_{\mathbf{w}\in\mathcal{W}_{\Gamma}}\mathcal{S}_{\mathbf{w}}$. For $v\in\Gamma$ let $P_{v}\in\mathbb{B}(\mathcal{H}_{\Gamma})$ be the projection onto $\mathcal{H}(v)^{\perp}$. For $(\mathbf{w}_{1},\mathbf{w}_{2},\mathbf{w}_{3})\in\mathcal{S}_{\Gamma}$ we let $\lambda_{(\mathbf{w}_{1},\mathbf{w}_{2},\mathbf{w}_{3})}:\mathbf{M}_{\Gamma}\to\mathbb{B}(\mathcal{H}_{\Gamma})$ be the linear map defined in [5, Equation (26)]. This map satisfies $\lambda_{(\mathbf{w}_{1},\mathbf{w}_{2},\mathbf{w}_{3})}(a)=0$ for $a\in\mathring{\mathbf{M}}_{\mathbf{v}}$ with $\mathbf{v}\not=\mathbf{w}$. Furthermore, for $a\in\mathring{\mathbf{M}}_{\mathbf{w}}$ it is given as follows. Write $\mathbf{w}_{i}=w_{i,1}\cdots w_{i,n_{i}}$ for $i=1,2,3$, and let $a_{i}:=a_{i,1}\otimes\cdots\otimes a_{i,n_{i}}\in\mathring{\mathbf{M}}_{\mathbf{w}_{i}}$ for $i=1,2,3$ be such that $\lambda(a)=\lambda(a_{1})\lambda(a_{2})\lambda(a_{3})$. Then
$\begin{split}\lambda_{(\mathbf{w}_{1},\mathbf{w}_{2},\mathbf{w}_{3})}(a)=&(P_{w_{1,1}}a_{1,1}P_{w_{1,1}}^{\perp})\cdots(P_{w_{1,n_{1}}}a_{1,n_{1}}P_{w_{1,n_{1}}}^{\perp})\cdot\\ &(P_{w_{2,1}}a_{2,1}P_{w_{2,1}})\cdots(P_{w_{2,n_{1}}}a_{2,n_{2}}P_{w_{2,n_{2}}})\cdot\\ &(P_{w_{3,1}}^{\perp}a_{3,1}P_{w_{3,1}})\cdots(P_{w_{3,n_{1}}}^{\perp}a_{3,n_{3}}P_{w_{3,n_{3}}}).\end{split}$
Intuitively, for $a\in\mathbf{M}_{\Gamma}$ the operator $\lambda_{(\mathbf{w}_{1},\mathbf{w}_{2},\mathbf{w}_{3})}(a)$ is the part of $\lambda(a)$ that acts as annihilation on a word of type $\mathbf{w}_{3}^{-1}$, acts diagonally on a word of type $\mathbf{w}_{2}$ and acts as creation on a word of type $\mathbf{w}_{1}$. We will use the following two result. Lemma 2.4 was originally proven in [15].
Lemma 2.4 (Lemma 2.2. in [5]).
Let $\mathbf{w}\in\mathcal{W}_{\Gamma}$ and $a\in\mathring{M}_{\mathbf{w}}$. We have
$\lambda=\sum_{(\mathbf{w}_{1},\mathbf{w}_{2},\mathbf{w}_{3})\in\mathcal{S}_{\Gamma}}\lambda_{(\mathbf{w}_{1},\mathbf{w}_{2},\mathbf{w}_{3})}$  $\quad\text{ and }\quad\lambda(a)=\sum_{(\mathbf{w}_{1},\mathbf{w}_{2},\mathbf{w}_{3})\in\mathcal{S}_{\mathbf{w}}}\lambda_{(\mathbf{w}_{1},\mathbf{w}_{2},\mathbf{w}_{3})}(a).$
Lemma 2.5.
Let $\mathbf{w},\mathbf{v}\in\mathcal{W}_{\Gamma}$, let $(\mathbf{w}_{1},\mathbf{w}_{2},\mathbf{w}_{3})\in\mathcal{S}_{\mathbf{w}}$ and let $a\in\mathring{M}_{\mathbf{w}}$ and $\eta\in\mathring{\mathcal{H}}_{\mathbf{v}}$. Then $\lambda_{(\mathbf{w}_{1},\mathbf{w}_{2},\mathbf{w}_{3})}(a)\eta\in\mathring{\mathcal{H}}_{\mathbf{u}}$ where $\mathbf{u}=\mathbf{w}_{1}\mathbf{w}_{3}\mathbf{v}$. Moreover, if $\lambda_{(\mathbf{w}_{1},\mathbf{w}_{2},\mathbf{w}_{3})}(a)\eta$ is non-zero, then $\mathbf{v}$ and $\mathbf{u}$ start with $\mathbf{w}_{3}^{-1}\mathbf{w}_{2}$ and $\mathbf{w}_{1}\mathbf{w}_{2}$ respectively.
Proof.
Let $\mathbf{w},\mathbf{v},(\mathbf{w}_{1},\mathbf{w}_{2},\mathbf{w}_{3}),a$ and $\eta$ be a stated. Suppose $\lambda_{(\mathbf{w}_{1},\mathbf{w}_{2},\mathbf{w}_{3})}(a)\eta$ is non-zero. We may assume that $a$ is of the form $a=a_{1}a_{2}a_{3}$ with $a_{i}\in\mathring{M}_{\mathbf{w}_{i}}$. We use the comments stated after [5, Equation (25)]. These imply that $\lambda_{(e,e,\mathbf{w}_{3})}(a_{3})\eta\in\mathring{\mathcal{H}}_{\mathbf{w}_{3}\mathbf{v}}$ and that $\mathbf{v}$ starts with $\mathbf{w}_{3}^{-1}$. Moreover the comments then imply that $\lambda_{(e,\mathbf{w}_{2},\mathbf{w}_{3})}(a_{2}a_{3})\eta=\lambda_{(e,\mathbf{w}_{2},e)}(a_{2})\lambda_{(e,e,\mathbf{w}_{3})}(a_{3})\eta\in\mathring{\mathcal{H}}_{\mathbf{w}_{3}\mathbf{v}}$ and that $\mathbf{w}_{3}\mathbf{v}$ starts with $\mathbf{w}_{2}$. This already shows that $\mathbf{v}$ starts with $\mathbf{w}_{3}^{-1}\mathbf{w}_{2}$ (observe that $\mathbf{w}_{2}=\mathbf{w}_{2}^{-1}$). Last, the comments imply that $\lambda_{(\mathbf{w}_{1},\mathbf{w}_{2},\mathbf{w}_{3})}(a)=\lambda_{(\mathbf{w}_{1},e,e)}(a_{1})\lambda_{(e,\mathbf{w}_{2},\mathbf{w}_{3})}(a_{2}a_{3})\eta\in\mathring{\mathcal{H}}_{\mathbf{w}_{1}\mathbf{w}_{3}\mathbf{v}}$ and that $\mathbf{w}_{1}\mathbf{w}_{3}\mathbf{v}$ starts with $\mathbf{w}_{1}$. Hence $\mathbf{u}=\mathbf{w}_{1}\mathbf{w}_{3}\mathbf{v}$ starts with $\mathbf{w}_{1}\mathbf{w}_{2}$. ∎
3. Rigid graphs
In this section we introduce the notion of rigid graphs.
Definition 3.1 (Rigid graphs).
We say that a simple graph $\Gamma$ is rigid if for every $v\in\Gamma$ we have $\textrm{Link}_{\Gamma}(\textrm{Link}_{\Gamma}(v))=\{v\}$. When $|\Gamma|\geq 2$ this means in particular for each $v\in\Gamma$ that $\textrm{Link}_{\Gamma}(v)$ is not empty.
Example 3.2.
We give some examples of rigid graphs which are easy to check:
  1. (1)
    By the convention $\textrm{Link}_{\Gamma}(\varnothing)=\Gamma$ it follows that if $|\Gamma|=1$ then $\Gamma$ is rigid.
  2. (2)
    Any complete graph is rigid.
  3. (3)
    For $n\geq 2$ let $\mathbb{Z}_{n}=\{1,\ldots,n\}$ be the cyclic graph of length $n$, i.e. $i,j$ share an edge if and only if |ij|=1|i-j|=1 or $\{i,j\}=\{1,n\}$. Then for $n\geq 5$ the graph $\mathbb{Z}_{n}$ is rigid. Note also that $\mathbb{Z}_{2}$ and $\mathbb{Z}_{3}$ are rigid, but $\mathbb{Z}_{4}$ is not.
  4. (4)
    Consider $\mathbb{Z}$ as the infinite cyclic graph, i.e. $i,j$ share an edge in $\mathbb{Z}$ if and only if |ij|=1|i-j|=1. Then $\mathbb{Z}$ is rigid.
We will now define the notion of graph products of graphs, and construct a large variety of rigid graphs.
Definition 3.3.
Let $\Gamma$ be a simple graph and for each $v\in\Gamma$ let $\Lambda_{v}$ be a simple graph. We denote $\Lambda_{\Gamma}:=*_{v,\Gamma}\Lambda_{v}$ for the graph product of the graphs $\{\Lambda_{v}\}_{v\in\Gamma}$. This is defined as the graph with vertices set
(3.1) $\{(v,s):v\in\Gamma,s\in\Lambda_{v}\},$
where vertices $(v,s)$ and $(w,t)$ share an edge in $\Lambda_{\Gamma}$ if either $v=w$ and $t,s$ share an edge in $\Lambda_{v}$ or $v\not=w$ and $v,w$ share an edge in $\Gamma$.
We observe that $\Lambda_{\Gamma}$ contains the graphs $\Lambda_{v}$ for $v\in\Gamma$ as (mutually disjoint) subgraphs. Furthermore, we observe that if we take $|\Lambda_{v}|=1$ for each $v\in\Gamma$ then $\Lambda_{\Gamma}=\Gamma$.
Remark 3.4.
For a simple graph $\Gamma$ and graphs $\{\Lambda_{v}\}_{v\in\Gamma}$, and groups $G_{w}$ and von Neumann algebras $(N_{w},\varphi_{w})$ with normal GNS-faithful state, with $w\in\Lambda_{v}$, we have
(3.2) $*_{w,\Lambda_{\Gamma}}G_{w}$ $=*_{v,\Gamma}(*_{w,\Lambda_{v}}G_{w})$
(3.3) $*_{w,\Lambda_{\Gamma}}(N_{w},\varphi_{w}),$ $=*_{v,\Gamma}(*_{w,\Lambda_{v}}(N_{w},\varphi_{w})).$
Indeed, this follows by the defining universal property of graph products of groups as well as its analogue for operator algebras that can be found in [14, Proposition 3.22].
Lemma 3.5.
Let $\Gamma$ be a simple graph and for each $v\in\Gamma$ let $\Lambda_{v}$ be a non-empty graph. Then the graph product graph $\Lambda_{\Gamma}$ is rigid if and only if for each vertex $v\in\Gamma$ the graph $\Lambda_{v}$ is rigid and the vertex $v$ satisfies at least one of the following conditions:
  1. (1)
    $\textrm{Link}_{\Gamma}(\textrm{Link}_{\Gamma}(v))=\{v\};$
  2. (2)
    $|\Lambda_{v}|\geq 2$.
Proof.
We may assume $\Gamma$ is non-empty. First, suppose the conditions in the lemma are satisfied. We show $\Lambda_{\Gamma}$ is rigid. Let $(v,j)\in\Lambda_{\Gamma}$ for some $v\in\Gamma$, $j\in\Lambda_{v}$. Let $(z,k)\in\textrm{Link}_{\Lambda_{\Gamma}}(\textrm{Link}_{\Lambda_{\Gamma}}(v,j))$. We need to show that $(z,k)=(v,j)$.
Suppose first that $|\Lambda_{v}|\geq 2$. Then, as $\Lambda_{v}$ is rigid, we have that $\textrm{Link}_{\Lambda_{v}}(j)$ is non-empty. Let $l\in\textrm{Link}_{\Lambda_{v}}(j)$. Then $(v,l)\in\textrm{Link}_{\Lambda_{\Gamma}}(v,j)$ and similarly $(z,k)\in\textrm{Link}_{\Lambda_{\Gamma}}(v,l)$. If $z\not=v$ then by the definition of the graph product graph this implies $z\in\textrm{Link}_{\Gamma}(v)$. But then, again by the definition of the graph product graph, we obtain $(z,k)\in\textrm{Link}_{\Lambda_{\Gamma}}(v,j)$. However, as $(z,k)\not\in\textrm{Link}_{\Lambda_{\Gamma}}(z,k)$, this contradicts that $(z,k)\in\textrm{Link}_{\Lambda_{\Gamma}}(\textrm{Link}_{\Lambda_{\Gamma}}(v,j))$. We conclude that $z=v$. Hence, since $(z,k)\in\textrm{Link}_{\Lambda_{\Gamma}}(v,l)$ we obtain that $k\in\textrm{Link}_{\Lambda_{v}}(l)$. Since this holds true for all $l\in\textrm{Link}_{\Lambda_{v}}(j)$, we obtain that $k\in\textrm{Link}_{\Lambda_{v}}(\textrm{Link}_{\Lambda_{v}}(j))$, so that $k=j$ by rigidity of $\Lambda_{v}$. Thus $(z,k)=(v,j)$.
Now suppose that $|\Lambda_{v}|<2$, i.e. $\Lambda_{v}=\{j\}$, and just assume that $\textrm{Link}_{\Gamma}(\textrm{Link}_{\Gamma}(v))=\{v\}$. If $|\Gamma|=1$ then $\Lambda_{\Gamma}=\Lambda_{v}$ is rigid. Thus we can assume $|\Gamma|\geq 2$. Then $\textrm{Link}_{\Gamma}(v)$ must be non-empty since $\textrm{Link}(\varnothing)=\Gamma\not=\{v\}$. Take $w\in\textrm{Link}_{\Gamma}(v)$. Then, as by assumption $\Lambda_{w}$ is non-empty, we can pick $i\in\Lambda_{w}$. Now $(w,i)\in\textrm{Link}_{\Lambda_{\Gamma}}(v,j)$, by the definition of the graph $\Lambda_{\Gamma}$. Thus $(z,k)\in\textrm{Link}_{\Lambda_{\Gamma}}(w,i)$. If $w=z$ then $z\in\textrm{Link}_{\Gamma}(v)$ and so also $(z,k)\in\textrm{Link}_{\Lambda_{\Gamma}}(v,j)$. But as $(z,k)\not\in\textrm{Link}_{\Lambda_{\Gamma}}(z,k)$, this contradicts that $(z,k)\in\textrm{Link}_{\Lambda_{\Gamma}}(\textrm{Link}_{\Lambda_{\Gamma}}(v,j))$. Thus $w\not=z$, and therefore, as $(z,k)\in\textrm{Link}_{\Lambda_{\Gamma}}(w,i)$, we obtain that $z\in\textrm{Link}_{\Gamma}(w)$. Therefore, since this holds for any $w\in\textrm{Link}_{\Gamma}(v)$, we obtain that $z\in\textrm{Link}_{\Gamma}(\textrm{Link}_{\Gamma}(v))=\{v\}$ and thus $z=v$. Thus as $k\in\Lambda_{z}=\Lambda_{v}=\{j\}$, we obtain $(z,k)=(v,j)$.
We now prove the reverse direction. First, suppose there is a vertex $v\in\Gamma$ such that $\Lambda_{v}$ is not rigid. Take $j\in\Lambda_{v}$ such that $\textrm{Link}_{\Lambda_{v}}(\textrm{Link}_{\Lambda_{v}}(j))\not=\{j\}$ so that we can choose $k\in\textrm{Link}_{\Lambda_{v}}(\textrm{Link}_{\Lambda_{v}}(j))$ with $k\not=j$. Now, one can check that $(v,k)\in\textrm{Link}_{\Lambda_{\Gamma}}(\textrm{Link}_{\Lambda_{\Gamma}}(v,j))$, hence $\Lambda_{\Gamma}$ is not rigid.
Now suppose there is vertex $v\in\Gamma$ such that $\textrm{Link}_{\Gamma}(\textrm{Link}_{\Gamma}(v))\not=\{v\}$ and $|\Lambda_{v}|=1$, i.e. $\Lambda_{v}=\{j\}$ for some $j$. Then $\textrm{Link}_{\Lambda_{\Gamma}}(v,j)=\bigcup_{w\in\textrm{Link}_{\Gamma}(v)}\{(w,i):i\in\Lambda_{w}\}$. We can choose a $z\in\textrm{Link}_{\Gamma}(\textrm{Link}_{\Gamma}(v))$ with $z\not=v$ and let $k\in\Lambda_{z}$. Then we see that $(z,k)\in\textrm{Link}_{\Lambda_{\Gamma}}(\textrm{Link}_{\Lambda_{\Gamma}}(v,j))$, which shows $\Lambda_{\Gamma}$ is not rigid. ∎
By the result of Lemma 3.5, it is possible to construct many different rigid graphs using the rigid graphs from Example 3.2.
Remark 3.6.
Let $\Gamma$ be a rigid graph. Then any connected component of $\Gamma$ is rigid and any irreducible component of $\Gamma$ is rigid. Indeed, if $\Lambda_{1},\ldots,\Lambda_{n}$ are the irreducible components of $\Gamma$ and we let $\Pi=\{1,\ldots,n\}$ be a complete graph, then $\Gamma=*_{v,\Pi}\Lambda_{v}=\Lambda_{\Pi}$. Hence, by Lemma 3.5 and rigidity of $\Gamma$ we obtain that the graphs $\Lambda_{1},\ldots,\Lambda_{n}$ are rigid. Similarly, if we let $\Lambda_{1}^{\prime},\ldots,\Lambda_{m}^{\prime}$ be connected components of $\Gamma$ and we let $\Pi^{\prime}=\{1,\ldots,m\}$ be a graph with no edges, then $\Gamma=*_{v,\Pi^{\prime}}\Lambda_{v}^{\prime}=\Lambda_{\Pi^{\prime}}^{\prime}$ so that by Lemma 3.5 and rigidity of $\Gamma$ we obtain that $\Lambda_{1}^{\prime},\ldots,\Lambda_{m}^{\prime}$ are rigid.
We now define the core of a graph.
Definition 3.7 (Core of a graph).
Let $\Gamma$ be a simple graph. We say that two vertices $v,w\in\Gamma$ are core equivalent, with notation $v\sim w$, if $\textrm{Star}(v)=\textrm{Star}(w)$. Let $\overline{v}$ be the core equivalence class of $v\in\Gamma$. We define the core of $\Gamma$, with notation $\mathcal{C}\Gamma$, as the graph whose vertices set is the set of all core equivalence classes of $\Gamma$. The edges set of $\mathcal{C}\Gamma$ is defined by declaring that $\overline{v},\overline{w}\in\mathcal{C}\Gamma$ with $\overline{v}\not=\overline{w}$ share an edge in $\mathcal{C}\Gamma$ if and only if $v,w$ share an edge in $\Gamma$.
We remark that $\mathcal{C}\mathcal{C}\Gamma=\mathcal{C}\Gamma$, that is, the core of the core of a graph is equal to the core of the graph. In the following lemma we show that any graph can be written as a graph product over its core.
Lemma 3.8.
Let $\Gamma$ be a simple graph. For $\overline{v}\in\mathcal{C}\Gamma$ let $\Lambda_{\overline{v}}$ be the complete graph of size $|\Lambda_{\overline{v}}|=|\overline{v}|$. Then $\Gamma\simeq\Lambda_{\mathcal{C}\Gamma}$. Furthermore, if $\mathcal{C}\Gamma$ is rigid, then so is $\Gamma$.
Proof.
Indeed, as for $\overline{v}\in\mathcal{C}\Gamma$ we have $|\overline{v}|=|\Lambda_{\overline{v}}|$, we can build a bijection $\iota_{\overline{v}}:\overline{v}\to\Lambda_{\overline{v}}$. We then define the bijection $\iota:\Gamma\to\Lambda_{\mathcal{C}\Gamma}$ as $\iota(v)=(\overline{v},\iota_{\overline{v}}(v))$. We show this is a graph isomorphism. Let $v\not=w\in\Gamma$. If $v,w$ do not share an edge in $\Gamma$ then $\overline{v}\not=\overline{w}$ and $\overline{v},\overline{w}$ do not share an edge in $\mathcal{C}\Gamma$. Hence $(\overline{v},\iota_{\overline{v}}(v))$ and $(\overline{w},\iota_{\overline{w}}(w))$ do not share an edge in $\Lambda_{\mathcal{C}\Gamma}$. Now suppose $v,w$ do share an edge in $\Gamma$. If $\overline{v}=\overline{w}$ then since $\Lambda_{\overline{v}}=\Lambda_{\overline{w}}$ is complete we obtain that $(\overline{v},\iota_{\overline{v}}(v))$ and $(\overline{w},\iota_{\overline{w}}(w))$ share an edge in $\Lambda_{\mathcal{C}\Gamma}$ . On the other hand, if $\overline{v}\not=\overline{w}$, then $\overline{v},\overline{w}$ share an edge in $\mathcal{C}\Gamma$ so that also $(\overline{v},\iota_{\overline{v}}(v))$ and $(\overline{w},\iota_{\overline{w}}(w))$ share an edge in $\Lambda_{\mathcal{C}\Gamma}$. This shows that $\iota$ is an isomorphism and hence $\Gamma\simeq\Lambda_{\mathcal{C}\Gamma}$.
We prove the last statement. Suppose $\mathcal{C}\Gamma$ is rigid. Since for each $\overline{v}\in\mathcal{C}\Gamma$ the graph $\Lambda_{\overline{v}}$ is rigid (since it is complete) and since by rigidity of $\mathcal{C}\Gamma$ we have $\textrm{Link}_{\mathcal{C}\Gamma}(\textrm{Link}_{\mathcal{C}\Gamma}(\overline{v}))=\overline{v}$, we obtain by Lemma 3.5 that $\Lambda_{\mathcal{C}\Gamma}$ is rigid. Thus $\Gamma\simeq\Lambda_{\mathcal{C}\Gamma}$ is rigid. ∎
We make two remarks on Lemma 3.8
Remark 3.9.
We remark that if a simple graph $\Gamma$ is rigid, then its core is, in general, not rigid. Indeed, let $\Pi=\{v,w\}$ denote the simple graph of size $2$ with no edges and let $\Lambda_{v},\Lambda_{w}$ denote complete graphs of size $|\Lambda_{v}|,|\Lambda_{w}|\geq 2$. Then the graph $\Gamma:=\Lambda_{\Pi}$ is rigid by Lemma 3.5 but $\mathcal{C}\Gamma=\Pi$ is not rigid.
Remark 3.10.
If a graph $\Gamma$ is in the class CC1 as described in [17] then $\Gamma$ is rigid. Indeed if $\Gamma$ is CC1 then its core $\mathcal{C}\Gamma$, which is in fact also CC1, is given by the graph of [17, Eqn. (1.1)]. This graph is rigid as can be checked directly from the very definition of rigidity. We can then apply Lemma 3.8 to obtain that $\Gamma$ is rigid. It thus follows that the graphs considered in the current paper form a much richer class than [17].
4. Graph products of nuclear C-algebras
The aim of this section is to give a sufficient condition for when the reduced graph product of nuclear C-algebras is nuclear again. Such a result cannot hold in full generality as it is clear from the fact that the free product of amenable discrete groups is non-amenable as soon as one group has at least 2 elements and the other group has at least 3 elements. Hence the stability result in this section requires particular conditions on the states with respect to which we take the graph product. Such a result was obtained by Ozawa in [53] for amalgamated free products and we use the amalgamated free product decomposition of graph products (2.1) to show that the same holds for graph products.
Let $\Gamma$ be a simple graph. Let $(A_{v},\varphi_{v})$ with $v\in\Gamma$ be unital C-algebras $A_{v}$, GNS-faithful states $\varphi_{v}$ and GNS-representation $\pi_{v}$ of $A_{v}$ on the Hilbert space $\mathcal{H}_{v}=L^{2}(A_{v},\varphi_{v})$.
For Hilbert C-modules we refer to [50]. Consider the reduced graph product C-algebras $(A_{\Lambda},\varphi_{\Lambda})$ for any $\Lambda\subseteq\Gamma$ which is a subalgebra of $(A_{\Gamma},\varphi_{\Gamma})$ with conditional expectation $\mathbb{E}_{\Lambda}$.
Definition 4.1.
We construct a Hilbert C-module $\mathcal{H}_{\mathbb{E}_{\Lambda}}$ as the completion of $A_{\Gamma}$ with respect to the $A_{\Lambda}$-valued inner product
$\langle a,b\rangle_{\mathbb{E}_{\Lambda}}=\mathbb{E}_{\Lambda}(b^{\ast}a)$
and the corresponding Hilbert $A_{\Lambda}$-module norm $\|a\|=\|\langle a,a\rangle^{\frac{1}{2}}\|$. Let $\pi_{\mathbb{E}_{\Lambda}}:A_{\Gamma}\rightarrow\mathbb{B}(\mathcal{H}_{\mathbb{E}_{\Lambda}})$ be the GNS-representation of $A_{\Gamma}$ on the Hilbert C-module $\mathcal{H}_{\mathbb{E}_{\Lambda}}$ by adjointable operators. Then $\pi_{\mathbb{E}_{\Lambda}}$ is given by extending left multiplication
$\pi_{\mathbb{E}_{\Lambda}}(x)a=xa,x\in A_{\Gamma},a\in A_{\Gamma}\subseteq\mathcal{H}_{\mathbb{E}_{\Lambda}}$
and we shall omit $\pi_{\mathbb{E}_{\Lambda}}$ in the notation if the module action is clear.
Definition 4.2.
An operator on the Hilbert $A_{\Lambda}$-module $\mathcal{H}_{\mathbb{E}_{\Lambda}}$ is called finite rank if it is in the linear span of operators of the form
$\theta_{\eta_{2},\eta_{1}}:\xi\mapsto\eta_{2}\langle\xi,\eta_{1}\rangle_{\mathbb{E}_{\Lambda}},\qquad\eta_{i}\in\mathcal{H}_{\mathbb{E}_{\Lambda}}.$
The closure of the space of all finite rank operators are defined as the space of compact operators $\mathbb{K}(\mathcal{H}_{\mathbb{E}_{\Lambda}})$.
Lemma 4.3.
Suppose there exists $v\in\Gamma$ such that $\Gamma=\textrm{Star}(v)$. If $\pi_{v}(A_{v})$ contains $\mathbb{K}(\mathcal{H}_{v})$ then $\pi_{\mathbb{E}_{\textrm{Link}(v)}}(A_{\textrm{Star}(v)})$ contains $\mathbb{K}(\mathcal{H}_{\mathbb{E}_{\textrm{Link}(v)}})$.
Proof.
We have that $A_{\textrm{Star}(v)}=A_{v}\otimes A_{\textrm{Link}(v)}$ where the tensor product is the minimal tensor product and under this correspondence we have
$\langle a\otimes b,c\otimes d\rangle_{\mathbb{E}_{\textrm{Link}(v)}}=\varphi_{v}(c^{\ast}a)d^{\ast}b,\qquad a,c\in A_{v},b,d\in A_{\textrm{Link}(v)}.$
We thus may identify $\mathcal{H}_{\mathbb{E}_{\textrm{Link}(v)}}$ as the closure of the algebraic tensor product $\mathcal{H}_{v}\otimes A_{\textrm{Link}(v)}$ with respect to the inner product $\langle\xi\otimes b,\eta\otimes d\rangle=\langle\xi,\eta\rangle d^{\ast}b$. Further, under this correspondence $\pi_{\mathbb{E}_{\textrm{Link}(v)}}=\pi_{v}\otimes\pi_{l}$ where $\pi_{l}(x)a=xa,x,a\in A_{\textrm{Link}(v)}$ is the left multiplication. Let $p_{v}$ be the projection of $\mathcal{H}_{v}$ onto $\mathbb{C}\xi_{v}$. Then $p_{v}\otimes 1$ equals the extension of $\mathbb{E}_{\textrm{Link}(v)}$ as a bounded map on $\mathcal{H}_{\mathbb{E}_{\textrm{Link}(v)}}$ identified with the closure of $\mathcal{H}_{v}\otimes A_{\textrm{Link}(v)}$. As by assumption $p_{v}$ lies in $\pi_{v}(A_{v})$ it thus follows that $p_{v}\otimes 1$ lies in $\pi_{\mathbb{E}_{\textrm{Link}(v)}}(A_{\textrm{Star}(v)})$. It thus follows that for $a,c,x\in A_{v},b,d,y\in A_{\textrm{Link}(v)}$ we have
$\theta_{a\otimes b,c\otimes d}(x\otimes y)=\varphi_{v}(c^{\ast}x)a\otimes bd^{\ast}y=\pi_{\mathbb{E}_{\textrm{Link}(v)}}(a\otimes b)(p_{v}\otimes 1)\pi_{\mathbb{E}_{\textrm{Link}(v)}}(c^{\ast}\otimes d^{\ast})(x\otimes y).$
The right hand side is contained in $\pi_{\mathbb{E}_{\textrm{Link}(v)}}(A_{\textrm{Star}(v)})$. Hence $\pi_{\mathbb{E}_{\textrm{Link}(v)}}(A_{\textrm{Star}(v)})$ contains a dense set of finite rank operators and hence must contain all compact operators. ∎
Theorem 4.4.
Let $\Gamma$ be a simple graph. If for each $v\in\Gamma$, $A_{v}$ is nuclear and $\pi_{v}(A_{v})$ contains the compact operators $\mathbb{K}(\mathcal{H}_{v})$, then $A_{\Gamma}$ is nuclear.
Proof.
It suffices to prove the theorem for $\Gamma$ a finite graph as inductive limits of inclusions of nuclear C-algebras are nuclear.
Our proof proceeds by induction to the number of vertices in $\Gamma$. So we assume that for any $\Lambda\subsetneq\Gamma$ we have proved that $A_{\Lambda}$ is nuclear. We shall prove that $A_{\Gamma}$ is nuclear.
If $\Gamma$ is complete then $A_{\Gamma}$ is the minimal tensor product of $A_{v},v\in\Gamma$ which is nuclear as each $A_{v}$ is nuclear.
Assume $\Gamma$ is not complete. Then we may take $v\in\Gamma$ such that $\textrm{Star}(v)\not=\Gamma$. In this case
[…]
5. Relative amenability, quasi-normalizers and embeddings in graph products
In this section we establish the required machinery we need throughout the paper. First in Section 5.1 we discuss how to calculate conditional expectations in graph products. This will be used in Section 5.2 to prove a result concerning relative amenability in graph products. The calculations from Section 5.1 will furthermore be used in Section 5.3 to keep control of certain quasi-normalizers in graph products. Last, in Section 5.4 we apply results from Section 5.3 to establish a unitary embedding of certain subalgebras in graph products.
5.1. Calculating conditional expectations in graph products
For a simple graph $\Gamma$, a graph product $(M_{\Gamma},\varphi_{\Gamma})=*_{v,\Gamma}(M_{v},\varphi_{v})$ and subgraphs $\Gamma_{1},\Gamma_{2}\subseteq\Gamma$ we discuss how to calculate iterated conditional expectations of the form $\mathbb{E}_{M_{\Gamma_{2}}}(a\mathbb{E}_{M_{\Gamma_{1}}}(x)b)$ for $a,b,x\in M_{\Gamma}$. Such calculations have been done in [6] in the setting of right-angled Coxeter groups (i.e. the setting $M_{v}=\mathcal{L}(\mathbb{Z}/2\mathbb{Z})$ for all $v\in\Gamma$) and in [16] for general graph products. We state Proposition 5.1 which largely follows from [16, Lemma 3.17] and also note that, only in this section, $\mathbb{E}_{M_{\Lambda}}$ denotes the conditional expectation that preserves the state $\varphi_{\Gamma}$.
Proposition 5.1.
Let $\Gamma$ be a simple graph and let $\Gamma_{1},\Gamma_{2}$ be its subgraphs. Let $\mathbf{u},\mathbf{v}\in\mathcal{W}_{\Gamma}$ and write $\mathbf{u}=\mathbf{u}_{l}\mathbf{u}_{c}\mathbf{u}_{r}$ and $\mathbf{v}=\mathbf{v}_{l}\mathbf{v}_{c}\mathbf{v}_{r}$ (both reduced) with $\mathbf{u}_{l},\mathbf{v}_{l}\in\mathcal{W}_{\Gamma_{1}}$, $\mathbf{u}_{r},\mathbf{v}_{r}\in\mathcal{W}_{\Gamma_{2}}$ and such that $\mathbf{u}_{c},\mathbf{v}_{c}$ do not start with letters from $\Gamma_{1}$ and do not end with letters from $\Gamma_{2}$.
For $v\in\Gamma$ let $(M_{v},\varphi_{v})$ be a von Neumann algebra with a normal faithful state. Let $a=a_{l}a_{c}a_{r}$ and $b=b_{l}b_{c}b_{r}$ where $a_{l}\in\mathring{M}_{\mathbf{u}_{l}}$, $a_{c}\in\mathring{M}_{\mathbf{u}_{c}}$, $a_{r}\in\mathring{M}_{\mathbf{u}_{r}}$ and $b_{l}\in\mathring{M}_{\mathbf{v}_{l}}$, $b_{c}\in\mathring{M}_{\mathbf{v}_{c}}$, $b_{r}\in\mathring{M}_{\mathbf{v}_{r}}$. Then for $x\in M_{\Gamma}$ we have
$\mathbb{E}_{M_{\Gamma_{2}}}(a^{*}\mathbb{E}_{M_{\Gamma_{1}}}(x)b)=\varphi(a_{c}^{*}b_{c})a_{r}^{*}\mathbb{E}_{M_{\Gamma_{1}\cap\Gamma_{2}\cap\textrm{Link}(\mathbf{u}_{c})}}(a_{l}^{*}xb_{l})b_{r}.$
Proof.
As $a_{l}^{*},b_{l}\in M_{\Gamma_{1}}$ and $a_{r}^{*},b_{r}\in M_{\Gamma_{2}}$ we have
(5.1) $\mathbb{E}_{M_{\Gamma_{2}}}(a^{*}\mathbb{E}_{M_{\Gamma_{1}}}(x)b)$ $=a_{r}^{*}\mathbb{E}_{M_{\Gamma_{2}}}(a_{c}^{*}\mathbb{E}_{M_{\Gamma_{1}}}(a_{l}^{*}xb_{l})b_{c})b_{r}.$
Now as $\mathbb{E}_{M_{\Gamma_{1}}}(a_{l}^{*}xb_{l})\in M_{\Gamma_{1}}$ and $a_{c}^{*}\in\mathring{M}_{\mathbf{u}_{c}^{-1}}$ and $b_{c}\in\mathring{M}_{\mathbf{v}_{c}}$ it follows from [16, Lemma 3.17] and linearity that
(5.2) $\begin{split}\mathbb{E}_{M_{\Gamma_{2}}}(a_{c}^{*}\mathbb{E}_{M_{\Gamma_{1}}}(a_{l}^{*}xb_{l})b_{c})&=\varphi(a_{c}^{*}b_{c})\mathbb{E}_{M_{\Gamma_{1}\cap\Gamma_{2}\cap\textrm{Link}(\mathbf{u}_{c})}}(\mathbb{E}_{M_{\Gamma_{1}}}(a_{l}^{*}xb_{l}))\\ &=\varphi(a_{c}^{*}b_{c})\mathbb{E}_{M_{\Gamma_{1}\cap\Gamma_{2}\cap\textrm{Link}(\mathbf{u}_{c})}}(a_{l}^{*}xb_{l}).\end{split}$
This proves the statement by combining (5.1) and (5.2).
5.2. Relative amenability in graph products
We state the definition of relative amenability for which we refer to [63, Proposition 2.4].
Definition 5.2.
Let $(M,\tau)$ be a tracial von Neumann algebra and let $P\subseteq 1_{P}M1_{P},Q\subseteq M$ be von Neumann subalgebras. We say that $P$ is amenable relative to $Q$ inside $M$ if there exists a $P$-central positive functional on $1_{P}\langle M,e_{Q}\rangle 1_{P}$ that restricts to the trace $\tau$ on $1_{P}M1_{P}$.
Using the calculations of conditional expectations we will prove Theorem 5.3 which asserts that when a von Neumann algebra $P\subseteq M_{\Gamma}$ is amenable relative to $M_{\Gamma_{i}}$ inside $M_{\Gamma}$ for some subgraphs $\Gamma_{i}\subseteq\Gamma$ for $i=1,2$, then $P$ is also amenable relative to $M_{\Gamma_{1}\cap\Gamma_{2}}$ inside $M_{\Gamma}$. Theorem 5.3 generalizes [6, Theorem 3.7] where the statement was proven in the setting of right-angled Coxeter groups. The proof of Theorem 5.3 is more or less identical to [6, Theorem 3.7]. In the proof of Theorem 5.3 we will therefore refer to the proof and notation from [6, Section 3, Theorem 3.7]. We remark that bimodule computations we do in the proof of Theorem 5.3 are also related to those done in [16, Section 5].
Theorem 5.3.
Let $\Gamma$ be a simple graph and let $\Gamma_{1},\Gamma_{2}\subseteq\Gamma$ be subgraphs. For $v\in\Gamma$ let $(M_{v},\tau_{v})$ be a von Neumann algebra with a normal faithful trace. Let $P\subset 1_{P}M_{\Gamma}1_{P}$ be a von Neumann subalgebra that is amenable relative to $M_{\Gamma_{i}}$ inside $M_{\Gamma}$ for $i=1,2$. Then $P$ is amenable relative to $M_{\Gamma_{1}\cap\Gamma_{2}}$ inside $M_{\Gamma}$.
Proof.
By [6, Remark 2.3] we may assume that the inclusion $P\subseteq M_{\Gamma}$ is unital. As in [6, Theorem 3.7] let $Q_{i}:=M_{\Gamma_{i}}$ for $i=1,2$, put $Q_{0}=Q_{1}\cap Q_{2}$ and put the Connes relative tensor product bimodule
$\mathcal{H}=L^{2}(M_{\Gamma},e_{Q_{1}})\otimes_{M_{\Gamma}}L^{2}(M_{\Gamma},e_{Q_{2}}).$
Similar to the proof of [6, Theorem 3.7] we obtain that the bimodule $_{M_{\Gamma}}L^{2}(M_{\Gamma})_{P}$ is weakly contained in $_{M_{\Gamma}}\mathcal{H}_{P}$. Denote
[…]
and define the subspace
$\mathcal{H}_{0}=\operatorname{Span}\{xe_{Q_{1}}y\otimes_{M_{\Gamma}}e_{Q_{2}}z:\mathbf{v}\in\mathcal{V},x,z\in M_{\Gamma},y\in\mathring{M_{\mathbf{v}}}\},$
which is dense in $\mathcal{H}$ (see [6, Theorem 3.7]). We define a bimodule $\mathcal{K}:=L^{2}(M_{\Gamma})\otimes_{Q_{0}}L^{2}(M_{\Gamma})$ and a map $U:\mathcal{H}_{0}\to L^{2}(M_{\Gamma})\otimes_{Q_{0}}\mathcal{K}$ as
$xe_{Q_{1}}y\otimes_{M_{\Gamma}}e_{Q_{2}}z\mapsto x\otimes_{Q_{0}}y\otimes_{Q_{0}}z\quad x,z\in M_{\Gamma},\mathbf{v}\in\mathcal{V},y\in\mathring{M}_{\mathbf{v}}.$
Let $x,x^{\prime},z,z^{\prime}\in M_{\Gamma}$, $\mathbf{v}\in\mathcal{V}$ and $y,y^{\prime}\in\mathring{M}_{\mathbf{v}}$. By Proposition 5.1 we have
(5.3) $\mathbb{E}_{Q_{0}}(y^{*}\mathbb{E}_{Q_{0}}(x^{*}x^{\prime})y^{\prime})=\tau(y^{*}y^{\prime})\mathbb{E}_{Q_{0}\cap\textrm{Link}(\mathbf{v})}(x^{*}x^{\prime})=\mathbb{E}_{Q_{1}}(y^{*}\mathbb{E}_{Q_{2}}(x^{*}x^{\prime})y^{\prime}).$
Similarly to [6, Theorem 3.7] we can use this to show that
(5.4) $\langle x^{\prime}\otimes_{Q_{0}}y^{\prime}\otimes_{Q_{0}}z^{\prime},x\otimes_{Q_{0}}y\otimes_{Q_{0}}z\rangle=\langle x^{\prime}e_{Q_{1}}y^{\prime}\otimes_{M_{\Gamma}}e_{Q_{2}}z^{\prime},xe_{Q_{1}}y\otimes_{M_{\Gamma}}e_{Q_{2}}z\rangle.$
As in [6, Theorem 3.7] we conclude $P$ is amenable relative to $Q_{0}=Q_{1}\cap Q_{2}$ inside $M_{\Gamma}$. ∎
5.3. Embeddings of quasi-normalizers in graph products
We prove Proposition 5.8 and Proposition 5.9 concerning embeddings in graph products. To prove Proposition 5.8 we need some auxiliary lemmas. We state Lemma 2.1 from [6], which was essentially proven in [69, Remark 3.8]. The result is surely known but for completeness we give the proof.
Lemma 5.4 (Lemma 2.1 in [6]).
Let $A,\{B_{k}\}_{k\in I},Q\subseteq M$ be von Neumann subalgebras with $B_{k}\subseteq Q$ for every $k\in I$ where $I$ is some index set. Assume that $A\prec_{M}Q$ but $A\not\prec_{M}B_{k}$ for any $k\in I$. Then there exist projections $p\in A,q\in Q$, a non-zero partial isometry $v\in qMp$ and a normal $\ast$-homomorphism $\theta:pAp\rightarrow qQq$ such that $\theta(x)v=vx,x\in pAp$ and such that $\theta(pAp)\not\prec_{Q}B_{k}$ for any $k\in I$. Moreover, it may be assumed that $p$ is majorized by the support of $\mathbb{E}_{A}(v^{*}v)$.
Proof.
Let $p\in A,q\in Q$ and $\theta:pAp\rightarrow qQq$ be a normal $\ast$-homomorphism such that there is a partial isometry $v\in qMp$ such that $\theta(x)v=vx$ for all $x\in pAp$. We first prove that without loss of generality we can assume that $p$ is majorized by the support of $\mathbb{E}_{A}(v^{*}v)$.
Let $z$ be the support of $\mathbb{E}_{A}(v^{*}v)$. As $p\mathbb{E}_{A}(v^{*}v)p=\mathbb{E}_{A}(pv^{*}vp)=\mathbb{E}_{A}(v^{*}v)$ it follows that $z\in pAp$. Further for $x\in pAp$ we have $x\mathbb{E}_{A}(v^{*}v)=\mathbb{E}_{A}(xv^{*}v)=\mathbb{E}_{A}(v^{*}\theta(x)v)=\mathbb{E}_{A}(v^{*}vx)=\mathbb{E}_{A}(v^{*}v)x$ so that $z\in(pAp)^{\prime}$. We conclude $z\in(pAp)^{\prime}\cap pAp$. Now let $p^{\prime}:=pz\in A$, let $\theta^{\prime}:p^{\prime}Ap^{\prime}\rightarrow qQq$ be the restriction of $\theta$ to $p^{\prime}Ap^{\prime}$ and let $v^{\prime}:=vz\in qMp^{\prime}$. Then for $x\in p^{\prime}Ap^{\prime}$ we have $\theta^{\prime}(x)v^{\prime}=\theta(x)vz=vxz=vzx$. We claim further that $v^{\prime}$ is non-zero. Indeed, $v^{\prime}=vz=0$ iff $zv^{\ast}vz=0$ iff $0=\mathbb{E}_{A}(zv^{\ast}vz)=z\mathbb{E}_{A}(v^{\ast}v)z$. But as $v$ is non-zero $\mathbb{E}_{A}(v^{\ast}v)$ is non-zero and hence $z\mathbb{E}_{A}(v^{\ast}v)z\not=0$ by construction of $z$. We conclude that $v^{\prime}\not=0$. In all the tuple $(\theta,p^{\prime},q,v^{\prime})$ witnesses that $A\prec_{M}Q$ and the support of $\mathbb{E}_{A}((v^{\prime})^{\ast}v^{\prime})$ majorizes $p^{\prime}$.
For the remainder of the proof one just follows [6, Lemma 2.1] which does not affect the assumption that $p$ is majorized by the support of $\mathbb{E}_{A}(v^{*}v)$. Althought the statement of [6, Lemma 2.1] is for $I$ with finite cardinality, the same proof can also be applied to $I$ with infinite cardinality. ∎
The following lemma is similar to [30, Remark 2.3].
Lemma 5.5.
Let $(M,\tau)$ be a tracial von Neumann algebra and let $A,\{B_{k}\}_{k\in I}$ be (possibly non-unital) von Neumann subalgebras of $M$. Here $I$ is an index set. Assume $A\not\prec_{M}B_{k}$ for every $k\in I$. Then there is a single net $(u_{i})_{i}$ of unitaries in $A$ such that for $k\in I$ and $a,b\in 1_{A}M1_{B_{k}}$ we have $\|\mathbb{E}_{B_{k}}(a^{*}u_{i}b)\|_{2}\to 0$ as $i\to\infty$
Proof.
Put
(5.5) $\widetilde{B}$ $=\bigoplus_{k\in I}B_{k},$ $\quad\widetilde{M}$ $=\bigoplus_{k\in I}M.$
Let $\pi:M\to\widetilde{M}$ be the (normal) diagonal embedding $\pi(x)=\bigoplus_{k\in I}x$. Suppose $\pi(A)\prec_{\widetilde{M}}\widetilde{B}$. Then there are projections $p\in\pi(A)$, $q\in\widetilde{B}$, a normal *-homomorphism $\theta:p\pi(A)p\to q\widetilde{B}q$ and a non-zero partial isometry $v\in q\widetilde{M}p$ s.t. $\theta(x)v=vx$ for $x\in p\pi(A)p$. For $k\in I$ let $\pi_{k}:\widetilde{M}\to M$ be the coordinate projections. Denote $p_{k}:=\pi_{k}(p)\in A$, $q_{k}:=\pi_{k}(q)\in B_{k}$ and $v_{k}:=\pi_{k}(v)\in\pi_{k}(q\widetilde{M}p)=q_{k}Mp_{k}$. Define a normal *-homomorphism $\theta_{k}:p_{k}Ap_{k}\to q_{k}Bq_{k}$ as $\theta_{k}(x)=\pi_{k}(\theta(\pi(x)))$. Then $\theta_{k}(x)v_{k}=\pi_{k}(\theta(\pi(x))v)=\pi_{k}(v\pi(x))=v_{k}\pi_{k}(\pi(x))=v_{k}x$. Since $0\not=v=\bigoplus_{k\in I}v_{k}$ there is $k_{0}\in I$ s.t. $v_{k_{0}}\not=0$. This then shows that $A\prec_{M}B_{k_{0}}$ which is a contradiction. We conclude that $\pi(A)\not\prec_{\widetilde{M}}\widetilde{B}$.
Thus, there is a net of unitaries $(\widetilde{u}_{i})_{i}$ in $\pi(A)$ s.t. for $a^{\prime},b^{\prime}\in 1_{\pi(A)}\widetilde{M}1_{\widetilde{B}}$ we have $\|\mathbb{E}_{\widetilde{B}}(a^{\prime*}\widetilde{u}_{i}b^{\prime})\|_{2}\to 0$ as $i\to\infty$. Let $u_{i}\in A$ be the unitary s.t. $\pi(u_{i})=\widetilde{u_{i}}$. Fix $1\leq k\leq n$ and let $a,b\in 1_{A}M1_{B_{k}}$. We can choose $\widetilde{a},\widetilde{b}\in 1_{\pi(A)}\widetilde{M}1_{\widetilde{B}}$ s.t. $\pi_{k}(\widetilde{a})=a$ and $\pi_{k}(\widetilde{b})=b$. We have
(5.6) $\|\mathbb{E}_{B_{k}}(a^{*}u_{i}b)\|_{2}=\|\mathbb{E}_{\pi_{k}(\widetilde{B})}(\pi_{k}(\widetilde{a}^{*}\widetilde{u}_{i}\widetilde{b}))\|_{2}\leq\|\mathbb{E}_{\widetilde{B}}(\widetilde{a}^{*}\widetilde{u}_{i}\widetilde{b})\|_{2}\to 0$  $\quad\text{ as }i\to\infty.$
This shows the net $(u_{i})_{i}$ satisfies the stated property. ∎
In order to have control over quasi-normalizers we need the following lemma. The lemma is stated in [72, Lemma D.3] for sequences, but holds equally well for nets.
Remark 5.6.
Consider an inclusion $B\subseteq 1_{B}M1_{B}$ of finite von Neumann algebras with conditional expectation $\mathbb{E}_{B}:1_{B}M1_{B}\rightarrow B$. We extend it to $\mathbb{E}_{B}:M\rightarrow B$ by setting $\mathbb{E}_{B}(x)=\mathbb{E}_{B}(1_{B}x1_{B})$. Fix a normal faithful tracial state $\tau$ on $M$. If $p\in B$ is a non-zero projection then $p$ is in the multiplicative domain of $\mathbb{E}_{B}$ and so $\mathbb{E}_{B}:pMp\rightarrow pBp$ is a conditional expectation. If $\mathbb{E}_{B}$ preserves $\tau$ then it also preserves the normal faithful tracial state $\tau(p)^{-1}\tau$ on $pMp$.
Lemma 5.7 (Lemma D.3 in [72]).
Let $(M,\tau)$ be a finite von Neumann algebra with normal faithful trace $\tau$ and let $B\subseteq 1_{B}M1_{B}$ and $A\subseteq 1_{A}B1_{A}$ be von Neumann subalgebras. Suppose there is a net of unitaries $(u_{i})_{i}$ in $A$ such that for all $a,b\in M$ with $\mathbb{E}_{B}(a)=\mathbb{E}_{B}(b)=0$ we have
(5.7) $\|\mathbb{E}_{B}(au_{i}b)\|_{2}\to 0\text{ as }i\to\infty.$
Then if $n\geq 1$, $x_{0},x_{1},\ldots,x_{n}\in M$ satisfy $Ax_{0}\subseteq\sum_{k=1}^{n}x_{k}B$ then we have that $1_{A}x_{0}1_{A}\in B$.
Proof.
We put $B_{0}=1_{A}B1_{A}$ and $M_{0}=1_{A}M1_{A}$ so that $A\subseteq B_{0}\subseteq M_{0}$ are unital inclusions. We observe $B_{0}=B\cap M_{0}$. Now let $a,b\in M_{0}$ be such that $\mathbb{E}_{B_{0}}(a)=\mathbb{E}_{B_{0}}(b)=0$. Then by Remark 5.6 with $p=1_{A}$ we find $\mathbb{E}_{B}(a)=\mathbb{E}_{B_{0}}(a)=0$ and similarly $\mathbb{E}_{B}(b)=0$. Thus by assumption $\|\mathbb{E}_{B}(au_{i}b)\|_{2}\to 0$ as $i\to\infty$. Hence, since $B_{0}=1_{A}B1_{A}$ we obtain $\|\mathbb{E}_{B_{0}}(au_{i}b)\|_{2}\to 0$ as $i\to\infty$. Choose a central projection $z\in B\cap B^{\prime}$ such that there exists $m\geq 1$ and partial isometries $v_{i}\in B$ for $1\leq i\leq m$ with $v_{i}v_{i}^{*}\leq 1_{A}$ and $\sum_{i=1}^{m}v_{i}^{*}v_{i}=z$. Now let $n\geq 1$, $x_{0},x_{1},\ldots,x_{n}\in M$ be such that $Ax_{0}\subseteq\sum_{k=1}^{n}x_{k}B$. Then
$A(1_{A}x_{0}z1_{A})=(Ax_{0}z)1_{A}\subseteq\sum_{k=1}^{n}x_{k}Bz1_{A}=\sum_{k=1}^{n}\sum_{i=1}^{m}x_{k}(v_{i}^{*}1_{A}v_{i})B1_{A}\subseteq\sum_{k=1}^{n}\sum_{i=1}^{m}x_{k}v_{i}^{*}B_{0}.$
Multiplying both sides from the left with $1_{A}$ gives $A(1_{A}x_{0}z1_{A})\subseteq\sum_{k=1}^{n}\sum_{i=1}^{m}1_{A}x_{k}v_{i}^{*}B_{0}$ where $1_{A}x_{k}v_{i}^{*}\in 1_{A}M1_{A}$. By the existence of the net $(u_{i})_{i}$ this implies, by applying [72, Lemma D.3] to the inclusions $A\subseteq B_{0}\subseteq M_{0}$, that $1_{A}x_{0}z1_{A}\in B$. As we may let $z$ approximate $1_{B}$ in the strong topology we find that $1_{A}x_{0}1_{A}\in B$. ∎
We are now able to show the following result which generalizes [6, Proposition 2.3] to general graph products. The second statement in the proposition should be compared to [43, Lemma 9.4]. While the inclusion $M_{\Lambda}\subseteq M_{\Gamma}$ is generally not mixing, we still have enough control over the (quasi)-normalizer of subalgebras. The proof of Proposition 5.8(1) uses Lemma 5.5, Lemma 5.7 and the results from Section 5.1 for calculating conditional expectations in graph products. The proof of Proposition 5.8(2) uses (1) and Lemma 5.4 which is analogues to [6], but we included it for convenience.
Proposition 5.8.
Let $\Gamma$ be a simple graph and for $v\in\Gamma$ let $(M_{v},\tau_{v})$ be a finite von Neumann algebra with normal faithful trace $\tau_{v}$. Let $\Lambda\subseteq\Gamma$ be a subgraph, and $\{\Lambda_{j}\}_{j\in\mathcal{J}}$ be a non-empty, collection of subgraphs of $\Gamma$. Define
(5.8) $\Lambda_{\operatorname{emb}}:=\Lambda\cup\bigcap_{j\in\mathcal{J}}\bigcup_{v\in\Lambda\setminus\Lambda_{j}}\textrm{Link}_{\Gamma}(v).$
Let $A\subseteq 1_{A}M_{\Gamma}1_{A}$ be a von Neumann subalgebra.
  1. (1)
    If $A\subseteq 1_{A}M_{\Lambda}1_{A}$ and $A\not\prec_{M_{\Gamma}}M_{\Lambda_{j}}$ for all $j\in\mathcal{J}$ then the following properties hold true:
    1. (a)
      There is a net $(u_{i})_{i}$ of unitaries in $A$ such that for all $a,b\in 1_{A}M_{\Gamma}1_{A}$ with $\mathbb{E}_{M_{\Lambda_{\operatorname{emb}}}}(a)=\mathbb{E}_{M_{\Lambda_{\operatorname{emb}}}}(b)=0$ we have $\|\mathbb{E}_{M_{\Lambda_{\operatorname{emb}}}}(au_{i}b)\|_{2}\rightarrow 0$;
    2. (b)
      $1_{A}\operatorname{qNor}_{M_{\Gamma}}(A)^{\prime\prime}1_{A}\subseteq M_{\Lambda_{\operatorname{emb}}}$;
    3. (c)
      For any unitary $u\in M_{\Gamma}$ satisfying $u^{*}Au\subseteq M_{\Lambda_{\operatorname{emb}}}$ we have $1_{A}u1_{A}\in M_{\Lambda_{\operatorname{emb}}}$.
  2. (2)
    Denote $P=\operatorname{Nor}_{M_{\Gamma}}(A)^{\prime\prime}$ and let $r\in P\cap P^{\prime}$ be a projection. If $rA\prec_{M_{\Gamma}}M_{\Lambda}$ and $rA\not\prec_{M_{\Gamma}}M_{\Lambda_{j}}$ for $j\in\mathcal{J}$ then $rP\prec_{M_{\Gamma}}M_{\Lambda_{\operatorname{emb}}}$.
We remark that if $\{\Lambda_{j}\}_{j\in\mathcal{J}}$ enumerates all strict subgraphs of $\Lambda$ then $\Lambda_{\operatorname{emb}}=\Lambda\cup\textrm{Link}_{\Gamma}(\Lambda)$.
Proof.
(1) By Lemma 5.5 we can build a net of unitaries $(u_{i})_{i}$ in $A$ such that for any $a,b\in M_{\Gamma}$ and any $j\in\mathcal{J}$ we have $\|\mathbb{E}_{M_{\Lambda_{j}}}(au_{i}b)\|_{2}\to 0$ when $i\to\infty$. We show the net $(u_{i})_{i}$ satisfies the properties of (1a). Let $b\in\mathring{M}_{\mathbf{v}}$ and $c\in\mathring{M}_{\mathbf{w}}$ for some $\mathbf{v},\mathbf{w}\in\mathcal{W}_{\Gamma}\setminus\mathcal{W}_{\Lambda_{\operatorname{emb}}}$. Write $\mathbf{v}=\mathbf{v}_{l}\mathbf{v}_{c}\mathbf{v}_{r}$ and $\mathbf{w}=\mathbf{w}_{l}\mathbf{w}_{c}\mathbf{w}_{r}$ with $\mathbf{v}_{l},\mathbf{w}_{l}\in\mathcal{W}_{\Lambda_{\operatorname{emb}}}$, $\mathbf{v}_{r},\mathbf{w}_{r}\in\mathcal{W}_{\Lambda}$ and such that $\mathbf{v}_{c}$ and $\mathbf{w}_{c}$ do not start with letter from $\Lambda_{\operatorname{emb}}$ nor do they end with letters from $\Lambda$. Now write $b=b_{l}b_{c}b_{r}$ and $c=c_{l}c_{c}c_{r}$ with $b_{l}\in\mathring{M}_{\mathbf{v}_{l}},c_{l}\in\mathring{M}_{\mathbf{w}_{l}}$, $b_{c}\in\mathring{M}_{\mathbf{v}_{c}},c_{c}\in\mathring{M}_{\mathbf{w}_{c}}$ and $b_{r}\in\mathring{M}_{\mathbf{v}_{r}},c_{r}\in\mathring{M}_{\mathbf{w}_{r}}$. Then as $\mathbf{v}\not\in\mathcal{W}_{\Lambda_{\operatorname{emb}}}$ and $\mathbf{v}_{l}\in\mathcal{W}_{\Lambda_{\operatorname{emb}}}$ and $\mathbf{v}_{r}\in\mathcal{W}_{\Lambda}\subseteq\mathcal{W}_{\Lambda_{\operatorname{emb}}}$, we have $\mathbf{v}_{c}\not\in\mathcal{W}_{\Lambda_{\operatorname{emb}}}$ and hence there is a letter $v$ of $\mathbf{v}_{c}$ such that $v\not\in\Lambda_{\operatorname{emb}}$. Thus, there is an index $j\in\mathcal{J}$ such that $v\not\in\bigcup_{w\in\Lambda\setminus\Lambda_{j}}\textrm{Link}_{\Gamma}(w)$. Hence $\textrm{Link}(v)\subseteq\Gamma\setminus(\Lambda\setminus\Lambda_{j})=\Lambda_{j}\cup(\Gamma\setminus\Lambda)$ and thus $\Lambda\cap\textrm{Link}(\mathbf{v}_{c})\subseteq\Lambda_{j}$. Using Proposition 5.1 we get,
$\begin{split}\|\mathbb{E}_{M_{\Lambda_{\operatorname{emb}}}}(b^{*}u_{i}c)\|_{2}=&\|\mathbb{E}_{M_{\Lambda_{\operatorname{emb}}}}(b^{*}\mathbb{E}_{M_{\Lambda}}(u_{i})c)\|_{2}\\ =&\|\tau(b_{c}^{*}c_{c})b_{r}^{*}\mathbb{E}_{M_{\Lambda\cap\textrm{Link}(\mathbf{v}_{c})}}(b_{l}^{*}u_{i}c_{l})c_{r}\|_{2}\\ =&\|\tau(b_{c}^{*}c_{c})b_{r}^{*}\mathbb{E}_{M_{\Lambda\cap\textrm{Link}(\mathbf{v}_{c})}}(\mathbb{E}_{M_{\Lambda_{j}}}(b_{l}^{*}u_{i}c_{l}))c_{r}\|_{2}\\ \leq&\|b_{c}\|_{2}\|c_{c}\|_{2}\|b_{r}\|\|c_{r}\|\|\mathbb{E}_{M_{\Lambda_{j}}}(b_{l}^{*}u_{i}c_{l})\|_{2}.\end{split}$
We see that this expression converges to 0 when $i\to\infty$. Thus, more generally, for $b,c\in M_{\Gamma}$ with $\mathbb{E}_{M_{\Lambda_{\operatorname{emb}}}}(b)=\mathbb{E}_{M_{\Lambda_{\operatorname{emb}}}}(c)=0$, we obtain $\|\mathbb{E}_{M_{\Lambda_{\operatorname{emb}}}}(b^{*}u_{i}c)\|_{2}\to 0$ when $i\to\infty$, which shows (1a).
(1b) Observe that if $x\in\operatorname{qNor}_{M_{\Gamma}}(A)$ then for some $n\geq 1$ and $x_{1},\ldots,x_{n}\in M_{\Gamma}$ we have $Ax\subseteq\sum_{k=1}^{n}x_{k}A\subseteq\sum_{k=1}^{n}x_{k}M_{\Lambda_{\operatorname{emb}}}$. Therefore by the existence of the net $(u_{i})_{i}$ shown by (1a) and by Lemma 5.7, we have that $1_{A}x1_{A}\in M_{\Lambda_{\operatorname{emb}}}$. This shows $1_{A}\operatorname{qNor}_{M_{\Gamma}}(A)1_{A}\subseteq M_{\Lambda_{\operatorname{emb}}}$ and thus proves (1b).
(1c) Let $u\in M_{\Gamma}$ be a unitary for which $u^{*}Au\subseteq M_{\Lambda_{\operatorname{emb}}}$. Then $Au\subseteq uM_{\Lambda_{\operatorname{emb}}}$ so again by the existence of the net $(u_{i})_{i}$ shown by (1a) and by Lemma 5.7, we obtain that $1_{A}u1_{A}\in M_{\Lambda_{\operatorname{emb}}}$.
(2) By replacing $\{\Lambda_{j}\}_{j\in\mathcal{J}}$ with $\{\Lambda_{j}\cap\Lambda\}_{j\in\mathcal{J}}$ we may assume that $\Lambda_{j}\subseteq\Lambda$ for $j\in\mathcal{J}$. We observe that $r$ is central in $A$, which we will use a number of times in the proof. By Lemma 5.4 the assumptions imply that there exist projections $p\in rA,q\in M_{\Lambda}$ a non-zero partial isometry $v\in qM_{\Gamma}p$ and a normal $\ast$-homomorphism $\theta:pAp\rightarrow qM_{\Lambda}q$ such that $\theta(x)v=vx$ for all $x\in pAp$ and such that moreover $\theta(pAp)\not\prec_{M_{\Lambda}}M_{\Lambda_{j}}$ for $j\in\mathcal{J}$. From (1) we see that $\theta(p)\operatorname{qNor}_{M_{\Gamma}}(\theta(pAp))\theta(p)\subseteq M_{\Lambda_{\operatorname{emb}}}$.
Now take $u\in\operatorname{Nor}_{M_{\Gamma}}(A)$. We follow the proof of [61, Lemma 3.5] or [43, Lemma 9.4]. Take $z\in A$ a central projection such that $z=\sum_{j=1}^{n}v_{j}v_{j}^{\ast}$ with $v_{j}\in A$ partial isometries such that $v_{j}^{\ast}v_{j}\leq p$. Then
$pzupz(pAp)\subseteq pzuA=pzAu=pAzu\subseteq\sum_{j=1}^{n}(pAv_{j})v_{j}^{\ast}u\subseteq\sum_{j=1}^{n}(pAp)v_{j}^{\ast}u,$
and similarly $(pAp)pzupz\subseteq\sum_{j=1}^{n}uv_{j}(pAp)$. We conclude that $pzupz\in\operatorname{qNor}_{pMp}(pAp)$.
Now if $x\in\operatorname{qNor}_{pM_{\Gamma}p}(pAp)$ then by direct verification we see that we have that $vxv^{\ast}\in\theta(p)\operatorname{qNor}_{qM_{\Gamma}q}(\theta(pAp))\theta(p)$. It follows that $vpzupzv^{\ast}$, with $u\in\operatorname{Nor}_{M_{\Gamma}}(A)$ as before, is contained in $\theta(p)\operatorname{qNor}_{qM_{\Gamma}q}(\theta(pAp))\theta(p)$ which was contained in $M_{\Lambda_{\operatorname{emb}}}$. We may take the projections $z$ to approximate the central support of $p$ and therefore $vuv^{\ast}=vpupv^{\ast}\in M_{\Lambda_{\operatorname{emb}}}$. Hence $v\operatorname{Nor}_{M_{\Gamma}}(A)^{\prime\prime}v^{\ast}\subseteq M_{\Lambda_{\operatorname{emb}}}$. Set $p_{1}=v^{\ast}v\in pA^{\prime}p$. Note that $p_{1}\leq p\leq r$. As both $A$ and $A^{\prime}$ are contained in $\operatorname{Nor}_{M_{\Gamma}}(A)^{\prime\prime}$ we find that $p_{1}\in\operatorname{Nor}_{M_{\Gamma}}(A)^{\prime\prime}$ (as $p\in A$). So we have the $\ast$-homomorphism $\rho:p_{1}\operatorname{Nor}_{M_{\Gamma}}(A)^{\prime\prime}p_{1}=p_{1}r\operatorname{Nor}_{M_{\Gamma}}(A)^{\prime\prime}p_{1}\rightarrow M_{\Lambda_{\operatorname{emb}}}:x\mapsto vxv^{\ast}$ with $v\in qM_{\Gamma}p_{1}$ and clearly $\rho(x)v=vx$. We conclude that $r\operatorname{Nor}_{M_{\Gamma}}(A)^{\prime\prime}\prec_{M_{\Gamma}}M_{\Lambda_{\operatorname{emb}}}$. ∎
We prove the following result concerning embeddings in graph products.
Proposition 5.9.
Let $\Gamma$ be a simple graph and for $v\in\Gamma$, and let $(M_{v},\tau_{v})$ be a tracial von Neumann algebra. Fix $v\in\Gamma$ and let $N\subseteq M_{v}$ be diffuse. If $N\prec_{M_{\Gamma}}M_{\Lambda}$ for some subgraph $\Lambda\subseteq\Gamma$, then $v\in\Lambda$. In particular if $\Lambda=\{w\}$, a singleton set, then $v=w$.
Proof.
Let $\Lambda\subseteq\Gamma$ be a subgraph with $v\not\in\Lambda$. We show that $N\not\prec_{M_{\Gamma}}M_{\Lambda}$. Since $N$ is diffuse, we can choose a net $(u_{k})_{k}$ of unitaries in $N$ such that $\tau(u_{k})=0$ and $u_{k}\rightarrow 0$ $\sigma$-weakly. Since $\lambda(\mathbf{M}_{\Gamma})$ is a dense subspace of $M_{\Gamma}$, it is sufficient to show for any reduced operators $x=x_{1}x_{2}\dots x_{m},y=y_{1}y_{2}\dots y_{n}$, s.t. $x_{i}\in\mathring{M}_{v_{i}}$, $y_{i}\in\mathring{M}_{w_{i}}$, we have $\|\mathbb{E}_{M_{\Lambda}}(xu_{k}y)\|_{2}\to 0$. Indeed, writing $x=x^{\prime}a$, $y=by^{\prime}$, where $a,b\in M_{v}$ and where $x^{\prime}$ respectively $y^{\prime}$ is a reduced operator without letter $v$ at the end respectively start. Then
$xu_{k}y=x^{\prime}au_{k}by^{\prime}=x^{\prime}\tau(au_{k}b)y^{\prime}+x^{\prime}(au_{k}b-\tau(au_{k}b))y^{\prime}.$
On the one hand, $\mathbb{E}_{M_{\Lambda}}(x^{\prime}\tau(au_{k}b)y^{\prime})=\tau(au_{k}b)\mathbb{E}_{M_{\Lambda}}(x^{\prime}y^{\prime})=\langle u_{k}b,a^{*}\rangle\mathbb{E}_{M_{\Lambda}}(x^{\prime}y^{\prime})\to 0$. On the other hand, we write $x^{\prime}=x^{\prime\prime}d$, $y^{\prime}=ey^{\prime\prime}$, where $d,e\in M_{\textrm{Link}(v)}$ and where $x^{\prime\prime}$ respectively $y^{\prime\prime}$ has no letter from $\textrm{Star}(v)$ at the end respectively at the start. Then we have
$\begin{split}x^{\prime}(au_{k}b-\tau(au_{k}b))y^{\prime}&=x^{\prime\prime}d(au_{k}b-\tau(au_{k}b))ey^{\prime\prime}\\ &=x^{\prime\prime}de(au_{k}b-\tau(au_{k}b))y^{\prime\prime}\\ &=\sum_{i}x^{\prime\prime}f_{i}(au_{k}b-\tau(au_{k}b))y^{\prime\prime},\end{split}$
where we write $de=\sum_{i}f_{i}$ and $f_{i}\in M_{Link(v)}$ reduced. Since $x^{\prime\prime}f_{i}(au_{k}b-\tau(au_{k}b))y^{\prime\prime}$ is reduced and $v\not\in\Lambda$ we obtain that $\mathbb{E}_{M_{\Lambda}}(x^{\prime\prime}f_{i}(au_{k}b-\tau(au_{k}b))y^{\prime\prime})=0$. Thus $\|\mathbb{E}_{M_{\Lambda}}(xu_{k}y)\|_{2}\to 0$, which completes the proof. ∎
Remark 5.10.
We remark in particular for any graph $\Gamma$, II1-factors $\{M_{v}\}_{v\in\Gamma}$ and a finite subgraph $\Lambda\subseteq\Gamma$ that $\operatorname{qNor}_{M_{\Gamma}}(M_{\Lambda})^{\prime\prime}=\operatorname{Nor}_{M_{\Gamma}}(M_{\Lambda})^{\prime\prime}=M_{\Lambda\cup\textrm{Link}(\Lambda)}$. Indeed, clearly $M_{\Lambda},M_{\textrm{Link}(\Lambda)}\subseteq\operatorname{Nor}_{M_{\Gamma}}(M_{\Lambda})^{\prime\prime}$ (as $M_{\textrm{Link}(\Lambda)}=M_{\Lambda}^{\prime}\cap M_{\Gamma}$) so that $M_{\Lambda\cup\textrm{Link}(\Lambda)}\subseteq\operatorname{Nor}_{M_{\Gamma}}(M_{\Lambda})^{\prime\prime}\subseteq\operatorname{qNor}_{M_{\Gamma}}(M_{\Lambda})^{\prime\prime}$. Furthermore, by Proposition 5.9 we have $M_{\Lambda}\not\prec_{M_{\Gamma}}M_{\widetilde{\Lambda}}$ for any strict subgraph $\widetilde{\Lambda}\subsetneq\Lambda$ so that by Proposition 5.8 we obtain $\operatorname{qNor}_{M_{\Gamma}}(M_{\Lambda})^{\prime\prime}\subseteq M_{\Lambda\cup\textrm{Link}(\Lambda)}$.
5.4. Unitary conjugacy in graph products
We prove Theorem 5.11 which gives sufficient conditions for a subalgebra $Q\subseteq M_{\Gamma}$ to unitarily embed in a subalgebra $M_{\Lambda_{\operatorname{emb}}}$. This can be seen as a generalization of [52, Theorem 3.3] where a unitary embedding is proven for free products. The proof of Theorem 5.11 combines (the second half of) the proof of [52, Theorem 3.3] with results of Section 5.3 concerning embeddings in graph products.
Theorem 5.11.
Let $\Gamma$ be a simple graph and for $v\in\Gamma$ let $(M_{v},\tau_{v})$ be a II1-factor with normal faithful trace $\tau_{v}$. Let $Q\subseteq M_{\Gamma}$ be a subfactor whose relative commutant $Q^{\prime}\cap M_{\Gamma}$ is also a factor. Let $\Lambda\subseteq\Gamma$ be a subgraph and let $\{\Lambda_{j}\}_{j\in\mathcal{J}}$ be a non-empty collection of subgraphs of $\Lambda$. Suppose $Q\prec_{M_{\Gamma}}M_{\Lambda}$ and $Q\not\prec_{M_{\Gamma}}M_{\Lambda_{j}}$ for $j\in\mathcal{J}$. Then there is a unitary $u\in M_{\Gamma}$ such that $u^{*}Qu\subseteq M_{\Lambda_{\operatorname{emb}}}$, where $\Lambda_{\operatorname{emb}}$ is defined as in (5.8).
Proof.
Since $Q\prec_{M_{\Gamma}}M_{\Lambda}$ and $Q\not\prec_{M_{\Gamma}}M_{\Lambda_{j}}$ for $j\in\mathcal{J}$ we have by Lemma 5.4 that there are projections $q\in Q$, $e\in M_{\Lambda}$, a normal *-homomorphism $\theta:qQq\to eM_{\Lambda}e$ and a non-zero partial isometry $v\in eM_{\Gamma}q$ such that $\theta(x)v=vx$ for $x\in qQq$ and such that moreover $\theta(qQq)\not\prec_{M_{\Gamma}}M_{\Lambda_{j}}$ for $j\in\mathcal{J}$. We may moreover assume that $q$ is majorized by the support of $\mathbb{E}_{Q}(v^{\ast}v)$. Let $q_{0}\in Q$ be a non-zero projection with $q_{0}\leq q$ and trace $\tau(q_{0})=\frac{1}{m}$ for some $m\geq 1$. Put $v_{0}:=vq_{0}$. Note that $v^{\ast}v\in(qQq)^{\prime}\cap qM_{\Gamma}q$. Then $\mathbb{E}_{Q}(v_{0}^{\ast}v_{0})=\mathbb{E}_{Q}(q_{0}v^{\ast}vq_{0})=\mathbb{E}_{Q}(q_{0}v^{\ast}v)=q_{0}\mathbb{E}_{Q}(v^{\ast}v)$ and the latter expression is non-zero by the assumption that the support of $\mathbb{E}_{Q}(v^{\ast}v)$ majorizes $q$. As $\mathbb{E}_{Q}$ is faithful $v_{0}\not=0$. Define $\theta_{0}:q_{0}Qq_{0}\to eM_{\Lambda}e$ as $\theta_{0}:=\theta|_{q_{0}Qq_{0}}$. Then for $x\in q_{0}Qq_{0}$ we have $\theta_{0}(x)v_{0}=\theta(x)vq_{0}=vxq_{0}=v_{0}x$. Automatically this implies $v_{0}^{\ast}v_{0}\in(q_{0}Qq_{0})^{\prime}\cap q_{0}M_{\Gamma}q_{0}$. Furthermore for $j\in\mathcal{J}$, the corner $\theta_{0}(q_{0}Qq_{0})=\theta_{0}(q_{0})\theta(qQq)\theta(q_{0})$ does not embed in $M_{\Lambda_{j}}$ inside $M_{\Gamma}$ since $\theta(qQq)$ does not embed in $M_{\Lambda_{j}}$ inside $M_{\Gamma}$. Hence, by Proposition 5.8(1b) we obtain $\theta(q_{0})\operatorname{Nor}_{M_{\Gamma}}(\theta_{0}(q_{0}Qq_{0}))^{\prime\prime}\theta(q_{0})\subseteq M_{\Lambda_{\operatorname{emb}}}$.
Since $Q$ is a factor and $\tau(q_{0})=\frac{1}{m}$ we can for $j=1,\ldots,m$ choose a partial isometry $u_{j}$ in $Q$ such that $u_{j}^{*}u_{j}=q_{0}$ and $\sum_{j=1}^{m}u_{j}u_{j}^{*}=1_{M_{\Gamma}}$. We may moreover assume that $u_{1}=q_{0}$. We define a projection $q^{\prime}:=\sum_{j=1}^{m}u_{j}v_{0}^{*}v_{0}u_{j}^{*}\in M_{\Gamma}$. We show that $q^{\prime}\in Q^{\prime}\cap M_{\Gamma}$. Indeed, let $y\in Q$. Then using that $v_{0}^{\ast}v_{0}\in(q_{0}Qq_{0})^{\prime}$ and $u_{j}^{*}yu_{j}\in q_{0}Qq_{0}$ for $j=1,\ldots,n$ we get
$\begin{split}q^{\prime}y&=\sum_{j=1}^{m}u_{j}(v_{0}^{\ast}v_{0})u_{j}^{*}y=\sum_{j=1}^{m}\sum_{i=1}^{m}u_{j}(v_{0}^{\ast}v_{0})(u_{j}^{*}yu_{i})u_{i}^{*}\\ &=\sum_{j=1}^{m}\sum_{i=1}^{m}u_{j}(u_{j}^{*}yu_{i})(v_{0}^{\ast}v_{0})u_{i}^{*}=\sum_{i=1}^{m}yu_{i}(v_{0}^{\ast}v_{0})u_{i}^{*}=yq^{\prime}.\end{split}$
and thus $q^{\prime}\in Q^{\prime}\cap M_{\Gamma}$. We observe that $v_{0}^{\ast}v_{0}=q_{0}q^{\prime}q_{0}=q_{0}q^{\prime}$ which shows in particular that $q^{\prime}$ is non-zero (since $v_{0}\not=0$). Since $Q^{\prime}\cap M$ is a (finite) factor and $q^{\prime}$ is a non-zero projection, we can choose a projection $q_{0}^{\prime}\in Q^{\prime}\cap M$ with $q_{0}^{\prime}\leq q^{\prime}$ and $\tau(q_{0}^{\prime})=\frac{1}{n}$ for some $n\geq 1$. Since $Q^{\prime}\cap M_{\Gamma}$ is a factor and since $\tau(q_{0}^{\prime})=\frac{1}{n}$ we can find partial isometries $u_{1}^{\prime},\ldots,u_{n}^{\prime}\in Q^{\prime}\cap M_{\Gamma}$ with $(u_{k}^{\prime})^{*}u_{k}^{\prime}=q_{0}^{\prime}$ for $k=1,\ldots,n$ and such that $\sum_{k=1}^{n}u_{k}^{\prime}(u_{k}^{\prime})^{*}=1_{M_{\Gamma}}$.
We then put $v_{00}:=v_{0}q_{0}^{\prime}=vq_{0}q_{0}^{\prime}\in eM_{\Gamma}q_{0}$. Observe that $v_{00}^{\ast}v_{00}=q_{0}^{\prime}v_{0}^{\ast}v_{0}q_{0}^{\prime}=q_{0}^{\prime}q_{0}$ has trace $\tau(v_{00}^{*}v_{00})=\tau(q_{0}^{\prime})\tau(q_{0})=\frac{1}{nm}$ so in particular $v_{00}$ is non-zero. Then for $x\in q_{0}Qq_{0}$ we have $\theta_{0}(x)v_{00}=\theta_{0}(x)v_{0}q_{0}^{\prime}=v_{0}xq_{0}^{\prime}=v_{00}x$. Therefore, we obtain $v_{00}v_{00}^{\ast}\in\theta_{0}(q_{0}Qq_{0})^{\prime}\cap M_{\Gamma}\subseteq\operatorname{Nor}_{M_{\Gamma}}(\theta_{0}(q_{0}Qq_{0}))^{\prime\prime}$. As $v_{00}v_{00}^{\ast}\leq\theta(q_{0})$ we obtain $v_{00}v_{00}^{\ast}\in\theta(q_{0})\operatorname{Nor}_{M_{\Gamma}}(\theta_{0}(q_{0}Qq_{0}))^{\prime\prime}\theta(q_{0})\subseteq M_{\Lambda_{\operatorname{emb}}}$ using the first paragraph.
Since $M_{\Lambda_{\operatorname{emb}}}$ is a factor (as it is a graph product of II1-factors), and since $\tau(v_{00}v_{00}^{\ast})=\frac{1}{nm}$ there exist for $j=1,\ldots,m$, $k=1,\ldots,n$ partial isometries $w_{j,k}\in M_{\Lambda_{\operatorname{emb}}}$ with $w_{j,k}w_{j,k}^{*}=v_{00}v_{00}^{\ast}$ and $\sum_{j=1}^{m}\sum_{k=1}^{n}w_{j,k}^{*}w_{j,k}=1_{M_{\Gamma}}$. Finally, we define the unitary $u:=\sum_{j=1}^{m}\sum_{k=1}^{n}u_{j}u_{k}^{\prime}v_{00}^{\ast}w_{j,k}\in M_{\Gamma}$. Then for $x\in Q$ we have
$u^{*}xu$ $=\sum_{j_{1}=1}^{m}\sum_{k_{1}=1}^{n}\sum_{j_{2}=1}^{m}\sum_{k_{2}=1}^{n}w_{j_{1},k_{1}}^{*}v_{00}(u_{k_{1}}^{\prime})^{*}(u_{j_{1}}^{*}xu_{j_{2}})u_{k_{2}}^{\prime}v_{00}^{*}w_{j_{1},k_{2}}$
$=\sum_{j_{1}=1}^{m}\sum_{k=1}^{n}\sum_{j_{2}=1}^{m}w_{j_{1},k}^{*}v_{00}(u_{j_{1}}^{*}xu_{j_{2}})q_{0}^{\prime}v_{00}^{\ast}w_{j_{1},k}$
$=\sum_{j_{1}=1}^{m}\sum_{k=1}^{n}\sum_{j_{2}=1}^{m}w_{j_{1},k}^{*}\theta_{0}(u_{j_{1}}^{*}xu_{j_{2}})v_{00}v_{00}^{\ast}w_{j_{1},k}$
$=\sum_{j_{1}=1}^{m}\sum_{k=1}^{n}\sum_{j_{2}=1}^{m}w_{j_{1},k}^{*}\theta_{0}(u_{j_{1}}^{*}xu_{j_{2}})w_{j_{1},k}\in M_{\Lambda_{\operatorname{emb}}}.$
Hence $u^{*}Qu\subseteq M_{\Lambda_{\operatorname{emb}}}$. ∎
6. Graph product rigidity
The aim of this section is to prove Theorem 6.18. This provides a rather general class of graphs and von Neumann algebras such that the graph product completely remembers the graph and the vertex von Neumann algebra up to stable isomorphism. Note that we cannot expect to cover all graphs as this would imply the free factor problem and which is beyond reach of our methods. The class of rigid graphs as presented in Section 3 is therefore natural.
6.1. Vertex von Neumann algebras
We define classes of von Neumann algebras for which we first recall a version of the Akemann-Ostrand property [41].
Definition 6.1.
Let $M$ be a von Neumann algebra with standard form $(M,L^{2}(M),J,L^{2}(M)^{+})$. We say that $M$ has strong property (AO) if there exist unital C-subalgebras $A\subseteq M$ and $C\subseteq\mathbb{B}(L^{2}(M))$ such that:
  • AA is $\sigma$-weakly dense in $M$,
  • $C$ is nuclear and contains $A$,
  • The commutators $C,JAJ]=\{[c,JaJ]\mid c\in C,a\in A\}$ are contained in the space of compact operators $\mathbb{K}(L^{2}(M))$.
We recall that a wide class of examples of von Neumann algebras with property strong (AO) comes from hyperbolic groups.
Theorem 6.2 (See Lemma 3.1.4 of [46] and remarks before).
Let $G$ be a discrete hyperbolic group. Consider the anti-linear isometry $J$ determined by
$J:\ell^{2}(G)\rightarrow\ell^{2}(G):\delta_{s}\mapsto\delta_{s^{-1}},\qquad s\in G.$
Then there is a nuclear C-algebra $C$ such that:
  1. (1)
    $C_{r}^{\ast}(G)\subseteq C\subseteq\mathbb{B}(\ell^{2}(G))$.
  2. (2)
    $C$ contains all compact operators.
  3. (3)
    The commutator $C,JC_{r}^{\ast}(G)J]$ is contained in the space of compact operators.
Remark 6.3.
In view of Section 4 it is worth to note that we may always assume without loss of generality that $C$ contains the space of compact operators by replacing $C$ by $C+\mathbb{K}(L^{2}(M))$ if necessary, see [41, Remark 2.7]. This fact also underlies Theorem 6.2.
Definition 6.4.
We define the following classes of von Neumann algebras:
  • Let $\mathcal{C}_{{\rm Vertex}}$ be the class of II1-factors $M$ with separable predual $M_{\ast}$ that satisfy condition strong (AO) and which are non-amenable;
  • Let $\mathcal{C}_{{\rm Complete}}$ be the class of all von Neumann algebraic graph products $(M_{\Gamma},\tau)=*_{v,\Gamma}(M_{v},\tau_{v})$ of tracial von Neumann algebras $(M_{v},\tau_{v})$ in $\mathcal{C}_{{\rm Vertex}}$ taken over non-empty, finite, complete graphs $\Gamma$;
  • Let $\mathcal{C}_{{\rm Rigid}}$ be the class of all von Neumann algebraic graph products $(M_{\Gamma},\tau)=*_{v,\Gamma}(M_{v},\tau_{v})$ of tracial von Neumann algebras $(M_{v},\tau_{v})$ in $\mathcal{C}_{{\rm Vertex}}$ taken over non-empty, rigid graphs $\Gamma$.
  • Let $\mathcal{C}_{{\rm Rigid}}^{f}$ be defined in the same way as $\mathcal{C}_{{\rm Rigid}}$ with the additional assumption that $\Gamma$ is finite.
Remark 6.5.
We remark that $\mathcal{C}_{{\rm Vertex}}\subseteq\mathcal{C}_{{\rm Complete}}\subseteq\mathcal{C}_{{\rm Rigid}}^{f}\subseteq\mathcal{C}_{{\rm Rigid}}$. Furthermore,
  1. (1)
    The class $\mathcal{C}_{{\rm Vertex}}$ is closed under taking (finitely many) free products (see [41, Example 2.8(5)]). Furthermore, all von Neumann algebras $M\in\mathcal{C}_{{\rm Vertex}}$ are solid and prime, see [54];
  2. (2)
    The class $\mathcal{C}_{{\rm Complete}}$ is closed under taking tensor products. Moreover, we observe that $\mathcal{C}_{{\rm Complete}}$ coincides with the class of tensor products of factors from $\mathcal{C}_{{\rm Vertex}}$;
  3. (3)
    The class $\mathcal{C}_{{\rm Rigid}}$ is closed under taking graph products over non-empty, rigid graphs by Remark 3.4 and Lemma 3.5. In particular, it is closed under tensor products;
  4. (4)
    The class $\mathcal{C}_{{\rm Rigid}}\setminus\mathcal{C}_{{\rm Vertex}}$ is closed under taking graph products over arbitrary non-empty, graphs by Remark 3.4 and Lemma 3.5. In particular, it is closed under tensor products and under free products.
Remark 6.6.
We show that it may happen that a graph product over a rigid graph is isomorphic to a graph product over a non-rigid graph; even if all vertex von Neumann algebras come from the class $\mathcal{C}_{{\rm Vertex}}$. Consider the graph $\mathbb{Z}_{4}$ defined in Example 3.2(3). The graph $\mathbb{Z}_{4}$ is rigid. For $v\in\mathbb{Z}_{4}$ let $G_{v}$ be a countable discrete group. Let $H_{v}=G_{v}\ast G_{v+2}$. We have for the graph products of groups that
$\ast_{v,\mathbb{Z}_{4}}G_{v}=(G_{0}\ast G_{2})\times(G_{1}\ast G_{3})=\ast_{v,\mathbb{Z}_{2}}H_{v}.$
We now set $G_{v}=\mathbb{F}_{2}$ and $H_{v}=\mathbb{F}_{4}$ to be free groups with 2 and 4 generators respectively. Set $M_{v}=\mathcal{L}(\mathbb{F}_{2}),v\in\mathbb{Z}_{4}$ and $N_{v}=\mathcal{L}(\mathbb{F}_{4}),v\in\mathbb{Z}_{2}$ equipped with their tracial Plancherel states $\tau_{v}$. Then $M_{v}$ and $N_{v}$ are in class $\mathcal{C}_{{\rm Vertex}}$ and $\ast_{v,\mathbb{Z}_{4}}(M_{v},\tau_{v})=\ast_{v,\mathbb{Z}_{2}}(N_{v},\tau_{v})$. We have thus given an example of a rigid and non-rigid graph that give isomorphic graph products.
6.2. Key result for embeddings of diffuse subalgebras in graph products
In this section we fix the following notation. Let $\Gamma$ be a simple graph. For $v\in\Gamma$ let $(M_{v},\tau_{v})$ be a tracial von Neumann algebra ($M_{v}\not$$\mathbb{C}$) that satisfies strong (AO) and has a separable predual. Let $(M_{\Gamma},\tau_{\Gamma})=\ast_{v,\Gamma}(M_{v},\tau_{v})$ be the von Neumann algebraic graph product. For $v\in\Gamma$ let $\mathcal{H}_{v}=L^{2}(M_{v},\tau_{v})$ and let $\mathcal{H}_{\Gamma}$ be the graph product of these Hilbert spaces, which is the standard Hilbert space of $M_{\Gamma}$ [14]. We denote by $J:\mathcal{H}_{\Gamma}\to\mathcal{H}_{\Gamma}$ the modular conjugation. Let $B_{v}=\mathbb{B}(\mathcal{H}_{v})$. Let $\Omega_{v}=1_{M_{v}}$ as a vector in $\mathcal{H}_{v}$ and let $\omega_{v}(x)=\langle x\Omega_{v},\Omega_{v}\rangle,x\in B_{v}$. Then $\omega_{v}$ is a GNS-faithful, but not faithful, state on $B_{v}$ and the GNS-space of $\omega_{v}$ can canonically be identified with $\mathcal{H}_{v}$. The reduced C-algebraic graph product $(B_{\Gamma},\omega_{\Gamma})=*^{\min}_{v,\Gamma}(B_{v},\omega_{v})$ gives then by construction a C-subalgebra $B$ of $\mathbb{B}(\mathcal{H}_{\Gamma})$. We let $\lambda_{v}:B_{v}\rightarrow B$ be the canonical embedding. Furthermore we let $\rho_{v}:B_{v}^{\operatorname{op}}\to B^{\operatorname{op}}$ be the map $\rho_{v}(x^{\operatorname{op}})=J\lambda_{v}(x)^{*}J$. As for $v\in\Gamma$ the von Neumann algebra $M_{v}$ has strong property (AO) by assumption (as $M_{v}\in\mathcal{C}_{{\rm Vertex}}$), there are unital $\operatorname{C^{*}}$-subalgebras $C_{v}\subseteq B_{v}$ and $A_{v}\subseteq M_{v}\cap C_{v}$ such that
  1. (1)
    The C-algebra $A_{v}$ are $\sigma$-weakly dense in $M_{v}$,
  2. (2)
    The C-algebra $C_{v}$ are nuclear,
  3. (3)
    The commutators $C_{v},J_{v}A_{v}J_{v}]$ are contained in $\mathbb{K}(\mathcal{H}_{v})$.
As in [41, Remarks 2.7 (1)] we may and will moreover assume that $\mathbb{K}(\mathcal{H}_{v})\subseteq C_{v}$. We let $(C_{\Gamma},\omega_{\Gamma})=*^{\min}_{v,\Gamma}(C_{v},\omega_{v})$ and $(A_{\Gamma},\omega_{\Gamma})=*^{\min}_{v,\Gamma}(A_{v},\omega_{v})$ be the reduced graph products of the C*-algebras. Observe that we now have
$A_{\Gamma}\subseteq M_{\Gamma}\subseteq B_{\Gamma}\text{ and }A_{\Gamma}\subseteq C_{\Gamma}\subseteq B_{\Gamma},$
and the states $\omega_{\Gamma}$ defined through the different graph products coincide.
Lemma 6.7.
$C_{\Gamma}$ is nuclear.
Proof.
The vector $\Omega_{v}$ is cyclic for $M_{v}$. Furthermore, $A_{v}$ is $\sigma$-weakly dense in $M_{v}$ by assumption and so $\Omega_{v}$ is also cyclic for $A_{v}$. It follows that the GNS-representation $\pi_{v}$ of $C_{v}$ with respect to $\omega_{v}$ is unitarily equivalent with the canonical representation given by the inclusion $C_{v}\subseteq\mathbb{B}(\mathcal{H}_{v})$, see [25, Theorem VIII.5.14 (b)]. We assumed that $\mathbb{K}(\mathcal{H}_{v})\subseteq C_{v}$ and that $C_{v}$ is nuclear and so we may apply Theorem 4.4 to conclude that $C_{\Gamma}$ is nuclear. ∎
Definition 6.8.
For $\Lambda\subseteq\Gamma$ we define the C-algebra
$D_{\Lambda}=U_{\Lambda}^{\prime}(\mathbb{K}(\mathcal{H}^{\prime}(\Lambda))\otimes\mathbb{B}(\mathcal{H}_{\Lambda}))(U^{\prime}_{\Lambda})^{*}.$
The tensor product in the definition of $D_{\Lambda}$ is understood as the spatial (minimal) tensor product, which is the norm closure of the algebraic tensors acting on the tensor product Hilbert space. In particular $D_{\varnothing}=\mathbb{K}(\mathcal{H}_{\Gamma})$.
Lemma 6.9.
Let $v\in\Gamma$. We have $B_{\Gamma}D_{\textrm{Link}(v)}B_{\Gamma}\subseteq D_{\textrm{Link}(v)}$.
Proof.
We note that the proof we give here in particular also works if $\textrm{Link}(v)$ is empty; though in that case the statement trivially follows from the fact that $D_{\varnothing}=\mathbb{K}(\mathcal{H}_{\Gamma})$ is an ideal in $\mathbb{B}(\mathcal{H}_{\Gamma})$. Take $x\in\mathbb{B}(\mathcal{H}_{w})$. Then if $w\not\in\textrm{Link}(v)$ we have that $\mathcal{H}^{\prime}(\textrm{Link}(v))$ is an invariant subspace of $x$ and
(6.1) $x=U_{\textrm{Link}(v)}^{\prime}(x\otimes 1)U_{\textrm{Link}(v)}^{\prime\ast}.$
Now suppose that $w\in\textrm{Link}(v)$. Let $P$ be the orthogonal projection of $\mathcal{H}^{\prime}(\textrm{Link}(v))$ onto $\mathcal{H}^{\prime}(\textrm{Link}(v))\cap\mathcal{H}_{\textrm{Link}(w)}$. Then
(6.2) $x=U_{\textrm{Link}(v)}^{\prime}(xP^{\perp}\otimes 1)U_{\textrm{Link}(v)}^{\prime\ast}+U_{\textrm{Link}(v)}^{\prime}(P\otimes x)U_{\textrm{Link}(v)}^{\prime\ast}.$
From the decompositions (6.1), (6.2) we see that $xD_{\textrm{Link}(v)},D_{\textrm{Link}(v)}x\subseteq D_{\textrm{Link}(v)}$. As $B_{\Gamma}$ is the closed linear span of products of elements in $\mathbb{B}(\mathcal{H}_{w}),w\in\Gamma$ the proof follows. ∎
Denote $P_{\Omega}$ for the orthogonal projection onto $\mathbb{C}\Omega$.
Lemma 6.10.
Let $v,w\in\Gamma$. Let $a\in B_{v}$, $b\in B_{w}$. Then
(6.3) $[a,JbJ]=\begin{cases}U_{\textrm{Star}(v)}^{\prime}(P_{\Omega}\otimes[a,JbJ])(U_{\textrm{Star}(v)}^{\prime})^{*},&v=w;\\ 0,&v\not=w.\end{cases}$
Proof.
If $v\not=w$ then the result follows from [14, Proposition 3.3]. Suppose $v=w$. Let $\mathbf{v}_{1}\in\mathcal{W}_{\Gamma}^{\prime}(\textrm{Star}(v))$ and $\mathbf{v}_{2}\in\mathcal{W}_{\textrm{Star}(v)}$, and put $\mathbf{v}=\mathbf{v}_{1}\mathbf{v}_{2}$. Let $\eta_{1}\in\mathring{\mathcal{H}}_{\mathbf{v}_{1}}$, $\eta_{2}\in\mathring{\mathcal{H}}_{\mathbf{v}_{2}}$ be pure tensors and denote $\eta:=U_{\textrm{Star}(v)}^{\prime}(\eta_{1}\otimes\eta_{2})\in\mathring{\mathcal{H}}_{\mathbf{v}}$. We claim
(6.4) $a\eta=\begin{cases}U_{\textrm{Star}(v)}^{\prime}(\eta_{1}\otimes(a\eta_{2})),&\text{ if }\mathbf{v}_{1}=e;\\ U_{\textrm{Star}(v)}^{\prime}((a\eta_{1})\otimes\eta_{2}),&\text{ if }\mathbf{v}_{1}\not=e.\end{cases}$
Indeed, if $\mathbf{v}_{1}=e$ then $\eta_{1}=\Omega$ and $\eta=\eta_{2}$ up to scalar multiplication, so that $U_{\textrm{Star}(v)}^{\prime}(\eta_{1}\otimes(a\eta_{2}))=a\eta_{2}=a\eta$. Thus suppose $\mathbf{v}_{1}\not=e$. Then it follows that $v\mathbf{v}_{1}\in\mathcal{W}^{\prime}(\textrm{Star}(v))$ since $\mathbf{v}_{1}\in\mathcal{W}^{\prime}(\textrm{Star}(v))$.
Suppose $v\mathbf{v}$ is reduced. Then also $v\mathbf{v}_{1}$ is reduced, and we have $a\eta=\lambda_{(v,e,e)}(a)\eta$ and $a\eta_{1}=\lambda_{(v,e,e)}(a)\eta_{1}$. It follows from [5, Lemma 2.3(iii)] that
$a\eta$ $=\lambda_{(v,e,e)}(a)\eta$
$=\mathcal{Q}_{(v\mathbf{v}_{1},\mathbf{v}_{2})}((\lambda_{(v,e,e)}(a)\eta_{1})\otimes\eta_{2})$
$=\mathcal{Q}_{(v\mathbf{v}_{1},\mathbf{v}_{2})}((a\eta_{1})\otimes\eta_{2}),$
and therefore, as $v\mathbf{v}_{1}\in\mathcal{W}^{\prime}(\textrm{Star}(v))$ and $\mathbf{v}_{2}\in\mathcal{W}_{\textrm{Star}(v)}$, we obtain $a\eta=U_{\textrm{Star}(v)}^{\prime}((a\eta_{1})\otimes\eta_{2})$.
Now, suppose $v\mathbf{v}$ is not reduced. Then also $v\mathbf{v}_{1}$ is not reduced as $\mathbf{v}_{1}\in\mathcal{W}^{\prime}(\textrm{Star}(v))$ and $\mathbf{v}_{1}\not=e$. We have $a\eta=\lambda_{(e,v,e)}(a)\eta+\lambda_{(e,e,v)}(a)\eta$ and $a\eta_{1}=\lambda_{(e,v,e)}(a)\eta_{1}+\lambda_{(e,e,v)}(a)\eta_{1}$. Again, using [5, Lemma 2.3(iii)] we obtain
$a\eta$ $=\lambda_{(e,v,e)}(a)\eta+\lambda_{(e,e,v)}(a)\eta$
$=\mathcal{Q}_{(\mathbf{v}_{1},\mathbf{v}_{2})}((\lambda_{(e,v,e)}(a)\eta_{1})\otimes\eta_{2})+\mathcal{Q}_{(v\mathbf{v}_{1},\mathbf{v}_{2})}((\lambda_{(e,e,v)}(a)\eta_{1})\otimes\eta_{2}).$
And thus
$a\eta=U_{\textrm{Star}(v)}^{\prime}(\lambda_{(e,v,e)}(a)\eta_{1})\otimes\eta_{2})+U_{\textrm{Star}(v)}^{\prime}(\lambda_{(e,e,v)}(a)\eta_{1})\otimes\eta_{2})=U_{\textrm{Star}(v)}^{\prime}((a\eta_{1})\otimes\eta_{2})).$
This shows (6.4).
We now claim that
(6.5) $JbJ\eta=U_{\textrm{Star}(v)}^{\prime}(\eta_{1}\otimes JbJ\eta_{2}).$
First, by [15, Proposition 2.20]we observe that $J\eta_{1}\in\mathring{\mathcal{H}}_{\mathbf{v}_{1}^{-1}}$, $J\eta_{2}\in\mathring{\mathcal{H}}_{\mathbf{v}_{2}^{-1}}$ and $J\eta=J\mathcal{Q}_{(\mathbf{v}_{1},\mathbf{v}_{2})}(\eta_{1}\otimes\eta_{2})=\mathcal{Q}_{(\mathbf{v}_{2}^{-1},\mathbf{v}_{1}^{-1})}(J\eta_{2}\otimes J\eta_{1})\in\mathring{\mathcal{H}}_{\mathbf{v}^{-1}}$. Furthermore, note that $v\mathbf{v}_{2}^{-1}=\mathbf{v}_{2}^{-1}v$ and $v\mathbf{v}_{2}\in W_{\textrm{Star}(v)}$.
Suppose that $v\mathbf{v}^{-1}$ is reduced. Then $v\mathbf{v}_{2}^{-1}$ is also reduced. Hence, similar as before we obtain $bJ\eta=\mathcal{Q}_{(v\mathbf{v}_{2}^{-1},\mathbf{v}_{1}^{-1})}((bJ\eta_{2})\otimes J\eta_{1})$. Hence
$JbJ\eta=\mathcal{Q}_{(\mathbf{v}_{1},v\mathbf{v}_{2})}(\eta_{1}\otimes(JbJ\eta_{2}))=U_{\textrm{Star}(v)}^{\prime}(\eta_{1}\otimes(JbJ\eta_{2})).$
Now, suppose that $v\mathbf{v}^{-1}$ is not reduced. Then $v\mathbf{v}_{2}^{-1}$ is not reduced. Similar as before we obtain
$bJ\eta=\mathcal{Q}_{(\mathbf{v}_{2}^{-1},\mathbf{v}_{1}^{-1})}((\lambda_{(e,v,e)}(b)J\eta_{2})\otimes J\eta_{1})+\mathcal{Q}_{(v\mathbf{v}_{2}^{-1},\mathbf{v}_{1}^{-1})}((\lambda_{(e,e,v)}(b)J\eta_{2})\otimes J\eta_{1}).$
Hence
$JbJ\eta$ $=\mathcal{Q}_{(\mathbf{v}_{1},\mathbf{v}_{2})}(\eta_{1}\otimes(\lambda_{(e,v,e)}(b)J\eta_{2}))+\mathcal{Q}_{(\mathbf{v}_{1},v\mathbf{v}_{2})}(\eta_{1}\otimes(\lambda_{(e,e,v)}(b)J\eta_{2}))$
$=U_{\textrm{Star}(v)}^{\prime}(\eta_{1}\otimes(J\lambda_{(e,v,e)}(b)J\eta_{2}))+U_{\textrm{Star}(v)}^{\prime}(\eta_{1}\otimes(J\lambda_{(e,e,v)}(b)J\eta_{2}))$
$=U_{\textrm{Star}(v)}^{\prime}(\eta_{1}\otimes(JbJ\eta_{2})),$
which shows (6.5). Now, combining (6.4) with (6.5) we obtain
$[a,JbJ]\eta=\begin{cases}U_{\textrm{Star}(v)}^{\prime}(\eta_{1}\otimes([a,JbJ]\eta_{2})),&\text{ if }\mathbf{v}_{1}=e;\\ 0,&\text{ if }\mathbf{v}_{1}\not=e\end{cases}$
and the statement follows. ∎
Lemma 6.11.
For $v,w\in\Gamma,c\in C_{v},a\in A_{w}$ we have $c,JaJ]\in D_{\textrm{Link}(v)}$.
Proof.
If $v\not=w$ it actually holds since by [14, Proposition 2.3] $[c,JaJ]=0$. So assume $v=w$. Lemma 6.10 gives that
(6.6) $c,JaJ]=U_{\textrm{Star}(v)}^{\prime}(P_{\Omega}\otimes[c,JaJ])(U_{\textrm{Star}(v)}^{\prime})^{*}.$
In what follows we will use the decomposition of Section 2.6 applied to $\textrm{Link}(v)$ as a subgraph of $\textrm{Star}(v)$, opposed to $\textrm{Link}(v)$ as a subgraph of $\Gamma$, and correspondingly define the Hilbert space $\mathcal{H}^{\prime}(\textrm{Link}(v))$ with respect to this inclusion. We thus have a natural unitary
$U_{\textrm{Link}(v)}^{\prime\prime}:\mathcal{H}^{\prime}(\textrm{Link}(v))\otimes\mathcal{H}_{\textrm{Link}(v)}\rightarrow\mathcal{H}_{\textrm{Star}(v)}.$
Further as $v$ commutes with all vertices in $\textrm{Link}(v)$ it follows that with respect to this decomposition we have $\mathcal{H}^{\prime}(\textrm{Link}(v))=\mathcal{H}_{v}$. So
$U_{\textrm{Link}(v)}^{\prime\prime}:\mathcal{H}_{v}\otimes\mathcal{H}_{\textrm{Link}(v)}\rightarrow\mathcal{H}_{\textrm{Star}(v)}.$
For $x\in\mathbb{B}(\mathcal{H}_{v})$ we get that
(6.7) $x=U_{\textrm{Link}(v)}^{\prime\prime}(x\otimes{\rm 1})(U_{\textrm{Link}(v)}^{\prime\prime})^{*}.$
Set the unitary
$U_{v}^{\prime\prime}:=U_{\textrm{Star}(v)}^{\prime}(1\otimes U_{\textrm{Link}(v)}^{\prime\prime}):\mathcal{H}^{\prime}(\textrm{Star}(v))\otimes\mathcal{H}_{v}\otimes\mathcal{H}_{\textrm{Link}(v)}\rightarrow\mathcal{H}_{\Gamma}.$
Combining (6.6) and (6.7) we have
$c,JaJ]=U_{v}^{\prime\prime}(P_{\Omega}\otimes[c,JaJ]\otimes 1)U_{v}^{\prime\prime\ast},$
where $[c,JaJ]$ on the left hand side acts on $\mathcal{H}_{\Gamma}$ and on the right hand side on $\mathcal{H}_{v}$. As we assumed $c,JaJ]\in\mathbb{K}(\mathcal{H}_{v})$ it follows that $[c,JaJ]$ is contained in
$U_{v}^{\prime\prime}(\mathbb{K}(\mathcal{H}^{\prime}(\textrm{Star}(v)))\otimes\mathbb{K}(\mathcal{H}_{v})\otimes 1)U_{v}^{\prime\prime\ast}=U_{\textrm{Link}(v)}^{\prime}(\mathbb{K}(\mathcal{H}^{\prime}(\textrm{Link}(v)))\otimes 1)U_{\textrm{Link}(v)}^{\prime\ast}\subseteq D_{\textrm{Link}(v)},$
and thus the lemma is proved. ∎
Let $Q\subseteq M_{\Gamma}$ be an amenable von Neumann subalgebra. As explained in [56, p. 228] there exists a conditional expectation $\Psi_{Q}:\mathbb{B}(\mathcal{H}_{\Gamma})\rightarrow Q^{\prime}$ that is proper in the sense that for any $a\in\mathbb{B}(\mathcal{H}_{\Gamma})$ we have that $\Psi_{Q}(a)$ is in the $\sigma$-weak closure of
${\rm Conv}\left\{uau^{\ast}\mid u\in\mathcal{U}(Q)\right\},$
where $\rm Conv$} denotes the convex hull.
Lemma 6.12.
Let $Q\subseteq M_{\Gamma}$ be an amenable von Neumann subalgebra. If there is $\Lambda\subseteq\Gamma$ such that $Q\not\prec_{M_{\Gamma}}M_{\Lambda}$, then $D_{\Lambda}$ is contained in $\ker\Psi_{Q}$.
Proof.
Let $p\in\mathbb{K}(\mathcal{H}^{\prime}(\Lambda))$ be a finite rank projection. We first claim that
$U_{\Lambda}^{\prime}(p\otimes 1)U^{\prime*}_{\Lambda}\in\ker\Psi_{Q}.$
We prove this claim by contradiction so suppose that $d:=\Psi_{Q}(U_{\Lambda}^{\prime}(p\otimes 1)U^{\prime*}_{\Lambda})\not=0$. First observe that for $a\in M_{\Lambda}$ we have
$JaJ=U_{\Lambda}^{\prime}(1\otimes J_{\Lambda}aJ_{\Lambda})U_{\Lambda}^{\prime\ast},$
where $J_{\Lambda}$ is the modular conjugation operator of $M_{\Lambda}$ acting on $\mathcal{H}_{\Lambda}$. It follows in particular that
$(JM_{\Lambda}J)^{\prime}=U_{\Lambda}^{\prime}(\mathbb{B}(\mathcal{H}^{\prime}(\Lambda))\bar{\otimes}M_{\Lambda})U_{\Lambda}^{\prime*}.$
Any $u\in\mathcal{U}(Q)$ commutes with $M_{\Gamma}^{\prime}=JM_{\Gamma}J$ and so certainly it commutes with $JM_{\Lambda}J$. As $\Psi_{Q}$ is proper we find that $d$ as defined above thus commutes with $JM_{\Lambda}J$. Thus $d\in U_{\Lambda}^{\prime}(\mathbb{B}(\mathcal{H}^{\prime}(\Lambda))\bar{\otimes}M_{\Lambda})U^{\prime*}_{\Lambda}$. Let $\rm Tr$} the trace on $\mathbb{B}(\mathcal{H}^{\prime}(\Lambda))$ and let $\Phi_{\Lambda}$ be the center valued trace of $M_{\Lambda}$ onto $\mathcal{Z}(M_{\Lambda})=M_{\Lambda}\cap M_{\Lambda}^{\prime}$. Using again that $\Psi_{Q}$ is proper we find by lower semi-continuity [67, Theorem VII.11.1] that for any normal (necessarily tracial) state $\tau$ on the center $\mathcal{Z}(M_{\Lambda})$ we have
$({\rm Tr}\otimes(\tau\circ\Phi_{\Lambda}))(U_{\Lambda}^{\prime\ast}dU_{\Lambda}^{\prime})\leq({\rm Tr}\otimes(\tau\circ\Phi_{\Lambda}))(p\otimes 1)<\infty.$
Let $e$ be a spectral projection of $d$ corresponding to the interval $\|d\|/2,\|d\|]$. Then
$({\rm Tr}\otimes(\tau\circ\Phi_{\Lambda}))(U_{\Lambda}^{\prime\ast}eU_{\Lambda}^{\prime})\leq 2({\rm Tr}\otimes(\tau\circ\Phi_{\Lambda}))(U_{\Lambda}^{\prime\ast}dU_{\Lambda}^{\prime})<\infty.$
Thus it follows that $({\rm Tr}\otimes\Phi_{\Lambda})(U_{\Lambda}^{\prime\ast}eU_{\Lambda}^{\prime})<\infty$. Then $\mathcal{K}:=e\mathcal{H}_{\Gamma}$ is a $Q$-$M_{\Lambda}$ sub-bimodule of $\mathcal{H}_{\Gamma}$ with $\dim_{M_{\Lambda}}(\mathcal{K})<\infty$ and $\mathcal{H}_{\Gamma}$ is the standard representation Hilbert space of $M_{\Gamma}$. It thus follows from Definition 2.1 (3) that $Q\prec_{M_{\Gamma}}M_{\Lambda}$. This contradicts the assumptions and the claim is proved.
Taking linear spans and closures it thus follows from the previous paragraph that
$U_{\Lambda}^{\prime}(\mathbb{K}(\mathcal{H}^{\prime}(\Lambda))\otimes 1)U^{\prime*}_{\Lambda}\subseteq\ker\Psi_{Q}.$
Using the multiplicative domain of $\Psi_{Q}$ it follows then that
$U_{\Lambda}^{\prime}(\mathbb{K}(\mathcal{H}^{\prime}(\Lambda))\otimes\mathbb{B}(\mathcal{H}_{\Lambda}))U^{\prime*}_{\Lambda}\subseteq\ker\Psi_{Q}.$
This concludes the proof.
Lemma 6.13.
Let $Q\subseteq M_{\Gamma}$ be an amenable von Neumann subalgebra. Assume that for every $v\in\Gamma$ we have $Q\not\prec_{M_{\Gamma}}M_{\textrm{Link}(v)}$. Then we have $C_{\Gamma},JA_{\Gamma}J]\subseteq\ker\Psi_{Q}$.
Proof.
The commutator $C_{\Gamma},JA_{v}J]$ is contained in the closed linear span of the sets
$C_{\Gamma}[C_{w},JA_{v}J]C_{\Gamma},\qquad v,w\in\Gamma.$
We have, as $C_{\Gamma}\subseteq B_{\Gamma}$, by Lemma 6.11 and Lemma 6.9 that
$C_{\Gamma}[C_{w},JA_{v}J]C_{\Gamma}\subseteq B_{\Gamma}D_{\textrm{Link}(v)}B_{\Gamma}\subseteq D_{\textrm{Link}(v)}.$
By Lemma 6.12 we see that $D_{\textrm{Link}(v)},v\in\Gamma$ is contained in the kernel of $\Psi_{Q}$. We thus conclude that $C_{\Gamma},JA_{v}J]$ is contained in $\ker\Psi_{Q}$.
Now $C_{\Gamma},JA_{\Gamma}J]$ is contained in the closed linear span of the sets
$JA_{\Gamma}J[C_{\Gamma},JA_{v}J]JA_{\Gamma}J,\qquad v\in\Gamma.$
Note $JA_{\Gamma}J$ is contained in $M_{\Gamma}^{\prime}$ so certainly in $Q^{\prime}$. As $\Psi_{Q}$ is a $Q^{\prime}$-bimodule map it follows that $JA_{\Gamma}J[C_{\Gamma},JA_{v}J]JA_{\Gamma}J$ is contained in $\ker\Psi_{Q}$. This finishes the proof. ∎
Lemma 6.14.
Let $Q\subseteq M_{\Gamma}$ be an amenable von Neumann subalgebra. Assume that for every $v\in\Gamma$ we have $Q\not\prec_{M_{\Gamma}}M_{\textrm{Link}(v)}$. The map
(6.8) $\begin{split}\Theta:A_{\Gamma}\otimes JA_{\Gamma}J&\to\mathbb{B}(\mathcal{H}_{\Gamma})\\ a\otimes JbJ&\mapsto\Psi_{Q}(aJbJ).\end{split}$
is continuous with respect to the minimal tensor norm.
Proof.
Observe that $\Psi_{Q}$ is a $Q^{\prime}$-bimodule map and we have $JA_{\Gamma}J\subseteq M_{\Gamma}^{\prime}\subseteq Q^{\prime}$. It thus follows from Lemma 6.13 that for $x\in C_{\Gamma}$ and $y\in JA_{\Gamma}J$ we have
$\Psi_{Q}(x)y=\Psi_{Q}(xy)=\Psi_{Q}(yx+[x,y])=\Psi_{Q}(yx)=y\Psi_{Q}(x).$
So $\Psi_{Q}(C_{\Gamma})\subseteq(JA_{\Gamma}J)^{\prime}=M_{\Gamma}$. Now consider the composition of maps, see [10, Theorem 3.3.7 and 3.5.3],
$\widetilde{\Theta}:C_{\Gamma}\otimes_{{\rm max}}JA_{\Gamma}J\rightarrow^{\Psi_{Q}\otimes{\rm Id}}M_{\Gamma}\otimes_{{\rm max}}JA_{\Gamma}J\rightarrow^{m}\mathbb{B}(\mathcal{H}),$
where $m$ is the multiplication map. Note that $C_{\Gamma}$ is nuclear by Lemma 6.7. Thus $C_{\Gamma}\otimes_{{\rm max}}JA_{\Gamma}J=C_{\Gamma}\otimes_{{\rm min}}JA_{\Gamma}J$. Then the restriction of $\widetilde{\Theta}$ to $A_{\Gamma}\otimes_{{\rm min}}JA_{\Gamma}J$ gives the map $\Theta$.
Theorem 6.15.
Let $\Gamma$ be a simple graph. Let $(M_{\Gamma},\tau)=\ast_{v,\Gamma}(M_{v},\tau_{v})$ be a graph product of finite von Neumann algebras $M_{v}$ ($\not$$\mathbb{C}$) that satisfy condition strong (AO) and have separable preduals. Let $Q\subseteq M_{\Gamma}$ be a diffuse von Neumann subalgebra. At least one of the following holds:
  1. (1)
    The relative commutant $Q^{\prime}\cap M_{\Gamma}$ is amenable;
  2. (2)
    There exists a non-empty $\Gamma_{0}\subseteq\Gamma$ such that $\textrm{Link}(\Gamma_{0})\not$$\varnothing$ and $Q\prec_{M_{\Gamma}}M_{\Gamma_{0}}$.
Proof.
We first show we can reduce it to the case that $Q$ is amenable. Indeed, suppose we have proven that every amenable diffuse subalgebra $Q_{0}\subseteq M_{\Gamma}$ satisfies (1) or (2). Let $Q\subseteq M_{\Gamma}$ be an arbitrary diffuse subalgebra. Then by [41, Corollary 4.7] there is an amenable diffuse von Neumann subalgebra $Q_{0}\subseteq Q$ such that for subgraphs $\Lambda\subseteq\Gamma$ we have $Q_{0}\not\prec_{M_{\Gamma}}M_{\Lambda}$ whenever $Q\not\prec_{M_{\Gamma}}M_{\Lambda}$. If $Q$ does not satisy (2), then neither does $Q_{0}$. Hence $Q_{0}$ satisfies (1), so $Q_{0}^{\prime}\cap M_{\Gamma}$ is amenable. Hence also the subalgebra $Q^{\prime}\cap M_{\Gamma}\subseteq Q_{0}^{\prime}\cap M_{\Gamma}$ is amenable, i.e. $Q$ satisfies (1), which shows the reduction.
We now prove the statement with the notation introduced in this section. Assume (2) does not hold and we shall prove (1). By assumption for $\Lambda\subseteq\Gamma$ with $\textrm{Link}(\Lambda)\not$$\varnothing$ we have $Q\not\prec_{M_{\Gamma}}M_{\Lambda}$. In particular we have for all $v\in\Gamma$ with $\textrm{Link}(v)$ non-empty that $v$ is contained in $\textrm{Link}(\textrm{Link}(v))$ and so $Q\not\prec_{M_{\Gamma}}M_{\textrm{Link}(v)}$. If $\textrm{Link}(v)$ is empty then $M_{\textrm{Link}(v)}=\mathbb{C}$ and so $Q\not\prec_{M_{\Gamma}}M_{\textrm{Link}(v)}$ as $Q$ is diffuse. It follows now from Lemma 6.14 that $\Theta$ defined in (6.8) is bounded for the minimal tensor norm.
Each $A_{v}$ is exact being included in the nuclear C-algebra $C_{v}$. Therefore the C-algebra $A_{\Gamma}$ is exact by [14, Corollary 3.17]. Furthermore, the inclusions $A_{\Gamma}\subseteq M_{\Gamma}$ and $JA_{\Gamma}J\subseteq M_{\Gamma}^{\prime}$ are $\sigma$-weakly dense.
The conclusions of the previous two paragraphs show that the assumptions of [52, Lemma 2.1] are satisfied and this lemma concludes that $Q^{\prime}\cap M_{\Gamma}$ is amenable.
We recall the following lemma about relative commutants which we shall use without further reference.
Lemma 6.16 (Lemma 3.5 of [69]).
If $A\subseteq 1_{A}M1_{A},B\subseteq 1_{B}M1_{B}$ are von Neumann subalgebras and $A\prec_{M}B$, then $B^{\prime}\cap 1_{B}M1_{B}\prec_{M}A^{\prime}\cap 1_{A}M1_{A}$.
6.3. Unique rigid graph product decomposition
We will prove our main result Theorem 6.18 which asserts for a graph product $M_{\Gamma}=*_{v,\Gamma}(M_{v},\tau_{v})\in\mathcal{C}_{{\rm Rigid}}$ with $M_{v}\in\mathcal{C}_{{\rm Vertex}}$ that we can retrieve the rigid graph $\Gamma$ and retrieve the vertex von Neumann algebras $M_{v}$ up to stable isomorphism. To prove the result we need the following lemma.
Lemma 6.17.
Let $\Gamma$ be a simple graph. For $v\in\Gamma$, let $M_{v},N_{v}$ be II1-factors and put $M_{\Gamma}=*_{v,\Gamma}(M_{v},\tau_{v})$ and $N_{\Gamma}=*_{v,\Gamma}(N_{v},\widetilde{\tau}_{v})$. Suppose $\iota:N_{\Gamma}\to M_{\Gamma}$ is a *-isomorphism and for $v\in\Gamma$ we have
$\iota(N_{v})\prec_{M_{\Gamma}}M_{v}$ $and$ $M_{v}\prec_{M_{\Gamma}}\iota(N_{v}).$
Then the following holds true:
  1. (1)
    For $v\in\Gamma$ there is a unitary $u_{v}\in M_{\Gamma}$ such that $u_{v}^{*}\iota(N_{\textrm{Star}(v)})u_{v}=M_{\textrm{Star}(v)}$.
  2. (2)
    Let $\Lambda_{0}\subseteq\Lambda\subseteq\Gamma$ be subgraphs such that $\iota(N_{\Lambda})=M_{\Lambda}$. Then $\iota(N_{\Lambda\cup\textrm{Link}_{\Gamma}(\Lambda_{0})})=M_{\Lambda\cup\textrm{Link}_{\Gamma}(\Lambda_{0})}$.
  3. (3)
    Let $P=(v_{1},\ldots,v_{n})$ be a path in $\Gamma$ and denote $\Gamma_{0}:=\bigcup_{i=1}^{n}\textrm{Star}(v_{i})$. If there exist $1\leq j\leq n$ and a subgraph $\Lambda\subseteq\Gamma_{0}$ such that $v_{j}\in\Lambda$ and $\iota(N_{\Lambda})=M_{\Lambda}$, then $\iota(N_{\Gamma_{0}})=M_{\Gamma_{0}}$.
  4. (4)
    Let $\Gamma_{0}$ be a connected component of $\Gamma$. If there is a non-empty subgraph $\Lambda\subseteq\Gamma_{0}$ with $\iota(N_{\Lambda})=M_{\Lambda}$ then $\iota(N_{\Gamma_{0}})=M_{\Gamma_{0}}$.
Proof.
(1) As $\iota(N_{v})\prec_{M_{\Gamma}}M_{v}$ and $\iota(N_{v})\not\prec_{M_{\Gamma}}M_{\varnothing}$ (since $N_{v}$ diffuse), and since $\iota(N_{v})$ and $\iota(N_{v})^{\prime}\cap M_{\Gamma}$ ($\iota(N_{\textrm{Link}(v)})$) are factors, we obtain by Theorem 5.11 a unitary $u_{v}\in M_{\Gamma}$ such that $u_{v}^{*}\iota(N_{v})u_{v}\subseteq M_{\textrm{Star}(v)}$. By assumption $M_{v}\prec_{M_{\Gamma}}\iota(N_{v})$ so that $M_{v}\prec_{M_{\Gamma}}u_{v}^{*}\iota(N_{v})u_{v}$. If $u_{v}^{*}\iota(N_{v})u_{v}\prec_{M_{\Gamma}}M_{\textrm{Link}(v)}$ then $u_{v}^{*}\iota(N_{v})u_{v}\prec_{M_{\Gamma}}^{s}M_{\textrm{Link}(v)}$ by Lemma 2.2 (2), since $M_{v}^{\prime}\cap M_{\Gamma}$ is a factor. Consequently, by Lemma 2.2 (1) we obtain $M_{v}\prec_{M_{\Gamma}}M_{\textrm{Link}(v)}$, which gives a contradiction by Proposition 5.9. We conclude that $u_{v}^{*}\iota(N_{v})u_{v}\not\prec_{M_{\Gamma}}M_{\textrm{Link}(v)}$. Now, since $u_{v}^{*}\iota(N_{v})u_{v}\subseteq M_{\textrm{Star}(v)}$ and $u_{v}^{*}\iota(N_{v})u_{v}\not\prec_{M_{\Gamma}}M_{\textrm{Link}(v)}$ we obtain by Proposition 5.8(1b) that $\operatorname{Nor}_{M_{\Gamma}}(u_{v}^{*}\iota(N_{v})u_{v})\subseteq M_{\textrm{Star}(v)}$, hence $u_{v}^{*}\iota(N_{\textrm{Star}(v)})u_{v}\subseteq M_{\textrm{Star}(v)}$.
By symmetry there is also a unitary $\widetilde{u}_{v}\in M_{\Gamma}$ such that $\widetilde{u}_{v}^{*}M_{\textrm{Star}(v)}\widetilde{u}_{v}\subseteq\iota(N_{\textrm{Star}(v)})$. Hence
(6.9) $u_{v}^{*}\widetilde{u}_{v}^{*}M_{\textrm{Star}(v)}\widetilde{u}_{v}u_{v}\subseteq u_{v}^{*}\iota(N_{\textrm{Star}(v)})u_{v}\subseteq M_{\textrm{Star}(v)}.$
Hence, since $M_{\textrm{Star}(v)}\not\prec_{M_{\Gamma}}M_{\widetilde{\Lambda}}$ for any strict subgraph $\widetilde{\Lambda}\subsetneq\textrm{Star}(v)$ we obtain by Proposition 5.8(1c) that $\widetilde{u}_{v}u_{v}\in M_{\textrm{Star}(v)}$. From this we conclude that the inclusions in (6.9) are in fact equalities so $u_{v}^{*}\iota(N_{\textrm{Star}(v)})u_{v}=M_{\textrm{Star}(v)}$.
(2) Let $\Lambda_{0}\subseteq\Lambda$ be a subgraph. Then $\iota(N_{\Lambda_{0}})\subseteq\iota(N_{\Lambda})=M_{\Lambda}$ and by the assumptions $\iota(N_{\Lambda_{0}})\not\prec_{M_{\Gamma}}M_{\widetilde{\Lambda}}$ for any strict subgraph $\widetilde{\Lambda}\subsetneq\Lambda_{0}$. Hence, by Proposition 5.8(1b) we obtain that $\iota(N_{\textrm{Link}(\Lambda_{0})})\subseteq\operatorname{Nor}_{M_{\Gamma}}(\iota(N_{\Lambda_{0}}))^{\prime\prime}\subseteq M_{\Lambda\cup\textrm{Link}(\Lambda_{0})}$. Thus $\iota(N_{\Lambda\cup\textrm{Link}(\Lambda_{0}})\subseteq M_{\Lambda\cup\textrm{Link}(\Lambda_{0})}$. By symmetry we also obtain that $M_{\Lambda\cup\textrm{Link}(\Lambda_{0})}\subseteq\iota(N_{\Lambda\cup\textrm{Link}(\Lambda_{0})})$ so we get the equality.
(3) As $v_{j}\in\Lambda$ and $\iota(N_{\Lambda})=M_{\Lambda}$, using (2) we obtain that $\iota(N_{\Lambda\cup\textrm{Star}(v_{j})})=\iota(N_{\Lambda\cup\textrm{Link}(v_{j})})=M_{\Lambda\cup\textrm{Link}(v_{j})}=M_{\Lambda\cup\textrm{Star}(v_{j})}$. Now for $1\leq i\leq n$ with |ij|=1|i-j|=1 we have $v_{i}\in\Lambda\cup\textrm{Star}(v_{j})$. Hence, applying (2) again we obtain $\iota(N_{\Lambda\cup\textrm{Star}(v_{j})\cup\textrm{Star}(v_{i})})=M_{\Lambda\cup\textrm{Star}(v_{j})\cup\textrm{Star}(v_{i})}$. Repeating the same argument at most $n$ times we obtain $\iota(N_{\Gamma_{0}})=M_{\Gamma_{0}}$.
(4) Let $P=(v_{1},\ldots,v_{n})$ be a path in $\Gamma$ traversing all vertices in $\Gamma_{0}$. Then $\Gamma_{0}$ is equal to $\bigcup_{i=1}^{n}\textrm{Star}(v_{i})$. Now since $\Lambda\subseteq\Gamma_{0}$ is non-empty, we can choose $1\leq j\leq n$ s.t. $v_{j}\in\Lambda$. Now by (3) we obtain $\iota(N_{\Gamma_{0}})=M_{\Gamma_{0}}$. ∎
Theorem 6.18.
Let $\Gamma$ be a rigid graph. For $v\in\Gamma$, let $M_{v}$ be von Neumann algebras in class $\mathcal{C}_{{\rm Vertex}}$. Let $M_{\Gamma}=\ast_{v,\Gamma}(M_{v},\tau_{v})$. Suppose $M_{\Gamma}=\ast_{w,\Lambda}(N_{w},\tau_{w})$ for another rigid graph $\Lambda$ and other von Neumann algebras $N_{w}\in\mathcal{C}_{{\rm Vertex}}$ for $w\in\Lambda$. Then there is a graph isomorphism $\alpha:\Gamma\rightarrow\Lambda$, and for each $v\in\Gamma$ there is a unitary $u_{v}\in M_{\Gamma}$ and a real number $0<t_{v}<\infty$ such that
(6.10) $M_{\textrm{Star}(v)}=u_{v}^{*}N_{\textrm{Star}(\alpha(v))}u_{v}$ $and$ $M_{v}\simeq N_{\alpha(v)}^{t_{v}}.$
Furthermore, for each vertex $v\in\Gamma$ its connected component $\Gamma_{v}\subseteq\Gamma$ satisfies $M_{\Gamma_{v}}=u_{v}^{*}N_{\alpha(\Gamma_{v})}u_{v}$.
Proof.
First we construct the graph isomorphism $\alpha$. Take $v\in\Gamma$. As the vertex von Neumann algebras are factors we have
$M_{\textrm{Link}(v)}^{\prime}\cap M=M_{\textrm{Link}(\textrm{Link}(v))}=M_{v}.$
In particular $M_{\textrm{Link}(v)}^{\prime}\cap M$ is non-amenable. Therefore Theorem 6.15 implies that there exists $\Lambda_{0}\subseteq\Lambda$ such that $M_{\textrm{Link}(v)}\prec_{N_{\Gamma}}N_{\Lambda_{0}}$ and $\textrm{Link}(\Lambda_{0})\not$$\varnothing$. Thus taking relative commutants (Lemma 6.16) we find that $N_{\textrm{Link}(\Lambda_{0})}\prec_{M_{\Gamma}}M_{v}$.
So we have shown that for every $v\in\Gamma$ there exists a subgraph $\alpha(v)\subseteq\Lambda$ that occurs as the link of a set such that $N_{\alpha(v)}\prec_{M_{\Gamma}}M_{v}$. Conversely, by symmetry, for every $w\in\Lambda$ there exists $\beta(w)\subseteq\Gamma$ that occurs as the link of a set such that $M_{\beta(w)}\prec_{M_{\Gamma}}N_{w}$.
Let again $v\in\Gamma$. Then for any $w\in\alpha(v)$ we have $N_{w}\prec_{M_{\Gamma}}M_{v}$ and consequently as $N_{w}^{\prime}\cap M_{\Gamma}$ is a factor $N_{w}\prec_{M_{\Gamma}}^{s}M_{v}$, see Lemma 2.2(2). Therefore, by transitivity of stable embeddings, i.e. Lemma 2.2(1), we find $M_{\beta(w)}\prec_{M_{\Gamma}}M_{v}$. Hence for any $v^{\prime}\in\beta(w)$ we have $M_{v^{\prime}}\prec_{M_{\Gamma}}M_{v}$. But then by Proposition 5.9 we see that $v^{\prime}=v$. Hence $\beta(w)=v$ for any $w\in\alpha(v)$ and in particular is a singleton set. So we have proved that for $v\in\Gamma$ we have $\beta(\alpha(v)):=\bigcup_{w\in\alpha(v)}\beta(w)=v$ and by symmetry for $w\in\Lambda$ we have $\alpha(\beta(w))=w$. But this can only happen if the values of $\alpha$ and $\beta$ are singletons and $\alpha$ and $\beta$ are inverses of each other.
If $v\in\Gamma$ then we know that $N_{\alpha(v)}\prec_{M_{\Gamma}}M_{v}$ and $M_{v}\prec_{M_{\Gamma}}N_{\alpha(v)}$. Taking relative commutants, using again factoriality of the vertex von Neumann algebras, we find
$M_{\textrm{Link}(v)}\prec_{M_{\Gamma}}N_{\textrm{Link}(\alpha(v))},\qquad N_{\textrm{Link}(\alpha(v))}\prec_{M_{\Gamma}}M_{\textrm{Link}(v)}.$
Now take $v^{\prime}\in\textrm{Link}(v)$ so that the first of these embeddings gives $M_{v^{\prime}}\prec_{M_{\Gamma}}N_{\textrm{Link}(\alpha(v))}$, hence $M_{v^{\prime}}\prec_{M_{\Gamma}}^{s}N_{\textrm{Link}(\alpha(v))}$ by Lemma 2.2(2). Then again by Lemma 2.2(1) we obtain $N_{\alpha(v^{\prime})}\prec_{N_{\Gamma}}N_{\textrm{Link}(\alpha(v))}$. This then implies by Proposition 5.9 that $\alpha(v^{\prime})\in\textrm{Link}(\alpha(v))$. So we conclude that $\alpha$ preserves edges. Similarly $\beta$ preserves edges, and it follows that $\alpha:\Gamma\to\Lambda$ is a graph isomorphism.
Since $\Gamma\simeq\Lambda$ we obtain by Lemma 6.17(1) that for each $v\in\Gamma$ there is a unitary $u_{v}\in M_{\Gamma}$ such that $u_{v}^{*}N_{\textrm{Star}(\alpha(v))}u_{v}=M_{\textrm{Star}(v)}$. Consider the *-isomorphism $\iota_{v}:=\operatorname{Ad}_{u_{v}^{*}}:N_{\Gamma}\to M_{\Gamma}$ which satisfies $\iota_{v}(N_{\textrm{Star}(\alpha(v))})=M_{\textrm{Star}(v)}$. Then by Lemma 6.17(4) we obtain for the connected component $\Gamma_{v}\subseteq\Gamma$ of $v$ that $u_{v}^{*}N_{\Gamma_{v}}u_{v}=\iota_{v}(N_{\Gamma_{v}})=M_{\Gamma_{v}}$.
We show the isomorphism of vertex von Neumann algebras up to amplification. Let $w\in\Gamma$. Since $\iota_{w}(N_{\textrm{Star}(\alpha(w))})=M_{\textrm{Star}(w)}$ and since $\iota_{w}(N_{\textrm{Link}(\alpha(w))})^{\prime}\cap M_{\textrm{Star}(w)}=\iota_{w}(N_{\alpha(w)})$ is non-amenable, we obtain by Theorem 6.15 that $\iota_{w}(N_{\textrm{Link}(\alpha(w))})\prec_{M_{\textrm{Star}(w)}}M_{\Gamma_{1}}$ for some subgraph $\Gamma_{1}\subseteq\textrm{Star}(w)$ with $\textrm{Link}_{\textrm{Star}(w)}(\Gamma_{1})\not$$\varnothing$. Thus, by Lemma 6.16 we obtain $M_{\textrm{Link}(\Gamma_{1})}\prec_{M_{\textrm{Star}(w)}}\iota_{w}(N_{\alpha(w)})$. Let $v\in\textrm{Link}(\Gamma_{1})$ (which is non-empty). Then $M_{v}\prec_{M_{\textrm{Star}(w)}}\iota_{w}(N_{\alpha(w)})$ and, as before, $\iota_{w}(N_{\alpha(w)})\prec_{M_{\Gamma}}^{s}M_{w}$. Hence $M_{v}\prec_{M_{\Gamma}}M_{w}$ and so $v=w$ by Proposition 5.9. Therefore $M_{w}\prec_{M_{\textrm{Star}(w)}}\iota_{w}(N_{\alpha(w)})$. Analogously, we obtain $\iota_{w}(N_{\alpha(w)})\prec_{M_{\textrm{Star}(w)}}M_{w}$.
Since we are dealing with II1-factors these embeddings are also with expectation, i.e. $\iota_{w}(N_{\alpha(w)})\\ \preceq_{M_{\textrm{Star}(w)}}M_{w}$ as in [41, Definition 4.1]. Thus, since $M_{\textrm{Star}(w)}=M_{w}\overline{\otimes}M_{\textrm{Link}(w)}$ we obtain by [41, Lemma 4.13] non-zero projections $p_{w},q_{w}\in M_{\textrm{Star}(w)}$ and a partial isometry $v_{w}\in M_{\textrm{Star}(v)}$ with $v_{w}^{*}v_{w}=p_{w}$ and $v_{w}v_{w}^{*}=q_{w}$ and a subfactor $P_{w}\subseteq q_{w}\iota_{w}(N_{\alpha(w)})q_{w}$ so that
$q_{w}\iota_{w}(N_{\alpha(w)})q_{w}=v_{w}M_{w}v_{w}^{*}\overline{\otimes}P_{w},$ $v_{w}M_{\textrm{Link}(w)}v_{w}^{*}=P_{w}\overline{\otimes}q_{w}\iota_{w}(N_{\textrm{Link}(\alpha(w))})q_{w}.$
Since $N_{w}$ is prime, so is $q_{w}\iota_{w}(N_{\alpha(w)})q_{w}$. Hence, as $v_{w}M_{w}v_{w}^{*}$ is a II1-factor, we obtain that $P_{w}$ is a factor of type In for some $n\in\mathbb{N}$. We conclude that $N_{\alpha(w)}$ is isomorphic to some amplification of $M_{w}$.
We state two corollaries that follow from Theorem 6.18. The following result tells us when a rigid graph product $M_{\Gamma}$ can decompose as graph product over another rigid graph $\Lambda$.
Corollary 6.19.
Let $\Gamma,\Lambda$ be rigid graphs. Let $M_{\Gamma}=*_{v,\Gamma}(M_{v},\tau_{v})$ be the graph product of factors $M_{v}\in\mathcal{C}_{{\rm Vertex}}$. The following are equivalent:
  1. (1)
    We can write $\Gamma=*_{w,\Lambda}\Gamma_{w}$ for some non-empty graphs $\Gamma_{w},w\in\Lambda$;
  2. (2)
    We can write $M_{\Gamma}=*_{w,\Lambda}(M_{w},\tau_{w})$ for some factors $M_{w}\in\mathcal{C}_{{\rm Rigid}},w\in\Lambda$.
Proof.
Suppose we can write $\Gamma=*_{w,\Lambda}\Gamma_{w}$ for non-empty graphs $\Gamma_{w}$ for $w\in\Lambda$. Note that $\Gamma_{w}$ is rigid by Lemma 3.5. Now by Remark 3.4 we have $M_{\Gamma}=*_{w,\Lambda}(M_{w},\tau_{w})$ where $M_{w}:=M_{\Gamma_{w}}\in\mathcal{C}_{{\rm Rigid}}$.
For the other direction, suppose $M_{\Gamma}=*_{w,\Lambda}(M_{w},\tau_{w})$ for some $M_{w}\in\mathcal{C}_{{\rm Rigid}}$ for $w\in\Lambda$. Then there are non-empty, rigid graphs $\Gamma_{w}$ and factors $N_{v}\in\mathcal{C}_{{\rm Vertex}}$ for $v\in\Gamma_{w}$ such that $M_{w}=*_{v,\Gamma_{w}}(N_{v},\tau_{v})$ for $w\in\Gamma$. Hence, by Remark 3.4 we obtain $M_{\Gamma}=*_{v,\Gamma_{\Lambda}}(N_{v},\tau_{v})$. Since $\Gamma_{\Lambda}$ is rigid by Lemma 3.5, we obtain by Theorem 6.18 that $\Gamma\simeq\Gamma_{\Lambda}=*_{w,\Lambda}\Gamma_{w}$. ∎
The following corollary states a unique prime factorization for the class $\mathcal{C}_{{\rm Complete}}$. This result recovers the result of [41] for a slightly smaller class.
Corollary 6.20.
Any von Neumann algebra $M\in\mathcal{C}_{{\rm Complete}}$ can decompose as tensor product
(6.11) $M=M_{1}\overline{\otimes}\cdots\overline{\otimes}M_{m}$
for some $1\leq m<\infty$ and prime factors $M_{1},\ldots,M_{m}\in\mathcal{C}_{{\rm Vertex}}$.
Furthermore, suppose $M\simeq N$ for
(6.12) $N=N_{1}\overline{\otimes}\cdots\overline{\otimes}N_{n},$
where $1\leq n<\infty$, and $N_{1},\ldots,N_{n}\in\mathcal{C}_{{\rm Vertex}}$ are other prime factors. Then $n=m$ and there is a permutation $\alpha$ of $\{1,\ldots,m\}$ such that $M_{i}$ is isomorphic to an amplification of $N_{\alpha(i)}$.
Proof.
Since $M\in\mathcal{C}_{{\rm Complete}}$, there is a non-empty complete graph $\Gamma$ and factors $M_{v}\in\mathcal{C}_{{\rm Vertex}}$ for $v\in\Gamma$ such that $M=*_{v,\Gamma}(M_{v},\tau_{v})$. Hence $M=\overline{\bigotimes}_{v\in\Gamma}M_{v}$ since $\Gamma$ is complete. Moreover, for each $v\in\Gamma$ the factor $M_{v}$ is prime by Remark 6.5(1). This shows (6.11) with $m=|\Gamma|$. Let $\Lambda$ be a complete graph with $n$ vertices. Then $N=\overline{\bigotimes}_{1\leq i\leq n}N_{i}=*_{v,\Lambda}(N_{i},\tau_{i})$. Since $\Gamma$ and $\Lambda$ are rigid we obtain by Theorem 6.18 a graph isomorphism $\alpha:\Gamma\to\Lambda$ such that $M_{i}$ is isomorphic to an amplification of $N_{\alpha(i)}$. In particular, $n=|\Lambda|=|\Gamma|=m$. ∎
7. Classification of strong solidity for graph products
We state the definition of strong solidity. We recall the assumption that inclusions of von Neumann algebras are understood as unital inclusions.
Definition 7.1.
A von Neumann algebra $M$ is called strongly solid if for any diffuse, amenable, von Neumann subalgebra $A\subseteq M$, $\operatorname{Nor}_{M}(A)^{\prime\prime}$ is also amenable.
Remark 7.2.
Note that a tracial von Neumann algebra that is not diffuse must be strongly solid as it contains no diffuse unital subalgebras at all.
In Section 7.1 we characterize strong solidity for graph products $M_{\Gamma}$ of tracial von Neumann algebras $(M_{v},\tau_{v})$. In Section 7.2 we then show that for many concrete cases this makes it possible to verify whether the graph product is strongly solid.
7.1. Strong solidity main result
We proof the main result Theorem 7.7. The overall proof method is similar to [6, Theorem 4.4], where strong solidity was classified for the group von Neumann algebras $\mathcal{L}(\mathcal{W}_{\Gamma})$ of right-angled Coxeter groups. To characterize strong solidity we use the following result concerning amalgamated free products.
Theorem 7.3 (Theorem A of [70]).
Let $(N_{1},\tau_{1})$$(N_{2},\tau_{2})$ be tracial von Neumann algebras with a common von Neumann subalgebra $B\subseteq N_{i}$ satisfying $\tau_{1}|_{B}=\tau_{2}|_{B}$, and denote $N:=N_{1}\ast_{B}N_{2}$ for their amalgamated free product. Let $A\subseteq 1_{A}N1_{A}$ be a von Neumann algebra that is amenable relative to $N_{1}$ or $N_{2}$ inside $N$. Put $P=\operatorname{Nor}_{1_{A}N1_{A}}(A)^{\prime\prime}$. Then at least one of the following is true:
  1. (i)
    $A\prec_{N}B$;
  2. (ii)
    $P\prec_{N}N_{i}$ for some $i=1,2$;
  3. (iii)
    $P$ is amenable relative to $B$ inside $N$.
Furthermore, we use the following results that are rather standard.
Proposition 7.4 (Proposition 4.2. in [6] or Proof of Corollary C in [70]).
Let $N\subseteq M$ be a von Neumann subalgebra and assume $N$ is strongly solid. Let $A\subseteq M$ be a diffuse amenable von Neumann subalgebra and $P=\operatorname{Nor}_{M}(A)^{\prime\prime}$ and $z\in P\cap P^{\prime}$ be a non-zero projection. Assume that $zP\prec_{M}N$. Then $zP$ has an amenable direct summand.
Recall that a von Neumann algebra $M$ is atomic if any projection in $M$ majorizes a minimal projection. If $M$ is atomic it is a direct sum of type I factors. We state the following proposition.
Proposition 7.5.
Let $N=N_{1}\overline{\otimes}N_{2}$ be a tensor product of finite von Neumann algebras $N_{1},N_{2}$. The following statements hold:
  1. (1)
    Suppose $N_{1}$ is non-amenable and diffuse and $N$ is strongly solid. Then $N_{2}$ is atomic;
  2. (2)
    Suppose $N_{1}$ is non-amenable and $N_{2}$ is diffuse. Then $N$ is not strongly solid;
  3. (3)
    Suppose $N_{1}$ is strongly solid and diffuse and $N_{2}$ is atomic. Then $N$ is strongly solid.
Proof.
1) Write $N_{2}=N_{c}\oplus N_{d}$ with $N_{c}$ either 0 or a diffuse von Neumann algebra and $N_{d}$ an atomic von Neumann algebra. Assume $N_{c}\not=0$. Let $A\subseteq N_{c},B\subseteq N_{1}$ be diffuse amenable von Neumann subalgebras. Then $C:=\mathbb{C}1_{N_{1}}\overline{\otimes}A\oplus B\overline{\otimes}\mathbb{C}1_{N_{d}}\subseteq N$ is diffuse and amenable. Furthermore, $\operatorname{Nor}_{N}(C)^{\prime\prime}$ contains $N_{1}\overline{\otimes}A\oplus B\overline{\otimes}\mathbb{C}1_{N_{d}}$ which is non-amenable. This contradicts that $N$ is strongly solid and we conclude that $N_{c}=0$.
2) Take any diffuse amenable subalgebra $A\subseteq N_{2}$, for instance we may take $A$ to be a maximal abelian subalgebra. Then $\mathbb{C}1_{N_{1}}\overline{\otimes}A$ is a diffuse amenable subalgebra of $N$ and $\operatorname{Nor}_{N}(\mathbb{C}1_{N_{1}}\overline{\otimes}A)^{\prime\prime}$ contains $N_{1}\overline{\otimes}A$ which is non-amenable. Hence $N$ is not strongly solid.
3) As $N_{2}$ is atomic we may identify $N_{2}$ with $\bigoplus_{\alpha\in I}\operatorname{Mat}_{n_{\alpha}}(\mathbb{C})$ where $I$ is some index set and $n_{\alpha}\in\mathbb{N}_{\geq 1}$. Let $1_{\alpha}$ be the unit of $\operatorname{Mat}_{n_{\alpha}}(\mathbb{C})$. Let $A\subseteq N_{1}\overline{\otimes}N_{2}$ be a diffuse amenable von Neumann subalgebra. Then $1_{\alpha}A\subseteq N_{1}\otimes\operatorname{Mat}_{n_{\alpha}}(\mathbb{C})$. So $\operatorname{Nor}_{N_{1}\otimes\operatorname{Mat}_{n_{\alpha}}(\mathbb{C})}(1_{\alpha}A)^{\prime\prime}$ is amenable by [40, Proposition 5.2] since $N_{1}$ is strong solid and diffuse. Since $\operatorname{Nor}_{N}(A)^{\prime\prime}=\bigoplus_{\alpha\in I}\operatorname{Nor}_{N_{1}\otimes\operatorname{Mat}_{n_{\alpha}}(\mathbb{C})}(1_{\alpha}A)^{\prime\prime}$ and direct sums of amenable von Neumann algebras are amenable we conclude that $\operatorname{Nor}_{N}(A)^{\prime\prime}$ is amenable. It follows that $N$ is strongly solid. ∎
We classify atomicity for graph products.
Proposition 7.6.
Let $\Gamma$ be a finite simple graph. Let $(M_{\Gamma},\tau_{\Gamma})=\ast_{v,\Gamma}(M_{v},\tau_{v})$ be a graph product of tracial von Neumann algebras over a simple graph $\Gamma$. Then $M_{\Gamma}$ is atomic if and only if $\Gamma$ is complete and each $M_{v}$ is atomic.
Proof.
Any subalgebra of an atomic von Neumann algebra is atomic again. It follows that each $M_{v}$ is atomic. If $\Gamma$ would not be complete then we may pick $v,w\in\Gamma$ not sharing an edge and $(M_{v},\tau_{v})\ast(M_{w},\tau_{w})\subseteq M_{\Gamma}$. However, $(M_{v},\tau_{v})\ast(M_{w},\tau_{w})$ is not atomic by [68]. So $\Gamma$ is complete. Conversely if $\Gamma$ is complete and each $M_{v}$ is atomic then $M=\overline{\bigotimes}_{v\in\Gamma}M_{v}$ is atomic. ∎
We now classify strong solidity for graph products in terms of conditions on subgraphs. These conditions can be verified in most cases (see Proposition 7.6, Proposition 7.8, Proposition 7.9 and Theorem 7.12).
Theorem 7.7.
Let $\Gamma$ be a finite graph and for each $v\in\Gamma$ let $M_{v}$ ($\not$$\mathbb{C}$) be a von Neumann algebra with normal faithful trace $\tau_{v}$. Then $M_{\Gamma}$ is strongly solid if and only if the following conditions are satisfied:
  1. (1)
    For every vertex $v\in\Gamma$ the von Neumann algebra $M_{v}$ is strongly solid;
  2. (2)
    For every subgraph $\Lambda\subseteq\Gamma$ with $M_{\Lambda}$ non-amenable, we have that $M_{\textrm{Link}(\Lambda)}$ is not diffuse;
  3. (3)
    For every subgraph $\Lambda\subseteq\Gamma$ with $M_{\Lambda}$ non-amenable and diffuse, we have moreover that $M_{\textrm{Link}(\Lambda)}$ is atomic.
Proof.
Suppose $M_{\Gamma}$ is strongly solid, we show that conditions (1), (2) and (3) are satisfied. Since strong solidity passes to subalgebras, as follows from its very definition, we obtain that (1) is satisfied. Furthermore, suppose $\Gamma_{0}\subseteq\Gamma$ is a subgraph for which $M_{\Gamma_{0}}$ is non-amenable. We have $M_{\Gamma_{0}\cup\textrm{Link}(\Gamma_{0})}=M_{\Gamma_{0}}\overline{\otimes}M_{\textrm{Link}(\Gamma_{0})}$ which is strongly solid being a von Neumann subalgebra of $M_{\Gamma}$. Hence, Proposition 7.5(2) shows that $M_{\textrm{Link}(\Gamma_{0})}$ cannot be diffuse. This concludes (2). If $M_{\Gamma_{0}}$ is diffuse then Proposition 7.5(1) shows that $M_{\textrm{Link}(\Gamma_{0})}$ is atomic. This concludes (3).
We now show the reverse direction. The proof is based on induction to the number of vertices of the graph. The statement clearly holds when $\Gamma=\varnothing$ since in that case $M_{\Gamma}=\mathbb{C}$ is strongly solid.
Induction. Let $\Gamma$ be a non-empty graph, and assume by induction that Theorem 7.7 is proved for any strictly smaller subgraph of $\Gamma$, i.e. with less vertices. Assume conditions (1), (2) and (3) are satisfied for $\Gamma$. Observe that condition (1), (2) and (3) are then satisfied for all subgraphs of $\Gamma$ as well. Hence by the induction hypothesis we obtain that $M_{\Gamma_{0}}$ is strongly solid for all strict subgraphs $\Gamma_{0}\subsetneq\Gamma$. We shall show that $M_{\Gamma}$ is strongly solid. Let $A\subseteq M$ be diffuse and amenable and denote $P=\operatorname{Nor}_{M}(A)^{\prime\prime}$. We will show that $P$ is amenable.
Suppose there is $v\in\Gamma$ with $\textrm{Star}(v)=\Gamma$. Then we can decompose the graph product as $M_{\Gamma}=M_{v}\overline{\otimes}M_{\Gamma\setminus\{v\}}$. Now $M_{v}$ is strongly solid by condition (1), and $M_{\Gamma\setminus\{v\}}$ is strongly solid by the induction hypothesis as $\Gamma\setminus\{v\}\subsetneq\Gamma$ is a strict subgraph. When both $M_{v}$ and $M_{\Gamma\setminus\{v\}}$ are amenable then $M_{\Gamma}=M_{v}\overline{\otimes}M_{\Gamma\setminus\{v\}}$ is also amenable, and hence $M_{\Gamma}$ is strongly solid. We can thus assume that $M_{v}$ or $M_{\Gamma\setminus\{v\}}$ is non-amenable. If $M_{v}$ is non-amenable we need to separate two cases.
  • If $M_{v}$ is non-amenable and not diffuse then by condition (2) neither $M_{\Gamma\setminus\{v\}}$ is diffuse and hence neither is $M_{\Gamma}=M_{v}\overline{\otimes}M_{\Gamma\setminus\{v\}}$. Then certainly $M_{\Gamma}$ is strongly solid by the absence of (unital) diffuse subalgebras.
  • If $M_{v}$ is non-amenable and diffuse then by condition (3) we obtain that $M_{\textrm{Link}(v)}$ ($M_{\Gamma\setminus\{v\}}$) is atomic, so that by Proposition 7.5(3) we have $M_{\Gamma}=M_{\textrm{Link}(v)}\overline{\otimes}M_{v}$ is strongly solid.
The case when $M_{\Gamma\setminus\{v\}}$ is non-amenable can be treated in the same way by swapping the roles of $M_{v}$ and $M_{\Gamma\setminus\{v\}}$ in the previous argument. We summarize that our proof is complete in case there is $v\in\Gamma$ with $\textrm{Star}(v)=\Gamma$.
Now we assume that for all $v\in\Gamma$ we have $\textrm{Star}(v)\not=\Gamma$. Pick $v\in\Gamma$ and set $\Gamma_{1}:=\textrm{Star}(v)$ and $\Gamma_{2}:=\Gamma\setminus\{v\}$. By (2.1) we can decompose $M_{\Gamma}=M_{\Gamma_{1}}\ast_{M_{\Gamma_{1}\cap\Gamma_{2}}}M_{\Gamma_{2}}$. Moreover, as $\Gamma_{1}$, $\Gamma_{2}$ and $\Gamma_{1}\cap\Gamma_{2}$ are strict subgraphs of $\Gamma$ we obtain by our induction hypothesis that $M_{\Gamma_{1}}$, $M_{\Gamma_{2}}$ and $M_{\Gamma_{1}\cap\Gamma_{2}}$ are strongly solid.
Let $z\in P\cap P^{\prime}$ be a central projection such that $zP$ has no amenable direct summand. Note that $zP\subseteq\operatorname{Nor}_{zM_{\Gamma}z}(zA)^{\prime\prime}$. As $zA$ is amenable, it is amenable relative to $M_{\Gamma_{1}}$ in $M_{\Gamma}$. Therefore by Theorem 7.3 at least one of the following three holds.
  1. (1)
    $zA\prec_{M_{\Gamma}}M_{\Gamma_{1}\cap\Gamma_{2}}$;
  2. (2)
    There is $i\in\{1,2\}$ such that $zP\prec_{M_{\Gamma}}M_{\Gamma_{i}}$;
  3. (3)
    zPzP is amenable relative to $M_{\Gamma_{1}\cap\Gamma_{2}}$ inside $M_{\Gamma}$.
We now analyse each of the cases.
Case (2). In Case (2) we have that Proposition 7.4 together with the induction hypothesis shows that $zP$ has an amenable direct summand in case $z\not=0$. This is a contradiction so we conclude $z=0$ and hence $P$ is amenable.
Case (1). In Case (1), since $zA\prec_{M_{\Gamma}}M_{\Gamma_{1}\cap\Gamma_{2}}$ but $zA\not\prec_{M_{\Gamma}}\mathbb{C}=M_{\varnothing}$, there is a subgraph $\Lambda\subseteq\Gamma_{1}\cap\Gamma_{2}$ such that $zA\prec_{M_{\Gamma}}M_{\Lambda}$ but $zA\not\prec_{M_{\Gamma}}M_{\widetilde{\Lambda}}$ for any strict subgraph $\widetilde{\Lambda}\subseteq\Lambda$. Put $\Lambda_{\operatorname{emb}}:=\Lambda\cup\textrm{Link}(\Lambda)$. Observe that $\Lambda_{\operatorname{emb}}$ contains at least $v$ and $\Lambda$. Furthermore, by Proposition 5.8(2) we obtain that $zP\prec_{M_{\Gamma}}M_{\Lambda_{\operatorname{emb}}}$. If $\Lambda_{\operatorname{emb}}\not=\Gamma$ then $M_{\Lambda_{\operatorname{emb}}}$ is strongly solid by the induction hypothesis. Therefore, in case $z\not=0$ we obtain by Proposition 7.4 that $zP$ has an amenable direct summand, which is a contradiction. Thus $z=0$, and $P$ is amenable. Hence $M_{\Gamma}$ is strongly solid.
We can thus assume that $\Lambda_{\operatorname{emb}}=\Gamma$. Suppose $M_{\Lambda}$ is non-amenable. Again we separate two cases:
  • Assume that $M_{\Lambda}$ is non-amenable and diffuse. Then by condition (3) we have that $M_{\textrm{Link}(\Lambda)}$ is atomic and by Proposition 7.6 we see that $\textrm{Link}(\Lambda)$ must be complete. But as $v\in\textrm{Link}(\Lambda)$ this implies that $\textrm{Link}(\Lambda)\subseteq\textrm{Star}(v)=\Gamma_{1}$ and thus $\Lambda_{\operatorname{emb}}\subseteq\Gamma_{1}$. Therefore $\Lambda_{\operatorname{emb}}$ is a strict subgraph of $\Gamma$, a contradiction. So this case does not occur;
  • Assume that $M_{\Lambda}$ is non-amenable and not diffuse. Then by (2) $M_{\textrm{Link}(\Lambda)}$ is not diffuse either. As $M_{\Gamma}=M_{\Lambda}\overline{\otimes}M_{\textrm{Link}(\Lambda)}$ we find that $M_{\Gamma}$ is not diffuse and thus strongly solid by absence of diffuse (unital) subalgebras.
Next suppose $M_{\textrm{Link}(\Lambda)}$ is non-amenable. Again we separate two cases:
  • Assume that $M_{\textrm{Link}(\Lambda)}$ is non-amenable and diffuse. Then $M_{\textrm{Link}(\textrm{Link}(\Lambda))}=M_{\Lambda}$ is atomic by (3). But then $zA\prec_{M_{\Gamma}}M_{\Lambda}$ with $zA$ diffuse leads to a contradiction;
  • Assume that $M_{\textrm{Link}(\Lambda)}$ is non-amenable and not diffuse. Then by (2) also $M_{\Lambda}$ is not diffuse and so $M_{\Gamma}=M_{\Lambda}\overline{\otimes}M_{\textrm{Link}(\Lambda)}$ is not diffuse and thus strongly solid.
So we are left with the case that $M_{\Lambda}$ and $M_{\textrm{Link}(\Lambda)}$ are amenable. In this case, $M_{\Gamma}=M_{\Lambda_{\operatorname{emb}}}=M_{\Lambda}\overline{\otimes}M_{\textrm{Link}(\Lambda)}$ is amenable and hence $M_{\Gamma}$ is strongly solid.
Remainder of the proof of the main theorem in the situation that Case (1) and Case (2) never occur. We first recall that if we can find a single vertex $v$ as above such that we are in case (1) or (2) then the proof is finished. Otherwise for any vertex $v\in\Gamma$ we are in case (3). So $zP$ is amenable relative to $M_{\textrm{Link}(v)}$ inside $M_{\Gamma}$. As $\bigcap_{v\in\Gamma}\textrm{Link}(v)\subseteq\bigcap_{v\in\Gamma}\Gamma\setminus\{v\}=\varnothing$ we obtain by iteratively using Theorem 5.3 that $zP$ is amenable relative to $\bigcap_{v\in V}M_{\textrm{Link}(v)}=\mathbb{C}$, that is $zP$ is amenable. So $z=0$ and we conclude again that $P$ is amenable. ∎
7.2. Classifying strong solidity in specific cases
We show that in many concrete cases that one can verify whether or not a graph product $M_{\Gamma}$ is strongly solid. Theorem 7.7 tells us how to decide whether $M_{\Gamma}$ is strongly solid. For this we need to know for each vertex $v$ whether or not $M_{v}$ is strongly solid. Furthermore, we need to know for each subgraph $\Lambda\subseteq\Gamma$ whether of not $M_{\Lambda}$ is atomic, diffuse, or non-amenable. We observe that in concrete cases we can verify whether $M_{\Lambda}$ is diffuse, atomic or non-amenable. Indeed, atomicity is classified in Proposition 7.6. Furthermore, amenability was classifed in [16]. Moreover, in [16] diffuseness was classified under the condition that each vertex algebra $M_{v}$ contains a unitary element of trace 0, i.e. a Haar unitary. This in particular applies to the case where $M_{v}$ is either diffuse or a group von Neumann algebra. We state these results here.
Proposition 7.8 (Proposition 6.3 of [16]).
Let $\Gamma$ be a simple graph. For $v\in\Gamma$ let $M_{v}$ ($\not$$\mathbb{C}$) be a von Neumann algebra with normal faithful state $\varphi_{v}$. Then the graph product $M_{\Gamma}=*_{v,\Gamma}(M_{v},\varphi_{v})$ is amenable if and only if the following conditions hold:
  1. (1)
    Each vertex von Neumann algebra $M_{v},v\in\Gamma$ is amenable;
  2. (2)
    If $v\not=w\in\Gamma$ share no edge, then $\dim M_{v}=\dim M_{w}=2$ and $\textrm{Link}(\{v,w\})=\Gamma\setminus\{v,w\}$.
Proposition 7.9 (Theorem E of [16]).
Let $(M_{\Gamma},\tau_{\Gamma})=\ast_{v,\Gamma}(M_{v},\tau_{v})$ be a graph product of tracial von Neumann algebras over a simple graph $\Gamma$. Assume that each $M_{v},v\in\Gamma$ contains a unitary $u_{v}$ with $\tau_{v}(u_{v})=0$. Then $M_{\Gamma}$ is diffuse if either (a) there is $v\in\Gamma$ with $M_{v}$ diffuse; (b) $\Gamma$ is not a complete graph.
In case not every vertex von Neumann algebra contain a unitary of trace 0 the situation becomes more subtle and the analysis becomes significantly more intricate. However, if the vertex von Neumann algebras are 2-dimensional then the results in [15], [33], [64] again yield a classification of diffuseness (and amenability) of graph products.
Definition 7.10.
Suppose that $M_{v,q_{v}},q_{v}\in(0,1]$ is the 2-dimensional Hecke algebra which is the $\ast$-algebra spanned by the unit $1_{v}$ and an element $T_{v,q_{v}}$ satisfying the Hecke relation
$(T_{v,q_{v}}-q_{v}^{\frac{1}{2}})(T_{v,q_{v}}+q_{v}^{-\frac{1}{2}})=0,\quad T_{v,q_{v}}^{\ast}=T_{v,q_{v}}.$
Define the tracial state $\tau_{v}$ by setting $\tau_{v}(T_{v,q_{v}})=0$ and $\tau_{v}(1_{v})=1$. For a simple graph $\Gamma$ and ${\bf q}:=(q_{v})_{v\in\Gamma}\in(0,1]^{\Gamma}$ we let $M_{\Gamma,{\bf q}}=\ast_{v,\Gamma}(M_{v},\tau_{v,q_{v}})$ be the graph product von Neumann algebra which is called the right-angled Hecke von Neumann algebra.
Remark 7.11.
Note that $(M_{v,q_{v}},\tau_{v})$ is isomorphic to $\mathbb{C}^{2}$ with tracial state $\tau_{\alpha}(x\oplus y):=\alpha x+(1-\alpha)y$ with $\alpha=\frac{1}{2}\left(1+\sqrt{1-\frac{4}{p_{v}(q)^{2}+4}}\right)$ where $p_{v}(q):=\frac{q_{v}-1}{\sqrt{q_{v}}}\in(-\infty,0]$. Hence a general 2-dimensional von Neumann algebra with a (necessarily tracial) faithful state is of the form $(M_{v,q_{v}},\tau_{q})$ and Hecke algebras correspond to a general graph product of 2-dimensional von Neumann algebras.
Let $L$ be the graph with 3 points and no edges and $L^{+}$ be the graph with 3 points and 1 edge between two of the points.
Theorem 7.12 (Theorm A of [64], Theorem 6.2 of [15]).
Let $\Gamma$ be a finite simple graph and ${\bf q}:=(q_{v})_{v\in\Gamma}\in(0,1]^{\Gamma}$. Then
  1. (1)
    The Hecke von Neumann algebra $M_{\Gamma,{\bf q}}$ is not diffuse if and only if the sum $\sum_{\mathbf{w}\in\mathcal{W}_{\Gamma}}q_{\mathbf{w}}$, converges where $q_{\mathbf{w}}=q_{w_{1}}\ldots q_{w_{n}}$ and $\mathbf{w}=w_{1}\ldots w_{n}$ reduced;
  2. (2)
    $M_{\Gamma,{\bf q}}$ is non-amenable if and only if $\mathcal{W}_{\Gamma}$ is non-amenable if and only if $L$ or $L^{+}$ is a subgraph of $\Gamma$.
Hence, by Theorem 7.7 and Proposition 7.6 and Theorem 7.12 the classification of strongly solid right-angled Hecke von Neumann algebras with finitely many generators is complete. Partial results toward this classification had been obtained earlier in [11] and [8].
8. Classification of primeness for graph products
We start by recalling the definition of primeness.
Definition 8.1.
A II1-factor $M$ is called prime if it can not factorize as a tensor product $M=M_{1}\overline{\otimes}M_{2}$ with $M_{1},M_{2}$ diffuse.
We study primeness for graph product $M_{\Gamma}=*_{v,\Gamma}(M_{v},\tau_{v})$ of tracial von Neumann algebras $M_{v}$. In Section 8.1 we prove Theorem 8.4 which characterizes primeness for graph products of II1-factors. In Section 8.2 we use this to prove Theorem 8.5 concerning irreducible components in rigid graph products. Moreover, we prove Theorem 8.6 which establishes UPF-results for the class $\mathcal{C}_{{\rm Rigid}}$. Last, in Section 8.3 we extend the primeness characterization from Theorem 8.4 to a larger class of graph products.
8.1. Primeness results for graph products of II1-factors
We prove Lemma 8.2 which we use in Lemma 8.3 to give sufficient conditions for a graph product to be either prime or amenable. For graph products of II1-factors we then characterize primeness in Theorem 8.4.
Lemma 8.2.
Let $\Gamma$ be a finite simple graph that is irreducible. For $v\in\Gamma$ let $M_{v}$ ($\not$$\mathbb{C}$) be a von Neumann algebra with a normal faithful trace $\tau_{v}$. Suppose $N\subseteq M_{\Gamma}$ is a diffuse von Neumann subalgebra. The following are equivalent:
  1. (1)
    $N\not\prec_{M_{\Gamma}}M_{\Gamma_{0}}$ for any strict subgraph $\Gamma_{0}\subsetneq\Gamma$;
  2. (2)
    $\operatorname{Nor}_{M_{\Gamma}}(N)^{\prime\prime}\not\prec_{M_{\Gamma}}M_{\Gamma_{0}}$ for any strict subgraph $\Gamma_{0}\subsetneq\Gamma$.
Proof.
As $N\subseteq\operatorname{Nor}_{M_{\Gamma}}(N)^{\prime\prime}$, it is clear that $(1)\implies(2)$. We will show that $(2)\implies(1)$.
As $N\subseteq M_{\Gamma}$ is a subalgebra, we have that $N\prec_{M_{\Gamma}}M_{\Gamma}$. Therefore, there is a (minimal) subgraph $\Lambda\subseteq\Gamma$ such that $N\prec_{M_{\Gamma}}M_{\Lambda}$ and $N\not\prec_{M_{\Gamma}}M_{\widetilde{\Lambda}}$ for all strict subgraphs $\widetilde{\Lambda}\subsetneq\Lambda$. By Proposition 5.8 (2) we obtain that $\operatorname{Nor}_{M_{\Gamma}}(N)^{\prime\prime}\prec_{M_{\Gamma}}M_{\Lambda_{\operatorname{emb}}}$ where $\Lambda_{\operatorname{emb}}=\Lambda\cup\textrm{Link}(\Lambda)$. Now by our assumption this implies that $\Lambda_{\operatorname{emb}}=\Gamma$. Now, as $\Gamma$ is irreducible and $\Gamma=\Lambda\cup\textrm{Link}(\Lambda)$ we have that $\Lambda$ or $\textrm{Link}(\Lambda)$ is empty. As $N\not\prec_{M_{\Gamma}}\mathbb{C}1_{M_{\Gamma}}$ (since $N$ is diffuse) and $N\prec_{M_{\Gamma}}M_{\Lambda}$ we must have that $\Lambda$ is non-empty, and thus that $\textrm{Link}(\Lambda)$ is empty. Thus $\Lambda=\Gamma$, and this proves the statement. ∎
Lemma 8.3.
Let $\Gamma$ be a finite irreducible graph with $|\Gamma|\geq 2$. For $v\in\Gamma$ let $M_{v}$ ($\not$$\mathbb{C}$) be a von Neumann algebra with a normal faithful trace $\tau_{v}$. Suppose the graph product $M_{\Gamma}=*_{v,\Gamma}(M_{v},\tau_{v})$ is a II1-factor and $M_{\Gamma}\not\prec_{M_{\Gamma}}M_{\Gamma_{0}}$ for any strict subgraph $\Gamma_{0}\subsetneq\Gamma$. Then $M_{\Gamma}$ is prime or amenable.
Proof.
Suppose that $M_{\Gamma}$ is not prime, we show it is amenable. As $M_{\Gamma}$ is not prime, we can write $M_{\Gamma}=N_{1}\overline{\otimes}N_{2}$ with $N_{1},N_{2}$ both diffuse. We observe that $\operatorname{Nor}_{M_{\Gamma}}(N_{1})^{\prime\prime}=M_{\Gamma}$. Therefore, using our assumption on $M_{\Gamma}$ and applying Lemma 8.2 we obtain that $N_{1}\not\prec_{M_{\Gamma}}M_{\Gamma_{0}}$ for any strict subgraph $\Gamma_{0}\subsetneq\Gamma$.
As $N_{2}$ is diffuse it contains a diffuse amenable von Neumann subalgebra $A\subseteq N_{2}$. Now observe that $\operatorname{Nor}_{M_{\Gamma}}(A)^{\prime\prime}$ contains $N_{1}$ and hence $\operatorname{Nor}_{M_{\Gamma}}(A)^{\prime\prime}\not\prec_{M_{\Gamma}}M_{\Gamma_{0}}$ for any strict subgraph $\Gamma_{0}\subsetneq\Gamma$. Thus, again by Lemma 8.2 we obtain that $A\not\prec_{M_{\Gamma}}M_{\Gamma_{0}}$ for any strict subgraph $\Gamma_{0}\subsetneq\Gamma.$
Let $v\in\Gamma$ and put $\Gamma_{1}:=\textrm{Star}(v)$ and $\Gamma_{2}:=\Gamma\setminus\{v\}$. We can write
(8.1) $M_{\Gamma}=M_{\Gamma_{1}}*_{M_{\textrm{Link}(v)}}M_{\Gamma_{2}}.$
As $A$ is amenable relative to $M_{\Gamma_{1}}$ inside $M_{\Gamma}$ (as $A$ is amenable), we obtain using Theorem 7.3 that at least one of the following holds:
  1. (1)
    $A\prec_{M_{\Gamma}}M_{\textrm{Link}(v)}$;
  2. (2)
    $\operatorname{Nor}_{M_{\Gamma}}(A)^{\prime\prime}\prec_{M_{\Gamma}}M_{\Gamma_{i}}$ for some $i\in\{1,2\}$;
  3. (3)
    $\operatorname{Nor}_{M_{\Gamma}}(A)^{\prime\prime}$ is amenable relative to $M_{\textrm{Link}(v)}$ inside $M_{\Gamma}$.
Now as $\Gamma_{1},\Gamma_{2}$ and $\textrm{Link}(v)$ are strict subgraphs of $\Gamma$ (as $\Gamma$ is irreducible and $|\Gamma|\geq 2$), we obtain that only option (3) is possible. Thus $\operatorname{Nor}_{M_{\Gamma}}(A)^{\prime\prime}$ is amenable relative to $M_{\textrm{Link}(v)}$ inside $M_{\Gamma}$. Note that $v\in\Gamma$ was chosen arbitarily. Thus, applying Theorem 5.3 repeatedly, and using that $\bigcap_{v\in\Gamma}\textrm{Link}(v)=\varnothing$, we obtain that $\operatorname{Nor}_{M_{\Gamma}}(A)^{\prime\prime}$ is amenable relative to $M_{\varnothing}$ ($\mathbb{C}1_{M_{\Gamma}}$) inside $M_{\Gamma}$, i.e. $\operatorname{Nor}_{M_{\Gamma}}(A)^{\prime\prime}$ is amenable. Hence the subalgebra $N_{1}\subseteq\operatorname{Nor}_{M_{\Gamma}}(A)^{\prime\prime}$ is amenable as well. Interchanging the roles of $N_{1}$ and $N_{2}$ we obtain that $N_{2}$ is also amenable, and hence $M_{\Gamma}=N_{1}\overline{\otimes}N_{2}$ is amenable. ∎
We characterize primeness for graph products of II1-factors.
Theorem 8.4.
Let $\Gamma$ be a finite simple graph of size $|\Gamma|\geq 2$. For each $v\in\Gamma$ let $M_{v}$ be a II1-factor. Then the graph product $M_{\Gamma}=*_{v,\Gamma}(M_{v},\tau_{v})$ is prime if and only if $\Gamma$ is irreducible.
Proof.
Take the finite simple graph $\Gamma$ with $|\Gamma|\geq 2$ and the II1-factors $(M_{v},\tau_{v})$ for $v\in\Gamma$. By [14, Theorem 1.2] the von Neumann algebra $M_{\Gamma}$ is a factor. Furthermore, by Proposition 5.9 we have that $M_{\Gamma}\not\prec_{M_{\Gamma}}M_{\Gamma_{0}}$ for any strict subgraph $\Gamma_{0}\subsetneq\Gamma$. Suppose that $\Gamma$ is irreducible. Then by applying Lemma 8.3 we obtain that $M_{\Gamma}$ is prime or amenable. Since $\Gamma$ is irreducible and has size $|\Gamma|\geq 2$ we obtain that $\Gamma$ is not complete. We then see by Proposition 7.8 that $M_{\Gamma}$ is non-amenable. Thus $M_{\Gamma}$ is prime, which shows one direction. Now suppose $\Gamma$ is reducible, so that we can decompose $\Gamma=\Gamma_{1}\cup\Gamma_{2}$ with $\Gamma_{1},\Gamma_{2}\subseteq\Gamma$ non-empty and such that $\textrm{Link}(\Gamma_{1})=\Gamma_{2}$. But then we can decompose $M_{\Gamma}=M_{\Gamma_{1}}\overline{\otimes}M_{\Gamma_{2}}$ as a tensor product and again by [14, Theorem 1.2] $M_{\Gamma_{1}}$ and $M_{\Gamma_{2}}$ are II1-factors. Thus $M_{\Gamma}$ is not prime. ∎
8.2. Unique prime factorization results
We proof Theorem 8.5 which strengthens the statement of Theorem 6.18 by showing for irreducible components $\Gamma_{0}$ that $M_{\Gamma_{0}}$ is isomorphic to an amplification of $N_{\alpha(\Gamma_{0})}$. We then use this result to proof Theorem 8.6 which establishes UPF results for the class $\mathcal{C}_{{\rm Rigid}}$.
Theorem 8.5.
Given a finite rigid graph $\Gamma$. For each $v\in\Gamma$ let $M_{v}\in\mathcal{C}_{{\rm Vertex}}$. Let $M_{\Gamma}=*_{v,\Gamma}(M_{v},\tau_{v})$ be the graph product. Suppose $M_{\Gamma}=*_{w,\Lambda}(N_{w},\tau_{w})$, with another rigid graph $\Lambda$ and other von Neumann algerbas $N_{w}\in\mathcal{C}_{{\rm Vertex}}$. Let $\alpha:\Gamma\to\Lambda$ be the graph isomorphism from Theorem 6.18. Then for any irreducible component $\Gamma_{0}\subseteq\Gamma$, $M_{\Gamma_{0}}$ is isomorphic to an amplification of $N_{\alpha(\Gamma_{0})}$.
Proof.
We observe that $M_{\Gamma\setminus\Gamma_{0}}^{\prime}\cap M_{\Gamma}=M_{\Gamma_{0}}$ is non-amenable. Hence, by Theorem 6.15 we obtain a subgraph $\Lambda_{0}\subseteq\Lambda$ such that $M_{\Gamma\setminus\Gamma_{0}}\prec_{M_{\Gamma}}N_{\Lambda_{0}}$ and $\textrm{Link}_{\Lambda}(\Lambda_{0})\not$$\varnothing$. Choose $\widetilde{\Lambda}_{0}\subseteq\Lambda_{0}$ minimal with the property that $M_{\Gamma\setminus\Gamma_{0}}\prec_{M_{\Gamma}}N_{\widetilde{\Lambda}_{0}}$. We show $\widetilde{\Lambda}_{0}=\alpha(\Gamma\setminus\Gamma_{0})$. By Proposition 5.8(2) we have $N_{\Lambda}=M_{\Gamma}=\operatorname{Nor}_{M_{\Gamma}}(M_{\Gamma\setminus\Gamma_{0}})^{\prime\prime}\prec_{M_{\Gamma}}N_{\Lambda_{\operatorname{emb}}}$ where $\Lambda_{\operatorname{emb}}=\widetilde{\Lambda}_{0}\cup\textrm{Link}_{\Lambda}(\widetilde{\Lambda}_{0})$. By Proposition 5.9 we conclude $\Lambda_{\operatorname{emb}}=\Lambda$. We note for $v\in\Gamma\setminus\Gamma_{0}$ that $N_{\alpha(v)}\prec_{M_{\Gamma}}M_{\Gamma\setminus\Gamma_{0}}$ and $M_{\Gamma\setminus\Gamma_{0}}\prec_{M_{\Gamma}}^{s}N_{\widetilde{\Lambda}_{0}}$ by Lemma 2.2(2). Hence by Lemma 2.2(1) we obtain $N_{\alpha(v)}\prec_{M_{\Gamma}}N_{\widetilde{\Lambda}_{0}}$. Thus $\alpha(\Gamma\setminus\Gamma_{0})\subseteq\widetilde{\Lambda}_{0}$ by Proposition 5.9. Put $S=\widetilde{\Lambda}_{0}\cap\alpha(\Gamma_{0})$. Then
$S\cup\textrm{Link}_{\alpha(\Gamma_{0})}(S)=(\widetilde{\Lambda}_{0}\cup\textrm{Link}_{\Lambda}(S))\cap\alpha(\Gamma_{0})\supseteq(\widetilde{\Lambda_{0}}\cup\textrm{Link}_{\Lambda}(\widetilde{\Lambda}_{0}))\cap\alpha(\Gamma_{0})=\alpha(\Gamma_{0}).$
Since the graph $\alpha(\Gamma_{0})$ is irreducible, we conclude that $S=\varnothing$ or $S=\alpha(\Gamma_{0})$. Now, if $S=\alpha(\Gamma_{0})$ then $\alpha(\Gamma_{0})\subseteq\widetilde{\Lambda}_{0}$, so that $\Lambda=\alpha(\Gamma_{0})\cup\alpha(\Gamma\setminus\Gamma_{0})\subseteq\widetilde{\Lambda}_{0}$ . But since $\widetilde{\Lambda}_{0}\subseteq\Lambda_{0}\subseteq\Lambda$ this implies $\Lambda_{0}=\Lambda$, which contradicts the fact that $\textrm{Link}_{\Lambda}(\Lambda_{0})\not$$\varnothing$. We conclude that $S=\varnothing$ and thus $\widetilde{\Lambda}_{0}=\alpha(\Gamma\setminus\Gamma_{0})$.
We have obtained $M_{\Gamma\setminus\Gamma_{0}}\prec_{M_{\Gamma_{0}}}N_{\alpha(\Gamma\setminus\Gamma_{0})}$. Taking relative commutants, by Lemma 6.16, we get $N_{\alpha(\Gamma_{0})}\prec_{M_{\Gamma}}M_{\Gamma_{0}}$. Since we are dealing with II1-factors, these embeddings are also with expectation, i.e. $N_{\alpha(\Gamma_{0})}\preceq_{M_{\Gamma}}M_{\Gamma_{0}}$ as in [41, Definition 4.1]. Thus, since $M_{\Gamma}=M_{\Gamma_{0}}\overline{\otimes}M_{\Gamma\setminus\Gamma_{0}}$ we obtain by [41, Lemma 4.13] non-zero projections $p,q\in M_{\Gamma}$ and a partial isometry $v\in M_{\Gamma}$ with $v^{*}v=p$ and $vv^{*}=q$ and a subfactor $P\subseteq qN_{\alpha(\Gamma_{0})}q$ so that
$qN_{\alpha(\Gamma_{0})}q=vM_{\Gamma_{0}}v^{*}\overline{\otimes}P$ $and$ $vM_{\Gamma\setminus\Gamma_{0}}v^{*}=P\overline{\otimes}qN_{\alpha(\Gamma\setminus\Gamma_{0})}q.$
By Theorem 8.4 we have that $N_{\alpha(\Gamma_{0})}$ is prime. Hence $qN_{\alpha(\Gamma_{0})}q$ is prime. Thus, since $vM_{\Gamma_{0}}v^{*}$ is a II1-factor, we obtain that $P$ is a type In factor for some $n\in\mathbb{N}$. We conclude that $N_{\alpha(\Gamma_{0})}$ is isomorphic to some amplification of $M_{\Gamma_{0}}$. ∎
Theorem 8.6.
Any von Neumann algebra $M\in\mathcal{C}_{{\rm Rigid}}^{f}$ have a prime factorization inside $\mathcal{C}_{{\rm Rigid}}^{f}$, i.e.
(8.2) $M=M_{1}\overline{\otimes}\cdots\overline{\otimes}M_{m},$
for some $1\leq m<\infty$ and prime factors $M_{1},\ldots,M_{m}\in\mathcal{C}_{{\rm Rigid}}^{f}$.
Suppose there is another prime factorization of $M$ inside $\mathcal{C}_{{\rm Rigid}}^{f}$, i.e.
(8.3) $M=N_{1}\overline{\otimes}\cdots\overline{\otimes}N_{n},$
for another $1\leq n<\infty$ and other prime factors $N_{1},\ldots,N_{n}\in\mathcal{C}_{{\rm Rigid}}^{f}$. Then $m=n$ and there is a permutation $\sigma$ of $\{1,\ldots,m\}$ such that $M_{i}$ is isomorphic to some amplification of $N_{\sigma(i)}$.
Proof.
Since $M\in\mathcal{C}_{{\rm Rigid}}^{f}$, we can write $M=*_{v,\Gamma}(M_{v},\tau_{v})$ for some finite rigid graph $\Gamma$ and some $M_{v}\in\mathcal{C}_{{\rm Vertex}}$ for $v\in\Gamma$. Let $\Gamma_{1},\ldots,\Gamma_{m}(1\leq m<\infty)$ be the irreducible components of $\Gamma$. Let $\Pi=\{1,\ldots,m\}$ be the complete graph with $m$ vertices and put $M_{i}=M_{\Gamma_{i}}$ for $i\in\Pi$. Then since $\Gamma=\Gamma_{\Pi}$ we have by Remark 3.4 that $M=*_{v,\Gamma}(M_{v},\tau_{v})=*_{i,\Pi}(*_{v,\Gamma_{i}}(M_{v},\tau_{v}))=M_{1}\overline{\otimes}\cdots\overline{\otimes}M_{m}$. Now, for $i\in\Pi$ we have by Theorem 8.4 that $M_{i}$ is prime since $\Gamma_{i}$ is irreducible. Note furthermore that $\Gamma_{i}$ is rigid by Remark 3.6 and hence $M_{i}\in\mathcal{C}_{{\rm Rigid}}^{f}$.
Now since $N_{i}\in\mathcal{C}_{{\rm Rigid}}^{f}$ for $i\in\{1,2,\ldots,n\}$, we can write $N_{i}=*_{v,\Lambda_{i}}(N_{v},\tau_{v})$ for some non-empty, rigid graph $\Lambda_{i}$. We note that $\Lambda_{i}$ is irreducible by Theorem 8.4 since $N_{i}$ is prime. Let $\Pi^{\prime}=\{1,\ldots,n\}$ be a complete graph with $n$ vertices and put $\Lambda:=\Lambda_{\Pi}$ which is rigid by Lemma 3.5. Then by Remark 3.4 we have $M=N_{1}\overline{\otimes}\cdots\overline{\otimes}N_{n}=*_{i,\Pi}(N_{i},\tau_{i})=*_{i,\Pi}(*_{v,\Lambda_{i}}(N_{v},\tau_{v}))=*_{v,\Lambda}(N_{v},\tau_{v})$. Hence, we can apply Theorem 6.18 to obtain a graph isomorphism $\alpha:\Gamma\to\Lambda$. We note that $\Lambda_{1},\ldots,\Lambda_{n}$ are the irreducible components of $\Lambda$ and that $\Gamma_{1},\ldots,\Gamma_{m}$ are the irreducible components of $\Gamma$. Since $\alpha$ is a graph isomorphism, this implies that $m=n$ and that there is a permutation $\sigma$ of $\{1,\ldots,m\}$ such that $\alpha(\Gamma_{i})=\Lambda_{\sigma(i)}$. Now, for $1\leq i<m$ we obtain by Theorem 6.18 a real number $0<t_{i}<\infty$ such that $M_{i}=M_{\Gamma_{i}}\simeq N_{\alpha(\Gamma_{i})}^{t_{i}}=N_{\Lambda_{\sigma(i)}}^{t_{i}}=N_{\sigma(i)}^{t_{i}}$. ∎
Remark 8.7.
In Fig. 1 we give an example of a von Neumann algebra for which we obtain a unique prime factorization. This example was not yet covered by [41, Theorem A] since the graph $\Gamma$ is not complete. The example is also not covered by [19, Theorem 6.16] in case the vertex von Neumann algebras $M_{v}\in\mathcal{C}_{{\rm Vertex}}$ are not known to be group von Neumann algebras. Examples of such $M_{v}$ can be found as von Neumann algebras of free orthogonal quantum groups [73] or $q$-Gaussian algebras of finite dimensional Hilbert spaces and $q\in(-1,1)$ sufficiently far away from 0, see [7, Remark 4.5] which is essentially proved in [49]. We emphasize that it is not known whether such von Neumann algebras are group von Neumann algebras; we do not make the more definite claim that they cannot be isomorphic to group von Neumann algebras.
aabbccddeeffgghhiijj
Figure 1. An example of a rigid graph $\Gamma$ is depicted. Put $M_{v}\in\mathcal{C}_{{\rm Vertex}}$ for $v\in\Gamma$. Then Theorem 8.6 obtains for $M_{\Gamma}=*_{v,\Gamma}(M_{v},\tau_{v})$ the unique prime factorization $M_{\Gamma}=M_{\Gamma_{1}}\overline{\otimes}M_{\Gamma_{2}}$, where $\Gamma_{1}=\{a,b,c,d,e\}$ and $\Gamma_{2}=\{f,g,h,i,j\}$ are the irreducible components of $\Gamma$.
8.3. Primeness results for other graph products
In case the von Neumann algebras $M_{v}$ are not (all) type II1-factors, it is interesting to know whether the condition $M_{\Gamma}\not\prec_{M_{\Gamma}}M_{\Gamma_{0}}$ for any strict subgraph $\Gamma_{0}\subsetneq\Gamma$, is satisfied. In Lemma 8.10 we will give sufficient conditions for the property to hold. To prove this, we need the following lemma.
Lemma 8.8.
Let $\Gamma$ be a simple graph and for $v\in\Gamma$ let $(M_{v},\tau_{v})$ be a tracial von Neumann algebra. Let $\Lambda\subseteq\Gamma$ be a subgraph and let $\mathbf{u}\in\mathcal{W}_{\Gamma}\setminus\mathcal{W}_{\Lambda}$. Let $\mathbf{v},\mathbf{v}^{\prime}\in\mathcal{W}_{\Gamma}$ be such that every letters at the start of $\mathbf{v}$ respectively $\mathbf{v}^{\prime}$ does not commute with any letters at the end of $\mathbf{u}^{-1}$ respectively $\mathbf{u}$. Let $\mathbf{w},\mathbf{w}^{\prime}\in\mathcal{W}_{\Gamma}$ with $|\mathbf{w}|\leq|\mathbf{v}|$ and $|\mathbf{w}^{\prime}|\leq|\mathbf{v}^{\prime}|$. Then
(8.4) $\mathbb{E}_{M_{\Lambda}}(axb)=0$  $\quad\quad\text{ for }a\in\mathring{M}_{\mathbf{w}},x\in\mathring{M}_{\mathbf{v}^{-1}\mathbf{u}\mathbf{v}^{\prime}},b\in\mathring{M}_{\mathbf{w}^{\prime}}.$
Proof.
Let $\mathbf{u},\mathbf{v},\mathbf{v}^{\prime},\mathbf{w},\mathbf{w}^{\prime}$ be given as stated. Observe by the assumptions on $\mathbf{v}$ and $\mathbf{v}^{\prime}$ that in particular $\mathbf{v}^{-1}\mathbf{u}\mathbf{v}^{\prime}$ is reduced. Denote
(8.5) $\mathcal{H}(\mathbf{u})$ $:=\bigoplus_{\mathbf{w}_{0}\in\mathcal{W}(\mathbf{u})}\mathring{\mathcal{H}}_{\mathbf{w}_{0}},$ $\mathbf{M}(\mathbf{u})$ $:=\bigoplus_{\mathbf{w}_{0}\in\mathcal{W}(\mathbf{u})}\mathring{\mathbf{M}}_{\mathbf{w}_{0}}.$
Observe for $y_{1}\in\lambda(\mathbf{M}(\mathbf{u}^{-1}))$, $y_{2}\in\mathring{M}_{\mathbf{u}}$ and $y_{3}\in\lambda(\mathbf{M}(\mathbf{u}))$ that if we denote $y:=y_{1}^{*}y_{2}y_{3}$ and write $y=\sum_{\mathbf{w}\in\mathcal{W}_{\Gamma}}y_{\mathbf{w}}$ where $y_{\mathbf{w}}\in\mathring{M}_{\mathbf{w}}$, then we have that $y_{\mathbf{w}}=0$ whenever $\mathbf{w}$ does not contain $\mathbf{u}$ as a subword. Thus, in particular $\mathbb{E}_{M_{\Lambda}}(y_{1}^{*}y_{2}y_{3})=0$. We will apply this to obtain the result.
[…]
By Lemma 2.5 we have $\eta_{\omega}:=\lambda_{(\mathbf{v}_{1}^{\prime},\mathbf{v}_{2}^{\prime},\mathbf{v}_{3}^{\prime})}(x_{3})b\Omega\in\mathring{\mathcal{H}}_{\mathbf{v}_{0}^{\prime}}$ where $\mathbf{v}_{0}^{\prime}=\mathbf{v}_{1}^{\prime}\mathbf{v}_{3}^{\prime}\mathbf{w}^{\prime}$. We show that $\eta_{\omega}\in\mathcal{H}(\mathbf{u})$. In particular, we can assume that $\eta_{\omega}$ is non-zero, so that $\mathbf{w}^{\prime}$ starts with $(\mathbf{v}_{3}^{\prime})^{-1}\mathbf{v}_{2}^{\prime}$ and $\mathbf{v}_{0}^{\prime}$ starts with $\mathbf{v}_{1}^{\prime}\mathbf{v}_{2}^{\prime}$. If $\mathbf{v}_{1}^{\prime}\mathbf{v}_{2}^{\prime}=e$ then $\mathbf{v}_{3}^{\prime}=\mathbf{v}^{\prime}$, so that $|\mathbf{v}_{3}^{\prime}|+|\mathbf{v}_{3}^{\prime}\mathbf{w}^{\prime}|=|\mathbf{w}^{\prime}|\leq|\mathbf{v}^{\prime}|=|\mathbf{v}_{3}^{\prime}|$ and therefore $\mathbf{v}_{3}^{\prime}\mathbf{w}^{\prime}=e$. We then conclude that $\eta_{\omega}\in\mathring{\mathcal{H}}_{e}\subseteq\mathcal{H}(\mathbf{u})$. Thus, suppose $\mathbf{v}_{1}^{\prime}\mathbf{v}_{2}^{\prime}\not=e$. Then $\mathbf{v}_{1}^{\prime}\mathbf{v}_{2}^{\prime}\mathbf{w}_{0}^{\prime}$ ($\mathbf{v}_{0}^{\prime}$) starts with a letter $v_{0}^{\prime}$ at the start of $\mathbf{v}^{\prime}$. Now, by the assumption on $\mathbf{v}^{\prime}$ we obtain that $v_{0}^{\prime}$ does not commute with elements at the end of $\mathbf{u}$. This implies that $\mathbf{u}\mathbf{v}_{0}^{\prime}$ is reduced and so $\eta_{\omega}\in\mathcal{H}(\mathbf{u})$. Now, as $\lambda(x_{3})\lambda(b)\Omega=\sum_{\omega\in\mathcal{S}_{\mathbf{v}^{\prime}}}\lambda_{\omega}(x_{3})\lambda(b)\Omega\in\mathcal{H}(\mathbf{u})$ we obtain that $y_{3}:=\lambda(x_{3})\lambda(b)\in\mathbf{M}(\mathbf{u})$. In a similar way we obtain $y_{1}:=\lambda(x_{1})\lambda(a)^{*}\in\mathbf{M}(\mathbf{u}^{-1})$. Hence, putting $y_{2}:=\lambda(x_{2})$ we obtain that $\mathbb{E}_{M_{\Lambda}}(\lambda(a)\lambda(x)\lambda(b))=\mathbb{E}_{M_{\Lambda}}(y_{1}^{*}y_{2}y_{3})=0$. The result now follows by density of $\lambda(\mathring{\mathbf{M}}_{\mathbf{z}})\subseteq\mathring{M}_{\mathbf{z}}$ for $\mathbf{z}\in\mathcal{W}_{\Gamma}$. ∎
Corollary 8.9.
Let $\Gamma$ be a simple graph, $\Lambda\subseteq\Gamma$ be a subgraph and let $\mathbf{u}\in\mathcal{W}_{\Gamma}\setminus\mathcal{W}_{\Lambda}$. Let $\mathbf{v},\mathbf{v}^{\prime}\in\mathcal{W}_{\Gamma}$ be such that every letters at the start of $\mathbf{v}$ respectively $\mathbf{v}^{\prime}$ does not commute with any letters at the end of $\mathbf{u}^{-1}$ respectively $\mathbf{u}$. Let $\mathbf{w},\mathbf{w}^{\prime}\in\mathcal{W}_{\Gamma}$ with $|\mathbf{w}|\leq|\mathbf{v}|$ and $|\mathbf{w}^{\prime}|\leq|\mathbf{v}^{\prime}|$. Then $\mathbf{w}\mathbf{v}^{-1}\mathbf{u}\mathbf{v}^{\prime}\mathbf{w}^{\prime}\not\in\mathcal{W}_{\Lambda}$.
Proof.
For $v\in\Gamma$ let $M_{v}:=\mathcal{L}(\mathbb{Z}/2\mathbb{Z})$, so that $M_{\Gamma}=\mathcal{L}(\mathcal{W}_{\Gamma})$. Take $a=\lambda_{\mathbf{w}}$, $x=\lambda_{\mathbf{v}^{-1}\mathbf{u}\mathbf{v}^{\prime}}$ and $b=\lambda_{\mathbf{w}^{\prime}}$. Then Lemma 8.8 shows that $\mathbb{E}_{M_{\Gamma\setminus\Lambda}}(\lambda_{\mathbf{w}\mathbf{v}^{-1}\mathbf{u}\mathbf{v}^{\prime}\mathbf{w}})=\mathbb{E}_{M_{\Lambda}}(axb)=0$. This means that $\mathbf{w}_{1}\mathbf{v}^{-1}\mathbf{u}\mathbf{v}\mathbf{w}_{2}\not\in\mathcal{W}_{\Lambda}$. ∎
Lemma 8.10.
Let $\Gamma$ be a simple graph of size $|\Gamma|\geq 3$ such that for any $v\in\Gamma$, $Star(v)\not=\Gamma$. For $v\in\Gamma$ let $(M_{v},\tau_{v})$ be a von Neumann algebra with a normal faithful trace. Suppose for any $v\in\Gamma$ there is a unitary $u_{v}\in M_{\Gamma}$ with $\tau_{v}(u_{v})=0$. Then $M_{\Gamma}\not\prec_{M_{\Gamma}}M_{\Lambda}$ for any strict subgraph $\Lambda\subsetneq\Gamma$.
Proof.
First observe that the Coxeter group $\mathcal{W}_{\Gamma}$ is icc since $|\Gamma|\geq 3$ and $\textrm{Star}(v)\not=\Gamma$ for all $v\in\Gamma$. Now let $\Lambda\subsetneq\Gamma$ be a strict subgraph and fix $v\in\Gamma\setminus\Lambda$. As the conjugacy class $\{\mathbf{v}^{-1}v\mathbf{v}:\mathbf{v}\in\mathcal{W}_{\Gamma}\}$ is infinite, we can for $n\in\mathbb{N}$ choose $\mathbf{v}_{n}\in\mathcal{W}_{\Gamma}$ such that $|\mathbf{v}_{n}^{-1}v\mathbf{v}_{n}|\geq 2n+1$. If a letters $s$ commuting with $v$ is at the start of $\mathbf{v}_{n}$ then we can replace $\mathbf{v}_{n}$ with $\widetilde{\mathbf{v}}_{n}:=s\mathbf{v}_{n}\in\mathcal{W}_{\Gamma}$ which does not start with $s$ and is such that $\widetilde{\mathbf{v}}_{n}^{-1}v\widetilde{\mathbf{v}}_{n}=\mathbf{v}_{n}^{-1}v\mathbf{v}_{n}$. Repeating the argument, we may thus assume that every letter at the start of $\mathbf{v}_{n}$ does not commute with $v$. Then in particular $\mathbf{v}_{n}^{-1}v\mathbf{v}_{n}$ is reduced and $|\mathbf{v}_{n}|\geq n$. Let $(v_{n,1},\ldots,v_{n,l_{n}})$ be a reduced expression for $\mathbf{v}_{n}^{-1}v\mathbf{v}_{n}$ and define $u_{n}:=u_{v_{n,1}}\ldots u_{v_{n,l_{n}}}\in\mathring{M}_{\mathbf{v}_{n}^{-1}\mathbf{u}\mathbf{v}_{n}}$. Then $u_{n}$ is a unitary and for any $\mathbf{w},\mathbf{w}^{\prime}\in\mathcal{W}_{\Gamma}$ with $|\mathbf{w}|,|\mathbf{w}^{\prime}|\leq n$ and $a\in\mathring{M}_{\mathbf{w}}$, $b\in\mathring{M}_{\mathbf{w}^{\prime}}$, we have by Lemma 8.8 that
(8.6) $\mathbb{E}_{M_{\Lambda}}(au_{n}b)=0.$
We take $x,y\in M_{\Gamma}$ and $\varepsilon>0$. We can choose $x_{0}\in M_{\Gamma}$ of the form $x_{0}=\sum_{i=1}^{K_{1}}x_{i}$ for some $K_{1}\geq 1$, $x_{i}\in\mathring{M}_{\mathbf{w}_{i}}$ with some $\mathbf{w}_{i}\in\mathcal{W}_{\Gamma}$, and with $\|y\|\cdot\|x_{0}-x\|_{2}\leq\varepsilon$. We can now also choose $y_{0}\in M_{\Gamma}$ of the form $y_{0}=\sum_{i=1}^{K_{2}}y_{i}$ for some $K_{2}\geq 1$, $y_{i}\in\mathring{M}_{\mathbf{w}_{i}^{\prime}}$, with some $\mathbf{w}_{i}^{\prime}\in W_{\Gamma}$ and $\|x_{0}\|\cdot\|y_{0}-y\|_{2}\leq\varepsilon$. Put $l_{1}:=\max_{1\leq i\leq K_{1}}|\mathbf{w}_{i}|$, $l_{2}:=\max_{1\leq i\leq K_{2}}|\mathbf{w}_{i}^{\prime}|$ and $l=\max\{l_{1},l_{2}\}$. Let $n\geq l$ so that by (8.6) and linearity we have $\mathbb{E}_{M_{\Lambda}}(x_{0}u_{n}y_{0})=0$ and hence
(8.7) $\mathbb{E}_{M_{\Lambda}}(xu_{n}y)=\mathbb{E}_{M_{\Lambda}}((x-x_{0})u_{n}y)+\mathbb{E}_{M_{\Lambda}}(x_{0}u_{n}(y-y_{0})).$
Furthermore,
(8.8) $\|(x-x_{0})u_{n}y\|_{2}$ $\leq\|x-x_{0}\|_{2}\cdot\|u_{n}y\|\leq\varepsilon,$
(8.9) $\|x_{0}u_{n}(y-y_{0})\|_{2}$ $\leq\|x_{0}u_{n}\|\cdot\|y-y_{0}\|_{2}\leq\varepsilon.$
Thus, as the conditional expectation $\mathbb{E}_{M_{\Lambda}}$ is $\|\cdot\|_{2}$-decreasing (this follows from the Schwarz inequality [57, Proposition 3.3] as $\mathbb{E}_{M_{\Lambda}}$ is trace-preserving and u.c.p.), we obtain for $n\geq l$ that
(8.10) $\|\mathbb{E}_{M_{\Lambda}}(xu_{n}y)\|_{2}\leq 2\varepsilon.$
This shows for any $x,y\in M_{\Gamma}$ that $\|\mathbb{E}_{M_{\Lambda}}(xu_{n}y)\|_{2}\to 0$ as $n\to\infty$. By Definition 2.1(2) this means that $M_{\Gamma}\not\prec_{M_{\Gamma}}M_{\Lambda}$. ∎
Theorem 8.11.
Let $\Gamma$ be a irreducible finite simple graph of size $|\Gamma|\geq 3$ and for $v\in\Gamma$, let $M_{v}$ ($\not$$\mathbb{C}$) be a von Neumann algebra with a normal faithful trace $\tau_{v}$ such that there exists a unitary $u_{v}\in M_{v}$ with $\tau_{v}(u_{v})=0$. Then $M_{\Gamma}$ is a prime factor.
Proof.
It follows from [16, Theorem E] and our assumptions that $M_{\Gamma}$ is a II1-factor. Furthermore, by Lemma 8.10 we have that $M_{\Gamma}\not\prec_{M_{\Gamma}}M_{\Lambda}$ for any strict subgraph $\Lambda\subsetneq\Gamma$. Hence, by Lemma 8.3 we obtain that $M_{\Gamma}$ is either prime or amenable. Since $\Gamma$ is irreducible and $|\Gamma|\geq 3$ it follows from Proposition 7.8 that $M_{\Gamma}$ is non-amenable. Hence, $M_{\Gamma}$ is prime. ∎
Theorem 8.12.
Let $\Gamma$ be a simple graph. For $v\in\Gamma$, let $M_{v}$ ($\not$$\mathbb{C}$) be a von Neumann algebra with a normal faithful trace $\tau_{v}$ and assume that $M_{\Gamma}=*_{v,\Gamma}(M_{v},\tau_{v})$ is a II1-factor. Then $M_{\Gamma}$ is prime if and only if there is an irreducible component $\Lambda\subseteq\Gamma$ for which $M_{\Lambda}$ is prime and $M_{\Gamma\setminus\Lambda}$ is finite-dimensional.
Proof.
Suppose there is an irreducible component $\Lambda\subseteq\Gamma$ with $M_{\Lambda}$ prime and with $\dim M_{\Gamma\setminus\Lambda}<\infty$. Then the factor $M_{\Gamma}=M_{\Lambda}\overline{\otimes}M_{\Gamma\setminus\Lambda}$ is prime since it is a matrix amplification of $M_{\Lambda}$.
For the other direction, suppose that $M_{\Gamma}$ is a prime factor. Denote
$\Lambda:=\{v\in\Gamma:\textrm{Star}_{\Gamma}(v)\not=\Gamma\text{ or }\dim M_{v}=\infty\}.$
If $w\in\Gamma\setminus\Lambda$ then $\textrm{Star}_{\Gamma}(w)=\Gamma$ and $\dim M_{w}<\infty$, so $w\in\textrm{Link}_{\Gamma}(\Lambda)$. Hence $\textrm{Link}_{\Gamma}(\Lambda)=\Gamma\setminus\Lambda$ and so $M_{\Gamma}=M_{\Lambda}\overline{\otimes}M_{\Gamma\setminus\Lambda}$. Now, since $\Gamma\setminus\Lambda$ is complete, and since $\dim M_{v}<\infty$ for $v\in\Gamma\setminus\Lambda$ we have that $M_{\Gamma\setminus\Lambda}$ is finite-dimensional. Hence, since $M_{\Gamma}$ is a prime factor also $M_{\Lambda}$ is a prime factor.
We now show that the graph $\Lambda$ is irreducible so that from $\textrm{Link}_{\Gamma}(\Lambda)=\Gamma\setminus\Lambda$ it follows that $\Lambda$ is an irreducible component of $\Gamma$. Suppose there is a non-empty subgraph $\Lambda_{1}\subseteq\Lambda$ s.t. $\Lambda_{2}:=\Lambda\setminus\Lambda_{1}$ is non-empty and $\textrm{Link}_{\Lambda}(\Lambda_{1})=\Lambda_{2}$. We show a contradiction. We can write $M_{\Lambda}=M_{\Lambda_{1}}\overline{\otimes}M_{\Lambda_{2}}$. Hence, by primeness of the factor $M_{\Lambda}$ there is $i\in\{1,2\}$ s.t. $\dim M_{\Lambda_{i}}<\infty$. Let $v\in\Lambda_{i}$. Since $\dim M_{\Lambda_{i}}<\infty$ we have $\dim M_{v}<\infty$. Hence, since $v\in\Lambda$ we have by definition of $\Lambda$ that $\textrm{Star}_{\Gamma}(v)\not=\Gamma$. Let $w\in\Gamma\setminus\textrm{Star}_{\Gamma}(v)$. Then $\textrm{Star}_{\Gamma}(w)\not=\Gamma$ so that $w\in\Lambda$. Furthermore, $w\not\in\textrm{Link}_{\Gamma}(v)$ so that $w\not\in\textrm{Link}_{\Lambda}(\Lambda_{i})=\Lambda\setminus\Lambda_{i}$, i.e. $w\in\Lambda_{i}$. Hence, since the vertices $v,w$ in $\Lambda_{i}$ share no edge we have $\dim M_{\Lambda_{i}}=\infty$, which is a contradiction. Thus $\Lambda$ is irreducible. ∎
9. Classification of free indecomposability for graph products
In this section we study free-indecomposability for graph product of II1-factors. In Theorem 9.1 we characterize for graph products of II1-factors (with separable predual) when they can decompose as tracial free products of II1-factors. In Theorem 9.2 we combine this result with Theorem 6.18 to show unique free product decompositions for von Neumann algebras in the class $\mathcal{C}_{{\rm Rigid}}\setminus\mathcal{C}_{{\rm Vertex}}$. Hereafter, we show that Theorem 9.1 and Theorem 9.2 really cover new examples. Indeed, in Proposition 9.3 we give sufficient conditions for a graph product to not posses a Cartan-subalgebra, which in Remark 9.4 we use to give examples of freely indecomposable von Neumann algebras $M\in\mathcal{C}_{{\rm Rigid}}\setminus\mathcal{C}_{{\rm Vertex}}$ that are not in the class $\mathcal{C}_{\text{anti-free}}$ from [42]. In Remark 9.5 we show that the unique free product decomposition from Theorem 9.2 also covers new examples.
Theorem 9.1.
Let $\Gamma$ be a simple graph of size $|\Gamma|\geq 2$, and for $v\in\Gamma$ let $(M_{v},\tau_{v})$ be tracial II1-factor with separable predual. Then the graph product $M_{\Gamma}:=*_{v,\Gamma}(M_{v},\tau_{v})$ can decompose as a tracial free product $M_{\Gamma}=(M_{1},\tau_{1})*(M_{2},\tau_{2})$ of II1-factors $M_{1}$$M_{2}$ if and only if $\Gamma$ is not connected.
Proof.
Let $\Gamma$ and $(M_{v},\tau_{v})_{v\in\Gamma}$ be given as stated. If $\Gamma$ is not connected then for any connected component $\Gamma_{0}$ of $\Gamma$ we have $M_{\Gamma}=(M_{\Gamma_{0}},\tau_{1})*(M_{\Gamma\setminus\Gamma_{0}},\tau_{2})$, which shows one direction.
For another direction suppose that $\Gamma$ is connected. Assume we can write $M=(M_{1},\tau_{1})*(M_{2},\tau_{2})$ for some II1-factors $M_{1},M_{2}$. Fix $v\in\Gamma$ and by [56, Proposition 13] let $N_{0}\subseteq M_{v}$ be an amenable II1-subfactor with $N_{0}^{\prime}\cap M_{\Gamma}=M_{v}^{\prime}\cap M_{\Gamma}$. Then $N_{0}$ is amenable relative to $M_{i}$ inside $M$ for $i=1,2$. Therefore, by Theorem 7.3 one of the following holds true:
  1. (1)
    $N_{0}\prec_{M_{\Gamma}}\mathbb{C}1_{M_{\Gamma}}$;
  2. (2)
    $\operatorname{Nor}_{M_{\Gamma}}(N_{0})^{\prime\prime}\prec_{M_{\Gamma}}M_{i}$ for some $1\leq i\leq 2$;
  3. (3)
    $\operatorname{Nor}_{M_{\Gamma}}(N_{0})^{\prime\prime}$ is amenable relative to $\mathbb{C}1_{M_{\Gamma}}$ inside $M_{\Gamma}$.
Since $N_{0}$ is diffuse, we can not have (1).
We show that (2) is also not satisfied. Suppose $\operatorname{Nor}_{M_{\Gamma}}(N_{0})^{\prime\prime}\prec_{M_{\Gamma}}M_{1}$. Since $N_{0}$ is diffuse, it can not embed in any type I factor. It follows that $\operatorname{Nor}_{M_{\Gamma}}(N_{0})^{\prime\prime}\not\prec_{M_{\Gamma}}M_{\varnothing}$. Therefore, since $N_{0}$ and $N_{0}^{\prime}\cap M_{\Gamma}=M_{v}^{\prime}\cap M_{\Gamma}=M_{\textrm{Link}(v)}$ are factors we obtain by Theorem 5.11 that $u^{*}\operatorname{Nor}_{M_{\Gamma}}(N_{0})^{\prime\prime}u\subseteq M_{1}$ for some unitary $u\in M_{\Gamma}$.
Now take $w\in\Gamma$ arbitrarily. Since $\Gamma$ is connected there is a path $P$ from $v$ to $w$, i.e. $P=(v_{0},v_{1},\ldots,v_{n})$ for some $n\geq 0$ and vertices $v_{0},v_{1},\ldots,v_{n}\in\Gamma$ such that $v_{i}\in\textrm{Link}(v_{i-1})$ for $1\leq i\leq n$ and such that $v_{0}=v$ and $v_{n}=w$. As $|\Gamma|\geq 2$ we can moreover choose this path such that it has length $n\geq 1$.
For $i\in\{1,\ldots,n\}$, denote $N_{i}:=M_{v_{i}}$. Then, since $v_{i}\in\textrm{Link}(v_{i-1})$ we obtain $N_{i}\subseteq\operatorname{Nor}_{M_{\Gamma}}(N_{i-1})^{\prime\prime}$. Since $u^{*}\operatorname{Nor}_{M_{\Gamma}}(N_{0})^{\prime\prime}u\subseteq M_{1}$ we obtain $u^{*}N_{1}u\subseteq M_{1}$. Then since $u^{*}N_{1}u\not\prec_{M_{\Gamma}}M_{\varnothing}$ (since $u^{*}N_{1}u$ is diffuse) we obtain by Proposition 5.8(1b) that $\operatorname{Nor}_{M_{\Gamma}}(u^{*}N_{1}u)^{\prime\prime}\subseteq M_{1}$. Now, observe that $\operatorname{Nor}_{M_{\Gamma}}(u^{*}N_{1}u)=u^{*}\operatorname{Nor}_{M_{\Gamma}}(N_{1})u$ and hence $\operatorname{Nor}_{M_{\Gamma}}(u^{*}N_{1}u)^{\prime\prime}=u^{*}\operatorname{Nor}_{M_{\Gamma}}(N_{1})^{\prime\prime}u$. We thus obtain that $u^{*}\operatorname{Nor}_{M_{\Gamma}}(N_{1})^{\prime\prime}u\subseteq M_{1}$. Continuing in this way we obtain $u^{*}\operatorname{Nor}_{M_{\Gamma}}(N_{i})^{\prime\prime}u\subseteq M_{1}$ for all $0\leq i\leq n$. Thus, in particular $u^{*}M_{w}u\subseteq u^{*}\operatorname{Nor}_{M_{\Gamma}}(N_{n-1})^{\prime\prime}u\subseteq M_{1}$. Since $w$ was arbitrary, we obtain that $M_{w}\subseteq uM_{1}u^{*}$ for each $w\in\Gamma$. But this implies $M_{\Gamma}=(\bigcup_{w\in\Gamma}M_{w})^{\prime\prime}\subseteq uM_{1}u^{*}$. Hence $M_{\Gamma}=M_{1}$, which is a contradiction. We conclude that $\operatorname{Nor}_{M_{\Gamma}}(N_{0})^{\prime\prime}\not\prec_{M_{\Gamma}}M_{1}$. By symmetry also $\operatorname{Nor}_{M_{\Gamma}}(N_{0})^{\prime\prime}\not\prec_{M_{\Gamma}}M_{2}$. We obtain that (2) is not satisfied.
We conclude that (3) is satisfied, i.e. $\operatorname{Nor}_{M_{\Gamma}}(N_{0})^{\prime\prime}$ is amenable. Hence $M_{\textrm{Link}(v)}\subseteq\operatorname{Nor}_{M_{\Gamma}}(N_{0})^{\prime\prime}$ is amenable as well. Therefore, by Proposition 7.8 we obtain that $\textrm{Link}(v)$ is a clique and that $M_{w}$ is amenable for any $w\in\textrm{Link}(v)$. We observe that $v\in\Gamma$ was arbitrary, thus for each vertex $z\in\Gamma$ its $\textrm{Link}(z)$ is a clique. Since $\Gamma$ is connected, it follows that $\Gamma$ is a complete graph. Moreover, for any $v\in\Gamma$ choose $z\in\Gamma\setminus\{v\}$ we have $M_{v}\subseteq M_{\textrm{Link}(z)}$, which shows that $M_{v}$ is amenable. Hence $M_{\Gamma}$ is a tensor product of amenable II1-factors and so $M_{\Gamma}$ is amenable. But the amenable II1-factor can not decompose as a free product of type II1-factors. This gives a contradiction and we conclude that $M_{\Gamma}$ can not decompose as free product of II1-factors. ∎
Theorem 9.2.
Any von Neumann algebra $M\in\mathcal{C}_{{\rm Rigid}}\setminus\mathcal{C}_{{\rm Vertex}}$ can decompose as tracial free product inside $\mathcal{C}_{{\rm Rigid}}\setminus\mathcal{C}_{{\rm Vertex}}$:
(9.1) $M=*_{i\in I}M_{i},$
for some index set $I$ and for every $i\in I$ a II1-factor $M_{i}\in\mathcal{C}_{{\rm Rigid}}\setminus\mathcal{C}_{{\rm Vertex}}$ that can not decompose as any tracial free product of II1-factors.
Furthermore, suppose $M$ has another free product decomposition:
$M=*_{j\in J}N_{j},$
for another index set $J$ and for every $j\in J$ a II1-factor $N_{j}\in\mathcal{C}_{{\rm Rigid}}\setminus\mathcal{C}_{{\rm Vertex}}$ that can not decompose as tracial free product of II1-factors. Then $|I|=|J|$ and there is a bijection $\sigma$ between $J$ and $I$ such that for each $j\in J$, $N_{j}$ is unitarily conjugate to $M_{\sigma(j)}$ in $M$.
Proof.
Since $M\in\mathcal{C}_{{\rm Rigid}}\setminus\mathcal{C}_{{\rm Vertex}}$ we can write $M=M_{\Gamma}=*_{v,\Gamma}(M_{v},\tau_{v})$ for some rigid graph $\Gamma$ of size $|\Gamma|\geq 2$ and some II1-factors $M_{v}\in\mathcal{C}_{{\rm Vertex}}$. Let $\{\Gamma_{i}\}_{i\in I}$ be the connected components of $\Gamma$ for some index set $I$, which are rigid by Remark 3.6. We let $\Pi=\{i\}_{i\in I}$ be the graph with $m$ vertices and no edges. We claim that $|\Gamma_{i}|\geq 2$ for all $i\in\Pi$. Indeed, if $|I|=1$ then $\Pi=\{1\}$ and $\Gamma_{1}=\Gamma$ so that $|\Gamma_{i}|=|\Gamma|\geq 2$ for all $i\in\Pi$. On the other hand, if $|I|\geq 2$ then $\textrm{Link}_{\Pi}(\textrm{Link}_{\Pi}(i))=\Pi\not=\{i\}$ for all $i\in\Pi$, so it follows from Lemma 3.5 and rigidity of $\Gamma_{\Pi}\simeq\Gamma$ that $|\Gamma_{i}|\geq 2$ for all $i\in\Pi$.
We denote $M_{i}:=M_{\Gamma_{i}}\in\mathcal{C}_{{\rm Rigid}}$ for $i\in\Pi$. By Theorem 6.18 and rigidity of $\Gamma_{i}$ and the fact that $|\Gamma_{i}|\geq 2$ it follows that $M_{i}\not\in\mathcal{C}_{{\rm Vertex}}$. Furthermore, since $\Gamma_{i}$ is connected we obtain by Theorem 9.1 that $M_{i}$ can not decompose as tracial free product of II1-factors. By Remark 3.4 we conclude that $M_{\Gamma}=*_{v,\Gamma}(M_{v},\tau_{v})=*_{i,\Pi}(M_{\Gamma_{i}},\tau_{i})=*_{i\in I}M_{i}$ which shows (9.1).
Since for every $j\in J$ $N_{j}\in\mathcal{C}_{{\rm Rigid}}\setminus\mathcal{C}_{{\rm Vertex}}$ we can write $N_{j}=*_{z,\Lambda_{j}}(N_{(j,z)},\tau_{(j,z)})$ where $\Lambda_{j}$ is a rigid graph and $(N_{(j,z)})_{z\in\Lambda_{j}}$ are II1-factors in $\mathcal{C}_{{\rm Vertex}}$. Observe for $j\in J$ that $|\Lambda_{j}|\geq 2$ since $N_{j}\not\in\mathcal{C}_{{\rm Vertex}}$ and that $\Lambda_{j}$ is connected by Theorem 9.1 since $N_{i}$ can not decompose as tracial free product of II1-factors. Let $\Pi^{\prime}=\{j\}_{j\in J}$ be the graph with vertices of all point in $J$ and no edges. Then by Remark 3.4 we have:
$M=*_{j\in J}N_{j}=*_{j,\Pi^{\prime}}(*_{v,\Lambda_{j}}(N_{(j,v)},\tau_{(j,v)}))\simeq*_{w,\Lambda_{\Pi^{\prime}}}(N_{w},\tau_{w})=N_{\Lambda_{\Pi^{\prime}}}.$
Then since $\Lambda_{\Pi^{\prime}}$ is rigid by Lemma 3.5, we obtain by Theorem 6.18 that $\Lambda_{\Pi^{\prime}}\simeq\Gamma$. The connected components of $\Lambda_{\Pi^{\prime}}$ respectively $\Gamma$ are $\{\Lambda_{j}\}_{j\in J}$ respectively $\{\Gamma_{i}\}_{i\in I}$. Hence $|I|=|J|$. Moreover, Theorem 6.18 asserts, for some bijection $\sigma$ of between $J$ and $I$, that $\widetilde{N}_{\Lambda_{j}}$ ($N_{j}$) is unitarily conjugate to $M_{\Gamma_{\sigma(j)}}$ ($M_{\sigma(j)}$) in $M_{\Gamma}$. ∎
We give sufficient conditions for absence of Cartan subalgebras in graph products. We note that in [11] absence of Cartan was studied for right-angled Hecke algebras and that in [23] absence of Cartan was fully characterized for von Neumann algebras associated to graph products of groups.
For a non-empty connected graph $\Gamma$ we define its radius as
(9.2) $\operatorname{Radius}(\Gamma):=\inf_{s\in\Gamma}\sup_{t\in\Gamma}\operatorname{Dist}_{\Gamma}(s,t),$
where $\operatorname{Dist}_{\Gamma}(s,t)$ denotes the minimal length of a path in $\Gamma$ from $s$ to $t$. Furthermore, we set $\operatorname{Radius}(\Gamma)=0$ if $\Gamma$ is empty and set $\operatorname{Radius}(\Gamma)=\infty$ if $\Gamma$ is not connected.
Proposition 9.3.
Let $\Gamma$ be a simple graph with $\operatorname{Radius}(\Gamma)\geq 3$ and for $v\in\Gamma$ let $M_{v}$ be a II1-factor with normal faithful trace $\tau_{v}$. Then $M_{\Gamma}=*_{v,\Gamma}(M_{v},\tau_{v})$ does not possess a Cartan-subalgebra.
Proof.
Suppose $M_{\Gamma}$ has a Cartan subalgebra $A\subseteq M_{\Gamma}$. Fix $v\in\Gamma$. Then $M_{\Gamma}=M_{\textrm{Star}(v)}*_{M_{\textrm{Link}(v)}}M_{\Gamma\setminus\{v\}}$. Since $A$ is amenable, one of the statements of Theorem 7.3 must hold. Since $\operatorname{Radius}(\Gamma)\geq 3$, we have $\textrm{Star}(v)\not=\Gamma$. Hence $\operatorname{Nor}_{M_{\Gamma}}(A)^{\prime\prime}=M_{\Gamma}\not\prec_{M_{\Gamma}}M_{\textrm{Star}(v)}$ and $\operatorname{Nor}_{M_{\Gamma}}(A)^{\prime\prime}=M_{\Gamma}\not\prec_{M_{\Gamma}}M_{\Gamma\setminus\{v\}}$ by Proposition 5.9. Thus we must either have $A\prec_{M_{\Gamma}}M_{\textrm{Link}(v)}$ or $\operatorname{Nor}_{M_{\Gamma}}(A)^{\prime\prime}$ is amenable relative to $M_{\textrm{Link}(v)}$ inside $M_{\Gamma}$. Suppose that $A\prec_{M_{\Gamma}}M_{\textrm{Link}(v)}$ then since $A\not\prec_{M_{\Gamma}}M_{\varnothing}$ we obtain by Proposition 5.8(1b) that $M_{\Gamma}=\operatorname{Nor}_{M_{\Gamma}}(A)^{\prime\prime}\subseteq M_{\Lambda_{\operatorname{emb}}}$ where $\Lambda_{\operatorname{emb}}=\textrm{Link}(v)\cup\bigcup_{w\in\textrm{Link}(v)}\textrm{Link}(w)$. We see that $\operatorname{Radius}(\Lambda_{\operatorname{emb}})\leq 2$ (indeed take as center $v$). Hence, since $\operatorname{Radius}(\Gamma)\geq 3$ we have $M_{\Gamma}=\operatorname{Nor}_{M_{\Gamma}}(A)^{\prime\prime}\subseteq M_{\Lambda_{\operatorname{emb}}}\subsetneq M_{\Gamma}$, a contradiction. We conclude that $\operatorname{Nor}_{M_{\Gamma}}(A)^{\prime\prime}$ ($M_{\Gamma}$) is amenable relative to $M_{\textrm{Link}(v)}$ in $M_{\Gamma}$. Since $v$ was arbitrary we obtain using Theorem 5.3 that $M_{\Gamma}$ is amenable. This is a contradiction since $M_{\Gamma}$ is non-amenable by Proposition 7.8 (since $\operatorname{Radius}(\Gamma)\geq 3$). Thus $M_{\Gamma}$ does not have a Cartan subalgebra. ∎
Remark 9.4.
We argue that we find new classes of finite von Neumann algebras that are freely indecomposable. More precisely we argue that Theorem 9.1 covers von Neumann algebras that are not in the class $\mathcal{C}_{\text{anti-free}}$ from [42]. Indeed, let $\Gamma$ be a graph with $\operatorname{Radius}(\Gamma)\geq 3$ (hence $\Gamma$ is irreducible) and for $v\in\Gamma$ let $M_{v}$ be a II1-factor with separable predual and possessing the Haagerup property. Then the II1-factor $M_{\Gamma}$ does not lie in the class $\mathcal{C}_{\text{anti-free}}$ from [42]. Indeed, (i) $M_{\Gamma}$ is prime by Theorem 8.4, (ii) $M_{\Gamma}$ is full (so no property Gamma) by [16, Theorem E], (iii) $M_{\Gamma}$ does not have a Cartan subalgebra by Proposition 9.3, and (iv) $M_{\Gamma}$ has the Haagerup property (so no property (T) by [24, Theorem 3]) by [14, Theorem 0.2]. If $\Gamma$ is moreover connected and rigid and if $M_{v}$ lies in $\mathcal{C}_{{\rm Vertex}}$ for each $v\in\Gamma$, then $M_{\Gamma}$ lies in $\mathcal{C}_{{\rm Rigid}}$ and can not decompose as free product of II1-factors. As a concrete example, take the cyclic graph $\Gamma=\mathbb{Z}_{n}$ for some $n\geq 6$ and for each $v\in\Gamma$ let $M_{v}=\mathcal{L}(\mathbb{F}_{2})\in\mathcal{C}_{{\rm Vertex}}$ which has the Haagerup property by [10, Theorem 12.2.5]. Then $M_{\Gamma}$ is a II1-factor in $\mathcal{C}_{\textrm{Rigid}}\setminus\mathcal{C}_{\text{anti-free}}$ that can not decompose as a (tracial) reduced free product of II1-factors.
Remark 9.5.
We argue that the unique free product decompositions from Theorem 9.2 are not covered by [42] nor [27]. Indeed, let $\Gamma$ be a simple graph whose connected components $\Gamma_{i}$ for $i=1,\ldots,m$ are of the form $\mathbb{Z}_{n_{i}}$ for some $n_{i}\geq 6$. Observe that $\Gamma$ is rigid. For $v\in\Gamma$ put $M_{v}=\mathcal{L}(\mathbb{F}_{2})\in\mathcal{C}_{{\rm Vertex}}$. Then Theorem 9.2 asserts the unique free product decomposition $M_{\Gamma}=M_{\Gamma_{1}}*\cdots*M_{\Gamma_{m}}$. Since the factors $M_{\Gamma_{i}}$ for $i=1,\ldots,m$ are not in the class $\mathcal{C}_{{\rm anti-free}}$, this result is not covered by [42]. Furthermore, we note for $i=1,\ldots,m$ that the group $_{v,\Gamma_{i}}\mathbb{F}_{2}$ is properly proximal by [26, Proposition 3.7] since $\operatorname{Radius}(\Gamma_{i})\geq 3$. Hence, also [27, Corollary 1.8] does not apply.
10. Graph radius rigidity
In this section we generalize the ideas from the proof of Theorem 9.1 and show that we can, in certain cases, retrieve the radius of the graph $\Gamma$ from the graph product $M_{\Gamma}$. In Section 10.1 we introduce the notion of the radius of a von Neumann algebra. Furthermore, we establish good estimates on $\operatorname{Radius}(M_{\Gamma})$ in terms of the radius of $\Gamma$ whenever the vertex algebras $M_{v}$ posses the property strong (AO). In Section 10.2 we establish similar estimates when the vertex algebras $M_{v}$ are group von Neumann algebras $\mathcal{L}(G_{v})$ of countable icc groups $G_{v}$.
10.1. Radius of von Neumann algebras
We introduce the following definition for a simple graph.
Definition 10.1.
Let $\Gamma$ be a simple graph and let $\Lambda\subseteq\Gamma$ be a subgraph. For $d\in\mathbb{Z}_{\geq 0}$ put
$B_{\Gamma}(\Lambda;d)=\{v\in\Gamma:\operatorname{Dist}_{\Gamma}(v,w)\leq d\text{ for some }w\in\Lambda\}.$
which is the closed ball of size $d$ around $\Lambda$. Furthermore, define $B_{\Gamma}(\Lambda,\infty)=\bigcup_{d\geq 1}B_{\Gamma}(\Lambda,d)$.
We will now introduce a similar definition for von Neumann algebras.
Definition 10.2.
Let $M$ be a diffuse von Neumann algebra and $A\subseteq M$ a diffuse von Neumann subalgebra. For $d\geq 0$ we define the von Neumann algebra $B_{M}(A;d)$ inductively. Put $B_{M}(A;0)=A$ and for $d\geq 1$ define
$B_{M}(A;d)=\left(\bigcup_{\begin{subarray}{c}B\subseteq B_{M}(A;d-1)\\ \text{ diffuse vNa}\end{subarray}}\operatorname{Nor}_{M}(B)\right)^{\prime\prime}$
Moreover, we also define
$B_{M}(A;\infty)=\left(\bigcup_{d\geq 0}B_{M}(A;d)\right)^{\prime\prime}$
We remark for $n,m\in\mathbb{Z}_{\geq 0}$ that $B_{M}(A;n+m)=B_{M}(B_{M}(A;n);m)$.
Recall that the radius of a graph $\Gamma$ was defined in (9.2) and note that it is equal to the infimum of all $d\in\mathbb{Z}_{\geq 0}$ for which there exists a vertex $v\in\Gamma$ with $B_{\Gamma}(v;d)=\Gamma$. In a similar way we can introduce the notion of the radius of a von Neumann algebras.
Definition 10.3.
Let $M$ be a diffuse von Neumann algebra. We define $\operatorname{Radius}(M)$ as the infimum of all $d\in\mathbb{Z}_{\geq 0}$ such that there exists a diffuse, amenable subfactor $A\subseteq M$ for which $A^{\prime}\cap M$ is a non-amenable factor and such that $B_{M}(A;d)=M$.
We remark that the definition of $\operatorname{Radius}(M)$ would be more natural with the relaxation that $A$ can be any diffuse amenable von Neumann subalgebra satisfying $B_{M}(A;d)=M$. However, we need the extra restrictions in order to get appropriate lower bounds on $\operatorname{Radius}(M)$.
Proposition 10.4.
Let $\Gamma$ be a simple graph and let $\Lambda\subseteq\Gamma$ be a subgraph. Let $M_{\Gamma}=*_{v,\Gamma}(M_{v},\tau_{v})$ be a graph product of II1-factors with separable preduals. Then
  1. (1)
    For $d\geq 0$ we have $B_{M_{\Gamma}}(M_{\Lambda};d)=M_{B_{\Gamma}(\Lambda;d)}$;
  2. (2)
    If $\Gamma$ is not complete then $\operatorname{Radius}(M_{\Gamma})\leq\max\{2,\operatorname{Radius}(\Gamma)\}$.
Proof.
(1) The statement holds true for $d=0$ since $B_{M_{\Gamma}}(M_{\Lambda},0)=M_{\Lambda}=M_{B_{\Gamma}(\Lambda,0)}$. We show the statement for $d=1$. Let $A\subseteq M_{\Lambda}$ be amenable and diffuse. Then $A\not\prec_{M_{\Gamma}}\mathbb{C}$. Let $\{\Lambda_{j}\}_{j\in\mathcal{J}}$ be the family $\{\varnothing\}$. Then by Proposition 5.8(1b) we obtain $\operatorname{qNor}_{M_{\Gamma}}(A)^{\prime\prime}\subseteq M_{\Lambda_{\operatorname{emb}}}$ where $\Lambda_{\operatorname{emb}}=\Lambda\cup\bigcup_{v\in\Lambda}\textrm{Link}_{\Gamma}(v)=B_{\Gamma}(\Lambda;1)$. Hence $B_{M}(M_{\Lambda};1)\subseteq M_{B_{\Gamma}(\Lambda;1)}$. To show equality, take $w\in B_{\Gamma}(\Lambda;1)\setminus\Lambda$ and let $v\in\Lambda$ such that $v$ and $w$ share an edge. Let $A\subseteq M_{v}$ be an amenable and diffuse. Then $\operatorname{Nor}_{M_{\Gamma}}(A)^{\prime\prime}\supseteq M_{\textrm{Link}(v)}\supseteq M_{w}$. Hence, $M_{w}\subseteq B_{M_{\Gamma}}(M_{\Lambda};1)$. Hence, we obtain equality. Now let $d\geq 1$ and suppose the statement holds true for d1d-1. Then
$B_{M_{\Gamma}}(M_{\Lambda};d)=B_{M_{\Gamma}}(B_{M_{\Gamma}}(M_{\Lambda};d-1);1)=B_{M_{\Gamma}}(M_{B_{\Gamma}(\Lambda;d-1)};1)=M_{B_{\Gamma}(B_{\Gamma}(\Lambda;d-1);1)}=M_{B_{\Gamma}(\Lambda;d)}$
This proves the statement by induction.
(2) Put $r=\operatorname{Radius}(\Gamma)$. We know $r\geq 1$ and furthermore we may assume $r<\infty$ since otherwise the statement is trivial. Let $v\in\Gamma$ such that $B_{\Gamma}(v;r)=\Gamma$. Observe, since $\Gamma$ is not complete, that $v$ can be chosen such that $\textrm{Link}_{\Gamma}(v)$ is not a clique in $\Gamma$. By [56, Proposition 13] we may let $A\subseteq M_{v}$ be a diffuse amenable subfactor for which $A^{\prime}\cap M_{\Gamma}=M_{v}^{\prime}\cap M_{\Gamma}=M_{\textrm{Link}(v)}$. Thus $A^{\prime}\cap M_{\Gamma}$ is a non-amenable factor. We show that $B_{M_{\Gamma}}(A;r)=M_{\Gamma}$. We see that
$M_{\textrm{Link}(v)}\subseteq\operatorname{Nor}_{M_{\Gamma}}(A)^{\prime\prime}\subseteq B_{M_{\Gamma}}(A;1)\subseteq B_{M_{\Gamma}}(M_{v};1)\subseteq M_{B_{\Gamma}(v;1)}$
Hence,
(10.1) $M_{B_{\Gamma}(\textrm{Link}_{\Gamma}(v);1)}=B_{M_{\Gamma}}(M_{\textrm{Link}(v)};1)\subseteq B_{M_{\Gamma}}(B_{M_{\Gamma}}(A;1);1)\subseteq B_{M_{\Gamma}}(M_{B_{\Gamma}(v;1)};1)=M_{B_{\Gamma}(v;2)}$
Now, observe that $B_{M_{\Gamma}}(A;2)=B_{M_{\Gamma}}(B_{M_{\Gamma}}(A;1);1)$ and $B_{\Gamma}(\textrm{Link}_{\Gamma}(v);1)=B_{\Gamma}(v;2)$. If $r\leq 2$ then $B_{\Gamma}(v;2)=\Gamma$ which shows that $\operatorname{Radius}(M_{\Gamma})\leq 2=\max\{2,r\}$. Thus assume $r\geq 2$. By (10.1) we obtain $B_{\Gamma}(A;2)=M_{B_{\Gamma}(v;2)}$. Thus we obtain
$B_{M_{\Gamma}}(A;r)=B_{M_{\Gamma}}(B_{M_{\Gamma}}(A;2);r-2)=B_{M_{\Gamma}}(M_{B_{\Gamma}(v;2)};r-2)=M_{B_{\Gamma}(v;r)}=M_{\Gamma}$
This shows $\operatorname{Radius}(M_{\Gamma})\leq r=\max\{2,r\}$.
Proposition 10.5.
Let $\Gamma$ be a simple graph. Let $M_{\Gamma}=*_{v,\Gamma}(M_{v},\tau_{v})$ be a graph product of II1-factors $M_{v}$. Let $K\geq 1$ be a constant. Suppose for any amenable diffuse subfactor $A\subseteq M_{\Gamma}$ with $A^{\prime}\cap M_{\Gamma}$ a non-amenable factor there is a subgraph $\Lambda\subseteq\Gamma$ with $\operatorname{Radius}(B_{\Gamma}(\Lambda,1))\leq K$ such that $A\prec_{M}M_{\Lambda}$. Then
$\operatorname{Radius}(\Gamma)-K\leq\operatorname{Radius}(M_{\Gamma}).$
Proof.
Denote $R=\operatorname{Radius}(M_{\Gamma})$. We may assume $R<\infty$. Let $A\subseteq M_{\Gamma}$ be an amenable, diffuse subfactor for which $A^{\prime}\cap M_{\Gamma}$ is a non-amenable factor and for which $B_{M_{\Gamma}}(A;R)=M_{\Gamma}$. By assumption $A\prec_{M_{\Gamma}}M_{\Lambda}$ for some subgraph $\Lambda\subseteq\Gamma$ with $\operatorname{Radius}(B_{\Gamma}(\Lambda;1))\leq K$. Let $\{\Lambda_{j}\}_{j\in\mathcal{J}}$ denote the non-empty familiy $\{\varnothing\}$. Then by Theorem 5.11 we obtain a unitary $u\in M_{\Gamma}$ so that $u^{*}Au\subseteq M_{\Lambda_{\operatorname{emb}}}$ where $\Lambda_{\operatorname{emb}}=B_{\Gamma}(\Lambda,1)$. Hence, for $d\geq 0$ we obtain
$u^{*}B_{M_{\Gamma}}(A,d)u=B_{M_{\Gamma}}(u^{*}Au,d)\subseteq B_{M_{\Gamma}}(M_{B_{\Gamma}(\Lambda;1)};d)=M_{B_{\Gamma}(B_{\Gamma}(\Lambda;1);d)}$
Then
$M_{\Gamma}=u^{*}B_{M_{\Gamma}}(A;R)u\subseteq M_{B_{\Gamma}(B_{\Gamma}(\Lambda;1);R)}$
so that $\Gamma=B_{\Gamma}(B_{\Gamma}(\Lambda;1);R)$. Therefore we obtain
$\operatorname{Radius}(\Gamma)\leq\operatorname{Radius}(B_{\Gamma}(\Lambda;1))+R=K+\operatorname{Radius}(M_{\Gamma}),$
which completes the proof. ∎
Theorem 10.6.
Let $\Gamma$ be a simple graph that is not complete. Let $M_{\Gamma}=*_{v,\Gamma}(M_{v},\tau_{v})$ be a graph product of II1-factors $M_{v}$ that satisfy condition strong (AO) and have separable predual. Then
$\operatorname{Radius}(\Gamma)-2\leq\operatorname{Radius}(M_{\Gamma})\leq\max\{2,\operatorname{Radius}(\Gamma)\}$
In particular this holds true when $M_{\Gamma}$ is a graph products of hyperfinite II1-factors.
Proof.
The upper bound is due to Proposition 10.4(2). To obtain the lower bound we show that the condition of Proposition 10.5 is satisfied with constant $K=2$. Let $A\subseteq M_{\Gamma}$ be amenable and diffuse and such that $A^{\prime}\cap M_{\Gamma}$ is non-amenable. By Theorem 6.15 we obtain $A\prec_{M_{\Gamma}}M_{\Lambda}$ for some non-empty subgraph $\Lambda\subseteq\Gamma$ with $\textrm{Link}(\Lambda)$ non-empty. Let $v\in\textrm{Link}(\Lambda)$. Then $\Lambda\subseteq\textrm{Link}(v)$. Hence, $B_{\Gamma}(\Lambda;1)$ equals $B_{\Gamma}(v,2)$ and has radius at most $2$. This proves the lower bound. ∎
Remark 10.7.
We remark that we can use Theorem 10.6 to distinguish von Neumann algebras coming from graph products. Indeed, let $\Gamma$ and $\Lambda$ be two simple graphs with $2\leq\operatorname{Radius}(\Gamma)<\operatorname{Radius}(\Lambda)-2$. Let $R_{\Gamma}=*_{v,\Gamma}(R_{v},\tau_{v})$ and $R_{\Lambda}=*_{v,\Lambda}(R_{v},\tau_{v})$ be graph products of hyperfinite II1-factors $R_{v}$. Then by Theorem 10.6 we obtain
$\operatorname{Radius}(R_{\Gamma})\leq\operatorname{Radius}(\Gamma)<\operatorname{Radius}(\Lambda)-2\leq\operatorname{Radius}(R_{\Lambda})$
Thus, in particular $R_{\Gamma}\not\simeq R_{\Lambda}$.
10.2. Radius estimates for graph products groups
We now show that the statement of Theorem 10.6 also holds true when the vertex von Neumann algebras $M_{v}$ are group von Neumann algebras $\mathcal{L}(G_{v})$ of countable icc groups (Theorem 10.11). We state the following definitions.
Definition 10.8.
Let $G$ be a countable discrete group and let $\mathcal{S}$ be a family of subgroups of $G$. Then a subset $F\subseteq G$ is called small relative to $\mathcal{S}$ if
$F\subseteq\bigcup_{i=1}^{k}g_{i}G_{i}h_{i}$
for some $k\geq 1$, groups $G_{1},\ldots,G_{k}\in\mathcal{S}$ and elements $g_{1},\ldots,g_{k},h_{1},\ldots,h_{k}\in G$.
Definition 10.9.
Let $G$ be a countable discrete group and let $\mathcal{S}$ be a family of subgroups of $G$. Let $V\subseteq\mathcal{L}(G)$ be a norm bounded subset. We write
$V\subseteq_{\text{approx}}\mathcal{L}(\mathcal{S})$
if for every $\varepsilon>0$ there is a subset $F\subseteq G$ that is small relative to $\mathcal{S}$ and satisfies for all $v\in V$ that $\|v-P_{F}(v)\|_{2}\leq\varepsilon$ (here $P_{F}:\ell^{2}(G)\to\ell^{2}(F)$ denotes the orthogonal projection).
The following proposition is similar to [19, Claim 6.15] and follows from the results in [71]. In the proof we write $(B)_{1}$ for the closed unit ball of the von Neumann algebra $B$.
Proposition 10.10.
Let $\Gamma$ be a simple graph and for $v\in\Gamma$ let $G_{v}$ be a countable icc group. Let $G_{\Gamma}=*_{v,\Gamma}G_{v}$ be the graph product and let $B\subseteq\mathcal{L}(G_{\Gamma})$ be a von Neumann subalgebra for which $B^{\prime}\cap\mathcal{L}(G_{\Gamma})$ is a factor. Let $\{\Lambda_{i}\}_{i\in I}$ be a collection of subgraphs of $\Gamma$ and let $\Lambda=\bigcap_{i}\Lambda_{i}$ be their intersection. If $B\prec_{\mathcal{L}(G_{\Gamma})}\mathcal{L}(G_{\Lambda_{i}})$ for all $i$ then $B\prec_{\mathcal{L}(G_{\Gamma})}\mathcal{L}(G_{\Lambda})$
Proof.
Assume $B\prec_{\mathcal{L}(G_{\Gamma})}\mathcal{L}(G_{\Lambda_{i}})$ for $i\in I$. We show $B\prec_{\mathcal{L}(G_{\Gamma})}\mathcal{L}(G_{\Lambda})$. For $i\in I$ we can by [71, Lemma 2.5] obtain a non-zero projection $q_{i}\in B^{\prime}\cap\mathcal{L}(G_{\Gamma})$ such that for $\mathcal{S}_{i}:=\{G_{\Lambda_{i}}\}$ we have
$(Bq_{i})_{1}\subset_{{\rm approx}}\mathcal{L}(\mathcal{S}_{i})$
Moreover, by [71, Proposition 2.6] we may assume $q_{i}\in\mathcal{Z}(\operatorname{Nor}_{\mathcal{L}(G_{\Gamma})}(B)^{\prime\prime})$. Note that
(10.2) $q_{i}\in\mathcal{Z}(\operatorname{Nor}_{\mathcal{L}(G_{\Gamma})}(B)^{\prime\prime})\cap(B^{\prime}\cap\mathcal{L}(G_{\Gamma}))\subseteq\mathcal{Z}(B^{\prime}\cap\mathcal{L}(G_{\Gamma}))=\mathbb{C}1$
Thus $q_{i}=1$. Denote
$\mathcal{S}=\{\bigcap_{i\in I}h_{i}G_{\Lambda_{i}}h_{i}^{-1}\mid h_{i}\in G_{\Gamma}\text{ for }i\in I\}.$
From [71, Lemma 2.7] it follows that $(B)_{1}\subset_{{\rm approx}}\mathcal{L}(\mathcal{S})$. Then from [1, Proposition 3.4] for each $(h_{i})_{i\in I},h_{i}\in G_{\Gamma}$ there is a subgraph $\Lambda_{0}\subseteq\Lambda$ and $k\in G_{\Gamma}$ such that
$\bigcap_{i\in I}h_{i}G_{\Lambda_{i}}h_{i}^{-1}=kG_{\Lambda_{0}}k^{-1}\subseteq kG_{\Lambda}k^{-1}.$
Thus, putting $\mathcal{S}_{0}=\{G_{\Lambda}\}$ it follows that $(B)_{1}\subseteq_{\text{approx}}\mathcal{L}(\mathcal{S}_{0})$ and hence by [71, Lemma 2.5] we obtain $B\prec_{\mathcal{L}(G_{\Gamma})}\mathcal{L}(G_{\Lambda})$. ∎
Theorem 10.11.
Let $\Gamma$ be a simple graph that is not complete. For $v\in\Gamma$ let $G_{v}$ be a countable icc group. Let $G_{\Gamma}=*_{v,\Gamma}G_{v}$ be the graph product. Then
$\operatorname{Radius}(\Gamma)-2\leq\operatorname{Radius}(\mathcal{L}(G_{\Gamma}))\leq\max\{2,\operatorname{Radius}(\Gamma)\}.$
Proof.
The upper bound on $\operatorname{Radius}(\mathcal{L}(G_{\Gamma}))$ follows immediately from Proposition 10.4 since $\mathcal{L}(G_{v})$ is a II1-factor for $v\in\Gamma$. To prove the lower bound we show the condition of Proposition 10.5 is satisfied with $K=2$. Put $M_{v}=\mathcal{L}(G_{v})$ and let $M_{\Gamma}=*_{v,\Gamma}(M_{v},\tau_{v})=\mathcal{L}(G_{\Gamma})$ be the graph product. Let $R\subseteq M_{\Gamma}$ be an amenable II1-factor for which $R^{\prime}\cap M_{\Gamma}$ is a non-amenable factor. We need to show that $R\prec_{M_{\Gamma}}M_{\Lambda}$ for some $\Lambda\subseteq\Gamma$ with $\operatorname{Radius}(B_{\Gamma}(\Lambda;1))\leq 2$. Let $I$ be the set of all vertices $v$ in $\Gamma$ for which $\operatorname{Nor}_{M_{\Gamma}}(R)^{\prime\prime}$ is amenable relative to $M_{\textrm{Link}_{\Gamma}(v)}$ inside $M_{\Gamma}$. By Theorem 5.3 we obtain that $\operatorname{Nor}_{M_{\Gamma}}(R)^{\prime\prime}$ is amenable relative to $M_{\textrm{Link}_{\Gamma}(I)}$ inside $M_{\Gamma}$. Since $\operatorname{Nor}_{M_{\Gamma}}(R)^{\prime\prime}$ is non-amenable (as it contains $R^{\prime}\cap M_{\Gamma}$), we obtain that $\textrm{Link}_{\Gamma}(I)$ is non-empty. Let $w\in\textrm{Link}_{\Gamma}(I)$. Then $I\subseteq B_{\Gamma}(w;1)$ so that $B_{\Gamma}(I;1)\subseteq B_{\Gamma}(w,2)$. Thus since $w\in B_{\Gamma}(I;1)$ we see that $B_{\Gamma}(I;2)$ has radius at most $2$.
Now let $J\subseteq\Gamma$ be the set of all $v\in\Gamma$ for which $R\prec_{M_{\Gamma}}M_{\Gamma\setminus\{v\}}$. Then since $R^{\prime}\cap\mathcal{L}(G_{\Gamma})$ is a factor we obtain by Proposition 10.10 that $R\prec_{M_{\Gamma}}M_{\Gamma\setminus J}$. Now, if $\Gamma\setminus J\subseteq I$ then $R\prec_{M_{\Gamma}}M_{I}$ which shows that we may take $\Lambda=I$. Thus assume $\Gamma\setminus J\not\subseteq I$. Take $v\in\Gamma\setminus J$ with $v\not\in I$. We can decompose
$M_{\Gamma}=M_{\textrm{Star}(v)}*_{M_{\textrm{Link}(v)}}M_{\textrm{Link}(v)}.$
Since $R$ is amenable we get by Theorem 7.3 that at least one of the following holds true
  1. (1)
    $R\prec_{M_{\Gamma}}M_{\textrm{Link}(v)}$
  2. (2)
    $\operatorname{Nor}_{M_{\Gamma}}(R)^{\prime\prime}\prec_{M_{\Gamma}}M_{\textrm{Star}(v)}$ or $\operatorname{Nor}_{M_{\Gamma}}(R)^{\prime\prime}\prec_{M_{\Gamma}}M_{\Gamma\setminus\{v\}}$.
  3. (3)
    $\operatorname{Nor}_{M_{\Gamma}}(R)^{\prime\prime}$ is amenable relative to $M_{\textrm{Link}(v)}$ inside $M_{\Gamma}$.
Since $v$ is not in $I\cup J$ we must have that $R\prec_{M_{\Gamma}}M_{\textrm{Link}(v)}$ or $\operatorname{Nor}_{M_{\Gamma}}(R)^{\prime\prime}\prec_{M_{\Gamma}}M_{\textrm{Star}(v)}$. Thus in particular we obtain $R\prec_{M_{\Gamma}}M_{\textrm{Star}(v)}$. Observe that $B_{\Gamma}(\textrm{Star}(v);1)$ has radius at most $2$. Hence we may take $\Lambda:=\textrm{Star}(v)$. This finishes the proof. ∎